### v1

# A Bayesian Dirichlet Auto-Regressive Conditional Heteroskedasticity Model for Compositional Time Series

Abstract
We analyze a high-frequency, real-world compositional dataset: Airbnb’s daily currency-fee proportions across multiple regions. These proportions sum to one and exhibit changing levels of volatility, particularly following disruptions such as the COVID-19 pandemic. Standard compositional models often assume constant variance, but our data reveal episodes of elevated uncertainty and clustering in volatility.
Motivated by these empirical patterns, we propose a Bayesian Dirichlet Auto-Regressive Moving Average framework extended with a Dirichlet Auto-Regressive Conditional Heteroskedasticity (DARCH) component—the B-DARMA-DARCH model. This novel approach retains the Dirichlet likelihood, ensuring valid predictions on the simplex, and introduces a time-varying precision parameter via an ARMA recursion. As a result, B-DARMA-DARCH captures both mean dynamics and stochastic volatility in compositional time series.
Simulations and empirical tests on the Airbnb currency data show that B-DARMA-DARCH outperforms simpler Dirichlet ARMA and Bayesian log-ratio VARMA models in forecasting accuracy, residual diagnostics, and interval coverage. By accommodating volatility clustering and structural breaks, B-DARMA-DARCH provides a robust framework for compositional data analysis in finance and beyond, supporting improved risk management and uncertainty quantification in settings where proportions and their volatility are of primary interest.
Keywords: Airbnb, Additive Log Ratio, Bayesian multivariate time series, Compositional time series, Currency Volatility, Data Science, Dirichlet distribution, Finance, GARCH, Generalized ARMA model, Heteroskedasticity, Hospitality Industry, Risk Management, Simplex, Vector ARMA model, Volatility Clustering
1 Introduction
Airbnb operates a two-sided marketplace with millions of listings worldwide, and it accommodates transactions in over 100 currencies. While individual hosts set their own nightly rates, the platform must still forecast these prices to predict total daily revenue. Complicating matters, travelers can pay in a currency of their choosing (often their local currency), resulting in a dynamic compositional time series of daily fee shares across USD, EUR, CAD, and other currencies. This currency composition is crucial for accurately converting booked stays into a consolidated U.S. Dollar (USD) amount—a requirement for company-wide financial reporting, treasury operations, and broader risk management.
To forecast revenue in USD, one must know (i) the nightly prices that hosts set in their local currency, and (ii) the fraction of total bookings ultimately paid in each currency. Even if EUR comprises only a small percentage of daily transactions, that fraction can swing markedly from day to day, amplified by large booking volumes and exchange-rate shifts. Such changes directly affect net USD receipts and thus any revenue forecast that fails to anticipate compositional shifts may lead to significant errors. Moreover, once daily compositions are known, treasury teams can plan FX conversions more efficiently, hedge against unexpected fluctuations, or analyze how macroeconomic conditions influence guest payment preferences.
Each day’s currency proportions must lie in a simplex (they are nonnegative and sum to one), a constraint that complicates standard time-series methods. One classical approach is to use log-ratio transformations, as pioneered by Aitchison (1982) and extended in works such as Cargnoni et al. (1997), Ravishanker et al. (2001), and Mills (2010), among others. Transformed additive log-ratio VARMA (tVARMA) models have been fruitfully applied to compositional data (Brunsdon and Smith, 1998; Mills, 2009, 2010).
In contrast, the Dirichlet distribution models compositional data directly on the simplex. Benjamin et al. (2003) proposed frequentist Generalized ARMA models that could be extended to the simplex. Contributions to Dirichlet time series include Bayesian compositional state space models by Grunwald et al. (1993), da Silva et al. (2011), and da Silva and Rodrigues (2015); the Bayesian Dirichlet Auto-Regressive Moving Average (B-DARMA) model by Katz et al. (2024, 2025); and frequentist Dirichlet ARMA models by Zheng and Chen (2017).
Seminal works in Bayesian time series analysis include West (1996), Prado and West (2010), Barber et al. (2011), Berliner (1996), Pole et al. (2018), and Koop and Korobilis (2010). Bayesian methods have been applied to Vector Auto-Regression (VAR) and Vector Auto-Regression Moving Average (VARMA) models (Spencer, 1993; Uhlig, 1997; Bańbura et al., 2010; Karlsson, 2013). Bayesian generalized linear models (GLMs) for time series data, suitable for counts and categorical responses, have been explored by Brandt and Sandler (2012), Roberts and Penny (2002), Nariswari and Pudjihastuti (2019), Chen and Lee (2016), McCabe and Martin (2005), Berry and West (2020), Fukumoto et al. (2019), Silveira de Andrade et al. (2015), and West (2013).
The generalized auto-regressive conditional heteroskedasticity (GARCH) model (Bollerslev, 1986) is fundamental in modeling financial time series volatility. Extensions include non-normal likelihoods to capture heavy tails and skewness (Nelson, 1991; Engle, 2001). Bauwens et al. (2006) surveyed advancements in multivariate GARCH models.
Motivated by the need to capture both shifting means and evolving volatility in compositional time series—especially under sudden disruptions such as COVID-19—we develop a new class of models called Bayesian Dirichlet Auto-Regressive Moving Average with Dirichlet Auto-Regressive Conditional Heteroskedasticity (B-DARMA-DARCH). Unlike existing Dirichlet ARMA methods that assume a fixed or purely deterministic precision parameter, our framework introduces a Dirichlet Auto-Regressive Conditional Heteroskedasticity (DARCH) component. This addition endows the standard Bayesian Dirichlet ARMA model with a GARCH-like recursion on the precision $\phi_{t}$, enabling it to:
  1. 1.
    Model volatility clustering by letting $\phi_{t}$ depend stochastically on recent innovations and prior volatility, thereby reflecting abrupt regime changes and prolonged uncertainty.
  2. 2.
    Accommodate severe shocks (e.g., the onset of COVID-19) that induce persistent or recurrent surges in variance—phenomena not well captured by constant-precision or deterministic-variance models.
  3. 3.
    Preserve compositional constraints via the Dirichlet likelihood, ensuring forecasted currency shares remain valid proportions over time.
  4. 4.
    Leverage a full Bayesian framework to deliver predictive distributions (and credible intervals) that adapt automatically to time-varying volatility, resulting in more robust risk assessments and interval coverage.
These innovations allow B-DARMA-DARCH to simultaneously handle both systematic and stochastic changes in compositional data, providing a flexible, unified approach to dynamic heteroskedasticity on the simplex—a capability hitherto missing in Bayesian compositional time-series models.
We apply the B-DARMA-DARCH model to Airbnb’s currency fee proportions across different regions, demonstrating its effectiveness in capturing complex temporal dynamics and volatility patterns in real-world compositional data. We conduct simulation studies comparing the B-DARMA-DARCH model with the standard B-DARMA and Bayesian transformed VARMA (B-tVARMA) models. The B-DARMA-DARCH model consistently outperforms the other models in forecast accuracy and residual diagnostics, highlighting its robustness and applicability in scenarios with heteroskedasticity, structural breaks, and regime changes. By directly modeling compositional data within the simplex and incorporating time-varying volatility, our proposed B-DARMA-DARCH model offers a powerful tool for analyzing dynamic compositional time series.
The next section presents the B-DARMA-DARCH model. Section 3 presents simulation studies comparing the B-DARMA-DARCH to both the B-DARMA data model and a Bayesian transformed data normal VARMA model. Section 4 presents analysis of the Airbnb data. The paper closes with a short discussion.
2 A Bayesian Dirichlet Auto-Regressive Moving Average - Dirichlet Auto-Regressive Conditional Heteroskedasticity Model
2.1 DARMA Data model
Consider a $J$-component compositional time series $\{\mathbf{y}_{t}\}_{t=1}^{T}$ of length $T$ indexed by $t=1,\dots,T$, where each observation $\mathbf{y}_{t}=(y_{t1},\ldots,y_{tJ})^{\top}$ lies in the simplex
$\mathcal{S}^{J-1}=\left\{\mathbf{y}\in(0,1)^{J}:\sum_{j=1}^{J}y_{j}=1\right\}.$
We model $\mathbf{y}_{t}$ using the Dirichlet distribution with mean vector $\boldsymbol{\mu}_{t}$ and precision parameter $\phi_{t}>0$
$\mathbf{y}_{t}\mid\boldsymbol{\mu}_{t},\phi_{t}\sim\operatorname{Dirichlet}\left(\phi_{t}\boldsymbol{\mu}_{t}\right),$ (1)
where $\boldsymbol{\mu}_{t}=(\mu_{t1},\ldots,\mu_{tJ})^{\top}\in\mathcal{S}^{J-1}$ with probability density function (pdf)
$p(\mathbf{y}_{t}\mid\boldsymbol{\mu}_{t},\phi_{t})=\frac{\Gamma\left(\phi_{t}\right)}{\prod_{j=1}^{J}\Gamma\left(\phi_{t}\mu_{tj}\right)}\prod_{j=1}^{J}y_{tj}^{\phi_{t}\mu_{tj}-1},$ (2)
for $\mathbf{y}_{t}\in\mathcal{S}^{J-1}$.
2.1.1 Modeling the Mean using the Additive Log-Ratio Transformation
We model $\boldsymbol{\mu}_{t}$ as a function of prior observations $\mathbf{y}_{1},\ldots,\mathbf{y}_{t-1}$, prior means $\boldsymbol{\mu}_{1},\ldots,\boldsymbol{\mu}_{t-1}$ and known covariates $\boldsymbol{x}_{t}$ in a generalized linear model framework. As $\boldsymbol{\mu}_{t}$ and $\mathbf{y}$ are constrained, we model $\boldsymbol{\mu}_{t}$ after reducing dimension using the additive log ratio (alr) link
$\boldsymbol{\eta}_{t}\equiv{\textnormal{alr}}(\boldsymbol{\mu}_{t})=\left(\log\left(\frac{\mu_{t1}}{\mu_{tj^{*}}}\right),\ldots,\log\left(\frac{\mu_{tJ}}{\mu_{tj^{*}}}\right)\right)$ (3)
where $\boldsymbol{\eta}_{t}$ is the linear predictor, a $J-1$-vector taking values in $\mathbb{R}^{J-1}$, $j^{*}$ is a chosen reference component $1\leq j^{*}\leq J$, and the element of $\boldsymbol{\eta}_{t}$ that would correspond to the $j^{*}$th element $\log(\mu_{j^{*}}/\mu_{j^{*}})=0$ is omitted. Given $\boldsymbol{\eta}_{t}$, $\boldsymbol{\mu}_{t}$ is defined by the inverse of equation (3) where $\mu_{tj}=\exp(\eta_{tj})/(1+\sum_{j=1}^{J-1}\exp(\eta_{tj}))$ for $j=1,\dots,J$, $j\neq j^{*}$ and for $j=j*$, $\mu_{tj*}=1/(1+\sum_{j=1}^{J-1}\exp(\eta_{tj}))$.
The alr-transformed mean vector $\mathbf{\eta}_{t}$ is modeled as a Vector Auto-Regressive Moving Average (VARMA) process
$\mathbf{\eta}_{t}=\sum_{p=1}^{P}\boldsymbol{A}_{p}(\text{alr}(\mathbf{y}_{t-p})-\mathbf{X}_{t-p}\boldsymbol{\beta})+\sum_{q=1}^{Q}\boldsymbol{B}_{q}(\text{alr}(\mathbf{y}_{t-q})-\mathbf{\eta}_{t-q})+\mathbf{X}_{t}\boldsymbol{\beta},$ (4)
where $P$ is the VAR lag, $Q$ is the VMA lag, and $t=m+1,\dots,T$ where $m=\max(P,Q)$, $\boldsymbol{A}_{p}$ and $\boldsymbol{B}_{q}$ are $(J-1)\times(J-1)$, $\mathbf{X}_{t}\in\mathbb{R}^{J-1\times r_{\beta}}$ is a matrix of deterministic covariates, and $\boldsymbol{\beta}\in\mathbb{R}^{r_{\beta}}$ is a vector of regression coefficients.
2.1.2 Precision Parameter $\phi_{t}$
In the B-DARMA model, precision parameter $\phi_{t}$ is modeled with log link as a function of an $r_{\gamma}$-vector of covariates $\boldsymbol{z}_{t}$,
$\phi_{t}=\exp(\boldsymbol{z}_{t}\boldsymbol{\gamma}),$ (5)
where $\boldsymbol{\gamma}$ is an $r_{\gamma}$-vector of coefficients. If there are no covariates, $\log\phi_{t}=\gamma$ for all $t$.
2.2 DARCH Process for Precision Parameter $\phi_{t}$
We introduce a DARCH process for the precision parameter $\phi_{t}$, modeled with a log link
$\log(\phi_{t})$ $=\sum_{l=1}^{L}\alpha_{l}(\log(\phi_{t-l})-\mathbf{z}_{t-l}^{\prime}\boldsymbol{\gamma})+\sum_{k=1}^{K}\tau_{k}({\textnormal{alr}}(\mathbf{y}_{t-k})-\boldsymbol{\eta}_{t-k})^{\prime}({\textnormal{alr}}(\mathbf{y}_{t-k})-\boldsymbol{\eta}_{t-k})$
$\quad+\mathbf{z}_{t}^{\prime}\boldsymbol{\gamma},$ (6)
where $\alpha_{l}\in\mathbb{R}$ are auto-regressive coefficients for the log-precision, $\tau_{k}\in\mathbb{R}$ are coefficients associated with past innovations, and $\mathbf{z}_{t}\in\mathbb{R}^{r_{\gamma}}$ are known covariates with coefficients $\boldsymbol{\gamma}\in\mathbb{R}^{r_{\gamma}}$.
For ease of notation, we set $P=L$ and $Q=K$, but this assumption is easily relaxed as shown in section 3. The DARCH component models the volatility in the precision parameter $\phi_{t}$, addressing non-constant variability over time. By capturing volatility clustering and external influences, the DARCH process enhances the model’s ability to understand compositional data dynamics, improving forecasting and risk management.
Alternative log-ratio transformations, such as the centered log-ratio (CLR) or the isometric log-ratio (ILR), can also be used. A detailed discussion of these transformations is provided in the Supplementary Material (Section Alternative Log-Ratio Transformations).
2.3 Joint Predictive Distribution
Define the consecutive observations $\mathbf{y}_{a:b}=(\mathbf{y}_{a},\dots,\mathbf{y}_{b})^{\prime}$ for positive integers $a<b$. To be well defined, linear predictor (4) requires having $m$ previous observations $\mathbf{y}_{(t-m):(t-1)}$, and corresponding linear predictors $\boldsymbol{\eta}_{t-m},\ldots,\boldsymbol{\eta}_{t-1}$. In computing posteriors, we condition on the first $m$ observations $\mathbf{y}_{1:m}$ which then do not contribute to the likelihood. For the corresponding first $m$ linear predictors, on the right hand side of (4), we set $\boldsymbol{\eta}_{1},\ldots,\boldsymbol{\eta}_{m}$ equal to ${\textnormal{alr}}(\mathbf{y}_{1}),\ldots,{\textnormal{alr}}(\mathbf{y}_{m})$ which effectively omits the VMA terms $\boldsymbol{B}_{l}({\textnormal{alr}}(\mathbf{y}_{t-l}))$ from (4) and the MA terms $\tau_{l}$ $({\textnormal{alr}}(\mathbf{y}_{t-l})-\boldsymbol{\eta}_{t-l})^{\prime}({\textnormal{alr}}(\mathbf{y}_{t-l})-\boldsymbol{\eta}_{t-k})$ from (6) when $t-l\leq m$. In contrast, in (4) the VAR terms and $\boldsymbol{X}_{t}\boldsymbol{\beta}$ and in (6) the AR terms and $\mathbf{z}_{t}^{\prime}\boldsymbol{\gamma}$ are well defined for $t=1,\ldots,m$.
Define the $C$-vector $\boldsymbol{\theta}$ of all unknown parameters $\boldsymbol{\theta}=(\boldsymbol{A}_{prs},\boldsymbol{B}_{qrs},\boldsymbol{\beta}^{\prime},\boldsymbol{\gamma}^{\prime},\tau_{j},\alpha_{i})^{\prime}$, where $r,s=1,\ldots,J-1$ index all elements of matrices $\boldsymbol{A}_{p}$ and $\boldsymbol{B}_{q}$, $p=1,\ldots,P$, $q=1,\ldots,Q$, $i=1,\dots,L$, $j=1,\dots,K$ and $C=(P+Q)*(J-1)^{2}+L+K+r_{\beta}+r_{\gamma}$. Our prior beliefs about $\boldsymbol{\theta}$ ($p(\boldsymbol{\theta})$) are updated by Bayes’ theorem to give the posterior
$p(\boldsymbol{\theta}|\mathbf{y}_{1:T})=\frac{p(\boldsymbol{\theta})p(\mathbf{y}_{(m+1):T}|\boldsymbol{\theta},\mathbf{y}_{1:m})}{p(\mathbf{y}_{(m+1):T}|\mathbf{y}_{1:m})},$
[…]
2.4 Differences between B-DARMA, B-tVARMA, and B-DARCH
The differences among the Bayesian-ALR-transformed VARMA (B-tVARMA), B-DARMA, and B-DARCH models lie in their handling of compositional constraints and volatility. The B-tVARMA model transforms compositional data to an unconstrained space, which may introduce distortions if model assumptions like normality are violated. In contrast, the B-DARMA and B-DARCH models directly model data within the simplex using the Dirichlet distribution. For volatility, the B-tVARMA model assumes constant variance unless extended to include heteroskedasticity. The B-DARMA model allows a time-varying precision parameter $\phi_{t}$ modeled by exogenous factors but lacks a stochastic structure, potentially missing stochastic volatility patterns. The B-DARCH model introduces a dynamic $\phi_{t}$ with an ARMA structure, capturing both systematic and stochastic changes in volatility, such as clustering and sudden shifts. While the B-tVARMA model is simpler, it may inadequately represent compositional data and volatility dynamics. The B-DARMA model offers a balance but may miss important volatility features. The B-DARCH model, though more complex, comprehensively accounts for mean dynamics and volatility patterns, enhancing model fit and forecasting accuracy when heteroskedasticity is present.
3 Simulation Study
3.1 Overview
We conducted six simulation studies to evaluate the performance of the B-DARMA(1,0)-DARCH(1,1) model compared to the B-DARMA(1,0) and B-tVARMA(1,0) models. Each study generates data from either a DARMA, DARMA-DARCH, or tVARMA and introduces random shocks to simulate misreported observations (1-3) or redraws values for all parameters (4-6) to simulate a regime shift. Each study simulates 50 datasets, and evaluates out-of-sample forecast accuracy and in-sample residual autocorrelation.
3.2 Simulation Studies Setup
In all six studies, datasets with five compositional components were simulated over $T=100$ time points. The first 60 observations were used for training, and the remaining 40 for testing. For each simulation, a new set of parameters was drawn, giving different intercept vectors, auto-regressive matrices, and precision parameters across simulations.
All six simulation studies have the same underlying mean vector formulation for the latent ALR-transformed process
$\boldsymbol{\eta}_{t}=\boldsymbol{\beta}+\boldsymbol{A}(\text{alr}(\mathbf{y}_{t-1})-\boldsymbol{\beta}).$
The auto-regressive coefficient matrix $\boldsymbol{A}$ was a $4\times 4$ matrix with elements drawn from a uniform distribution, $A_{ij}\sim\mathcal{U}(-0.75,0.75)$, independently for each simulation. For each of the 50 simulations, the elements of the intercept vector $\boldsymbol{\beta}$ were generated by drawing five elements $\beta_{j}^{*},j=1,\dots,5$ from $\mathcal{N}(0.2,0.03^{2})$, then normalized to sum to 1 ($\beta_{j}^{*}=\frac{\beta_{j}^{*}}{\sum_{j^{*}=1}^{5}\beta_{j}^{*}})$, transformed using the ALR transformation to give the $4\times 1$ intercept vector $\boldsymbol{\beta}=\text{alr}(\boldsymbol{\beta}^{*})$. For each simulation, the initial values were
$y_{1j}=.2+\epsilon_{j},\ \epsilon_{j}\sim N(0,.01^{2}),\ j=1,\dots,5$
followed by $y_{1}=\frac{y_{1}}{\sum y_{1}}.$
3.2.1 Simulation Studies 1 & 4: DARMA DGP
In Simulations 1 and 4, the data-generating process (DGP) was a DARMA model, $\mathbf{y}_{t}|\mathbf{\mu}_{t},\phi_{t}\sim\text{Dirichlet}(\phi_{t}\mathbf{\mu}_{t})$ with a constant precision parameter $\log(\phi_{t})=\phi_{0}$. The precision parameter $\phi_{0}$ was drawn independently for each simulation from a uniform distribution between 6 and 7.5, $\phi_{0}\sim\mathcal{U}(6,7.5).$
3.2.2 Simulation Studies 2 & 5: DARCH DGP
The data in the second and fifth simulations were generated using a DARMA-DARCH process, where $\phi_{t}$ was generated
$\log(\phi_{t})=\phi_{0}+\alpha(\log(\phi_{t-1})-\phi_{0})+\tau\sum_{i=1}^{4}(\text{alr}(y_{t-1,i})-\eta_{t-1,i})^{2}.$
The true intercept $\phi_{0}$ was drawn independently for each simulation from a uniform distribution between 6 and 7.5, $\phi_{0}\sim\mathcal{U}(6,7.5).$ We fixed parameters $\alpha=0.8$ and $\tau=-0.95$ to simulate data with high volatility persistence and strong responsiveness to past innovations, creating heteroskedasticity patterns that test the B-DARCH model’s ability to capture dynamic volatility.
3.2.3 Simulation Studies 3 & 6: tVARMA DGP
The data in the third and sixth simulations were generated using a tVARMA process with
$\boldsymbol{\eta}_{t}\sim\mathcal{N}(\boldsymbol{\beta}+\boldsymbol{A}(\text{alr}(\mathbf{y}_{t-1})-\boldsymbol{\beta}),\sigma^{2}\boldsymbol{\Sigma})$
$\mathbf{y}_{t}=\text{alr}^{-1}(\boldsymbol{\eta}_{t}).$
A single precision parameter $\sigma$ was drawn independently for each simulation from $\sigma\sim\mathcal{U}(0.05,0.5).$ The covariance matrix $\boldsymbol{\Sigma}$ was generated to be positive-definite with complex interdependencies $\boldsymbol{\Sigma}=\mathbf{M}^{\top}\mathbf{M},$ where $\mathbf{M}$ was a $4\times 4$ matrix with elements drawn independently from $\mathcal{U}(-0.3,0.3)$ for each simulation.
3.2.4 Simulated Shocks
In simulations 1-3, random shocks were introduced into the training data to simulate misreported observations. Time points were selected by sampling intervals from a Poisson(6) distribution and cumulatively summing them. At these times $l$, $\mathbf{y}_{t}$ was replaced with a random composition generated by independently drawing each component from a Uniform(0,1) distribution and normalizing the vector to sum to 1,
$y_{lj}\sim\mathcal{U}(0,1),\ \mathbf{y}_{\text{misreported}[l]}=\frac{y_{lj}}{\sum_{j=1}^{5}y_{lj}},j=1,\dots,5.$
3.2.5 Regime Shifts
In Simulation 4-6, we introduced regime shifts at a random time point $t_{\text{shift}}$ uniformly selected from [10, 50] by redrawing the intercept vector $\boldsymbol{\beta}$, the auto-regressive matrix $A$ and the precision parameters (either $\phi_{0}$, $\sigma$, $\Sigma$) for the DARMA/DARCH models or tVARMA model. For $t=t_{\text{shift}}$ to $t=t_{\text{shift}}+9$, data were generated using these new parameters, simulating a regime shift in the data-generating process. After this period, parameters reverted to their original values, allowing us to evaluate the models’ robustness to temporary structural changes.
3.3 Model Comparisons
We fitted three models to each simulated dataset for comparison: the B-DARMA(1,0)-DARCH(1,1), the Bayesian transformed VARMA (B-tVARMA(1,0)), and the B-DARMA(1,0) models. All models shared the following priors: the intercepts $\boldsymbol{\beta}$ were given $\text{Normal}(0,0.3^{2})$ priors, and the elements of the auto-regressive matrices $\boldsymbol{A}_{p}$ had $\text{Normal}(0,1)$ priors.
For the B-DARMA(1,0)-DARCH(1,1) model, the initial precision parameter $\phi_{0}$ was assigned a $\text{Normal}(7,1.5^{2})$ prior, and the AR and MA terms for $\phi_{t}$ had $\text{Normal}(0.35,0.5^{2})$ and $\text{Normal}(-0.75,0.5^{2})$ priors, respectively.
[…]
All models were fit using Stan with four chains of 1,000 iterations each (500 warm-up), yielding 2,000 posterior samples.
3.4 Metrics
We assess model performance using Forecast Root Mean Squared Error (FRMSE) and Forecast Mean Absolute Error (FMAE) for each of the five components. These metrics quantify the out-of-sample forecast accuracy
${\text{FRMSE}}_{j}=\left(\frac{1}{40}\sum_{t=61}^{100}\left(y_{tj}-\bar{\mu}_{tj}\right)^{2}\right)^{\frac{1}{2}}$
${\text{FMAE}}_{j}=\frac{1}{40}\sum_{t=61}^{100}\left|y_{tj}-\bar{\mu}_{tj}\right|$
where $\bar{\mu}_{tj}$ represents the posterior mean of the predicted value at time $t$ for component $j$. We then summarize by averaging them across all five components ${\text{FRMSE}}_{mean}=\frac{1}{5}\sum_{j=1}^{5}{\text{FRMSE}}_{j}$ and ${\text{FMAE}}_{mean}=\frac{1}{5}\sum_{j=1}^{5}{\text{FMAE}}_{j}$.
Additionally, we assess the Partial Autocorrelation Function (PACF) (see Supplementary Section Definition of the Partial Autocorrelation Function (PACF)) of the sums of squared standardized residuals to evaluate how well each model captures residual dynamics during the in-sample period. For each model, residuals are standardized according to the model’s underlying distribution (see Supplementary Section Standardized Residuals).
3.5 Results
3.5.1 Simulation Studies 1-3: Shocks
The left side of table 1 shows the forecast performance FRMSE and FMAE metrics for the B-DARMA, B-DARCH, and B-tVARMA models across 50 simulated datasets in three different scenarios.
Across simulation studies 1–3, the B-DARCH model consistently had the lowest FRMSE and FMAE values compared to the second-best model, the B-tVARMA. In simulation study 1, FRMSE and FMAE were 1.6% and 1.4% lower, respectively. In simulation study 2, they were 12.6% and 7.7% lower. In simulation study 3, FRMSE and FMAE were .8% and 3.0% lower than B-tVARMA.
Figure S1 shows the mean PACF values across all simulations for B-DARCH, B-DARMA, and B-tVARMA models. The mean PACF values for the B-DARCH model consistently remain closer to zero across the majority of lags, with fewer pronounced PACF values. Notably, the B-DARMA and B-tVARMA models show large negative PACF values, particularly for early lags (lags 1-5). These large values indicate that residuals in these models have substantial autocorrelation, suggesting less efficient capture of the underlying dependencies across time within the data. In contrast, B-DARCH shows minimal autocorrelation for early lags, reflecting an overall improvement in residual behavior and a better fit to the true data-generating process.
3.5.2 Simulation Studies 4-6: Regime Shifts
The right side of Table 1 shows the forecast performance metrics FRMSE and FMAE for the B-DARMA, B-DARCH, and B-tVARMA models across 50 simulations with regime shifts. The B-DARCH model consistently had the lowest FRMSE and FMAE values, outperforming the second-best model (B-tVARMA) with 14.6% lower FRMSE and 8.4% lower FMAE in simulation study 4, 4.7% and 4.6% lower in simulation study 5, and 10.5% and 4.8% lower in simulation study 6. These results highlight the B-DARCH model’s comparative advantage in adapting to sudden changes in the underlying dynamics.
Figure S2 shows the mean PACF values for simulations 4-6 for B-DARCH, B-DARMA, and B-tVARMA models. The B-DARCH model has positive peaks for early lags, with autocorrelation generally settling closer to zero across subsequent lags. Notably, the magnitude of the PACF values for B-DARCH is smaller than those of the other models, except for Simulation 4, where the B-DARMA model shows a smaller magnitude at lag 1.
In contrast, both B-DARMA and B-tVARMA have strong positive PACF values for the first few lags, which then gradually decrease to negative values at later lags, indicating greater persistence of autocorrelation effects. This pattern suggests that B-DARMA and B-tVARMA models struggle more with handling shifts in the underlying process, resulting in larger autocorrelations in early lags compared to the B-DARCH model.
These results indicate that the B-DARCH model’s dynamic volatility component enables it to adapt more effectively to structural breaks, capturing both systematic and stochastic changes in volatility. In contrast, the B-DARMA and B-tVARMA models, lacking a stochastic volatility structure, were less effective in handling the regime shifts, leading to higher forecast errors.
Model Simulation 1 (DARMA Shock) Simulation 4 (DARMA Regime Shift)
Average FRMSE Average FMAE Average FRMSE Average FMAE
B-DARMA 38.6 2.94 25.3 2.58
B-DARCH 38.0 2.87 18.1 2.19
B-tVARMA 38.6 2.91 21.2 2.37
Model Simulation 2 (DARCH Shock) Simulation 5 (DARCH Regime Shift)
Average FRMSE Average FMAE Average FRMSE Average FMAE
B-DARMA 11.2 1.84 20.4 2.28
B-DARCH 9.79 1.69 18.3 2.09
B-tVARMA 11.3 1.83 19.2 2.19
Model Simulation 3 (tVARMA Shock) Simulation 6 (tVARMA Regime Shift)
Average FRMSE Average FMAE Average FRMSE Average FMAE
B-DARMA 25.6 2.43 32.5 2.59
B-DARCH 24.1 2.24 22.9 2.18
B-tVARMA 24.3 2.31 25.7 2.29
Table 1: Summary of Model Performance Metrics times 100 on test set ($T=40$) across 50 simulations for Simulation Studies 1–6. Average FRMSE is the average Forecasted Root Mean Squared Error across all five components; Average Forecasted Mean Absolute Error is the average FMAE across all five components.
4 Airbnb Data Analysis
The dataset comprises daily records from January 1, 2017, to December 31, 2020, detailing the proportion of fees in U.S. Dollars (USD) originating from various billing currencies. The data is categorized into four anonymized diverse geographic regions, referred to as Regions 1-4. Within each region, the fees are broken down by the top five currencies specific to that region, with the remaining currencies grouped into an “other” category. The top five currencies for each region are fixed for the entire period. The training period spans from January 1, 2017, to June 30, 2020. The subsequent three months (July 1, 2020 to September 30, 2020) are used for validation and model specification, while the final 3 months (October 1, 2020, to December 31, 2020) constitute the test set.
4.1 Exploratory Data Analysis
4.1.1 Regional Currency Dynamics
The currency proportions for each region are shown in Figures 1 and Supplementary Figures S3 and S4. Using the U.S. Dollar (USD) as the reference category, 30-day rolling ALR means are shown in Supplementary Figures S5 and S6 while 30-day rolling ALR variances are shown in Supplementary Figure S7 and Figure 2. The correlations of the ALR values with their lag 1 values is shown in Supplementary Figure S8.
Across all four regions, currencies exhibit seasonal cycles and trends relative to USD. Regions 1 and 3 display pronounced yearly seasonality for multiple currencies, with Region 4 exhibiting the strongest seasonal patterns. Pre-COVID-19 (before 2020), currencies in Regions 1 and 2 generally showed steady levels or consistent declines relative to USD, while in Region 3, the Australian Dollar (AUD) and other currencies displayed distinct upward trends. During the COVID-19 period, starting around January or February of 2020, significant shifts are seen: in Region 1, all currencies except the Canadian Dollar (CAD) saw large declines in their means at the expense of the USD; in Region 2, the Euro (EUR) and Swiss Franc (CHF) experienced the largest increases as the USD decline; in Region 3, AUD continued its sharp rise, while the New Zealand Dollar (NZD) initially declined but rapidly recovered; in Region 4, the Brazilian Real (BRL) experienced sharp increases in its rolling mean before being overtaken by USD, and the Mexican Peso (MXN) peaked before a drastic decline.
Lag 1 correlations show strong autocorrelations within currencies, with high self-correlations observed in EUR and GBP in Region 1 and in BRL, Chilean Peso (CLP), and MXN in Region 4. Positive lag 1 cross correlations were found among currencies in Region 2, except for CAD, while in Region 3, NZD and AUD were strongly correlated, and EUR was negatively correlated with other currencies. In Region 4, negative correlations between BRL and EUR, as well as between CLP and EUR, suggest opposing movements between these currency pairs.
Figure 2 and Supplementary Figure S7 show the 30-day rolling variance of each currency’s ALR-transformed proportions. Rather than a single abrupt change, we observe recurrent episodes of elevated variability, often persisting for several weeks before subsiding. For instance, Region 1 (top left) shows multiple surges for EUR and GBP, culminating in early 2020 and tapering mid-year, while Region 2 (top right) begins with a high-variance phase (notably in CAD and CHF), settles to moderate levels around 0.1–0.2, and then rises again at the onset of 2020. Region 3 (bottom left) experiences shorter bursts around 2018 and late 2019, each followed by another peak in early 2020, suggesting that once volatility increases, it may linger rather than immediately revert. Finally, Region 4 (bottom right) transitions from moderate day-to-day swings to abrupt climbs in BRL and CLP near 2019, with additional spikes in BRL and MXN emerging after 2020. In every region, these heightened-variance intervals emerge at distinct times and exhibit varying durations, indicating that volatility in currency shares is shaped by more than one-off shocks, and can arise repeatedly under evolving market conditions. Such patterns underscore the importance of accommodating dynamic, recurrent shifts in variance for effective modeling and forecasting.
Figure 1: Airbnb data analysis - proportion of fees by billing currency for four regions from Jan 1, 2017 to June 30, 2020. AUD is the Australian dollar, BRL is the Brazillian Real, CAD is the Canadian Dollar,CHF is the Swiss Franc, CLP is the Chilean Peso,EUR is the European Euro, GBP is the Great British Pound, MXN is the Mexican Peso, NZD is the New Zealand Dollar, and USD is the US Dollar.
Figure 2: Airbnb data analysis - proportion of fees by currency- 30-day rolling ALR variance for four regions from Jan 1, 2017 to Dec 31, 2019. USD is the reference component.
4.2 Objective
The primary goal of this analysis is to compare the performance of three model classes: Bayesian Dirichlet Auto-Regressive Moving Average (B-DARMA), Bayesian transformed Vector Auto-Regressive Moving Average (B-tVARMA), and Bayesian Dirichlet Auto-Regressive Moving Average with Dirichlet Auto-Regressive Conditional Heteroskedasticity (B-DARMA-DARCH).
4.3 Comparison Models
The B-DARCH and B-DARMA models share the same model for $\boldsymbol{\eta}$ but differ in their treatment of the precision parameter. In the B-DARCH model, the precision parameter has its own AR and MA components, allowing for dynamic heteroskedasticity that evolves over time. In contrast, the B-DARMA model does not include these components in the precision parameter, relying solely on auto-regressive and moving average terms in the mean structure to account for time dependence.
The third model, B-tVARMA, is a Bayesian normal VARMA model applied to transformed data. It transforms the observed multivariate time series $\mathbf{y}_{t}$ using the additive log-ratio (alr) transformation, resulting in ${\textnormal{alr}}(\mathbf{y}_{t})\in\mathbb{R}^{J-1}$. The transformed data is then modeled as ${\textnormal{alr}}(\mathbf{y}_{t})\sim N(\boldsymbol{\eta}_{t},\mathbf{\Sigma})$, where $\boldsymbol{\eta}_{t}$ has time-varying mean defined in (4), and $\bf{\Sigma}$ is an unknown $J-1\times J-1$ positive definite covariance matrix. The covariance matrix $\bf{\Sigma}$ is modeled through its Cholesky factor $L_{\Sigma}$, scaled by $\sigma$, $\bf{\Sigma}$ $=\sigma^{2}\Omega$.
All models are fit with STAN (Stan Development Team, 2022) in R. We run $4$ chains with $2000$ iterations each with a warm up of $1000$ iterations for a posterior sample of $4000$. Initial values for all parameters are selected randomly from the interval $[-1,1]$.
4.4 Seasonal and Trend Terms
Figure 1 and Supplementary figure S3 show varying levels of yearly and weekly seasonality in the currency data in different regions. There is a recurring yearly pattern driven by popular travel periods and a clear weekly seasonality reflecting variations between days of the week. Consequently, we include weekly and yearly seasonal variables and a trend variable in the predictors $\mathbf{X}_{t}$ and the same weekly and yearly seasonal variables in $\boldsymbol{z}_{t}$. Both $\phi_{t}$ and the elements of $\boldsymbol{\eta}_{t}$ are modeled with a linear trend and Fourier terms for seasonal variables, using pairs of $\left(\sin{\frac{2k\pi t}{w_{\text{season}}}},\cos{\frac{2k\pi t}{w_{\text{season}}}}\right)$ for $k=1,\dots,K_{\text{season}}\leq\frac{w_{\text{season}}}{2}$, with $w_{\text{season}}=7$ for weekly seasonality and $w_{\text{season}}=365.25$ for yearly seasonality. The orthogonality of the Fourier terms aids in model convergence. The B-DARCH, B-DARMA, and B-tVARMA all share the same $\mathbf{X}_{t}$ while the B-DARCH and B-DARMA share the same $\boldsymbol{z}_{t}$.
4.5 Model Specification and Validation
The models are fit using data from January 1, 2017, to June 30, 2020, as the training set. The validation set, spanning July 1, 2020, to September 30, 2020, is used to determine optimal lags and exogenous covariates. The test set from October 1, 2020, to December 31, 2020, is used for final model comparison.
For the initial specification, we fix $P=1$, $Q=0$, and $K_{\text{week}}=3$ for each region, allowing the intra-year Fourier terms to vary. The optimal number of Fourier terms ($K_{\text{year}}$) is selected based on the model’s performance on the validation set. With the optimal ($K_{\text{year}}$) established, we vary $P$ and $Q$ to identify the best-performing orders of the auto-regressive (AR) and moving average (MA) components on the validation set. Using these specifications, the models are retrained up until September 30, 2020, and tested on the independent test set for the B-DARMA, B-DARMA-DARCH, and B-tVARMA models.
4.6 Priors for B-DARCH, B-DARMA, and B-tVARMA
For all models (B-DARCH, B-DARMA, and B-tVARMA), we assume each component’s auto-regressive (AR) and moving average (MA) elements, $a_{rs}$ and $b_{rs}$, are strongest with their own lags. Thus, we assign weakly informative normal priors $N(.4,.5^{2})$ to $a_{rrp},b_{rrq}$ for diagonal elements ($r=1,\dots,4,q=1,\dots,3$), and $N(0,.5^{2})$ for non-diagonal elements in the $\boldsymbol{A}$ and $\boldsymbol{B}$ matrices.
The priors for the Fourier, trend, and intercept regression coefficients are $N(0,1)$, $N(0,.1^{2})$, and $N(0,2^{2})$ respectively. For the B-DARMA and B-DARCH, the precision model predictors $\boldsymbol{z}_{t}$ have the same seasonal terms as the mean models, with corresponding priors for the coefficients in $\boldsymbol{\gamma}$ matching those of the $\boldsymbol{\eta}$ terms. The B-DARCH has a $N(0,1)$ prior on its AR $\alpha$ and MA $\tau$ terms.
For the B-tVARMA, we impose a half normal prior $N^{+}(0,.5^{2})$ on the scale parameter $\sigma$ and the correlation matrix $\Omega$ has a $\text{LKJ}(3)$ prior on its Cholesky factor $L_{\Omega}$.
4.7 Results
4.7.1 Model Specification
Out-of-sample validation results (Supplementary Table S1) show that Region 1 performed best with $K_{\text{year}}=8$ and $(P,Q)=(1,0)$, yielding a average FMAE of 0.0036 and total FRSS of .0173. Region 2 achieved optimal performance with $K_{\text{year}}=10$ and $(P,Q)=(1,1)$, resulting in a average FMAE of .0127 and tRSS of .2018. In Region 3, the optimal model had $K_{\text{year}}=10$ and $(P,Q)=(1,1)$, with a average FMAE of .0283 and total FRSS of 1.1395. Region 4 was best modeled with $K_{\text{year}}=8$ and $(P,Q)=(3,2)$, with a FMAE of .0116 and FRSS of .1553.
4.7.2 Performance on Test Set
The Forecast Mean Absolute Error (FMAE) and Forecast Residual Sum of Squares (FRSS) for each region and major currency are shown in Table 2, while Figures 36 display the corresponding out‐of‐sample forecasts $\bar{\mu}_{tj}$. When errors are aggregated over all currencies within each region, the B-DARCH model consistently delivers the lowest total FMAE and FRSS. For example, in Region 1 the total FMAE (FRSS) for B-DARCH is 2.84 (3.15), which is markedly lower than 7.65 (19.11) for B-DARMA and 5.20 (10.62) for B-tVARMA. In Region 2, the advantage of B-DARCH is even more pronounced: its total FMAE is 13.55 (with an average of 2.26) and total FRSS is 73.93 (average 12.32), compared to 14.71 (2.45) and 81.87 (13.65) for B-DARMA, and 17.13 (2.86) and 104.43 (17.41) for B-tVARMA. Regions 3 and 4 exhibit similar aggregate advantages for B-DARCH.
At the individual-currency level, however, there are still several cases in which either B-DARMA or B-tVARMA yield lower FMAE—particularly in Regions 3 and 4. In Region 2, by contrast, the B-DARCH yields the smallest FMAE across all major currencies (CAD, CHF, EUR, GBP, USD, and “other”). In Regions 3 and 4, there remain a few currencies that favor one of the alternative models. For instance, in Region 3, B-DARMA produces slightly lower FMAE than B-DARCH for JPY (2.17 vs. 2.20) and NZD (2.91 vs. 3.57), while in Region 4, B-DARMA outperforms B-DARCH for EUR (0.58 vs. 1.29) but not for most of the other currencies. Notably, B-tVARMA is especially competitive for certain currencies in Region 1 (e.g., EUR), likely because its assumption of constant variance aligns well with periods of relatively stable volatility. Nonetheless, when the full basket of currencies is considered, the time-varying volatility framework of B-DARCH leads to the most accurate overall forecasts.
Supplementary Table S2 shows the empirical coverage rates of the 95% credible intervals for the forecasts across regions and currencies. The B-DARCH model exhibits relatively stable coverage that is generally closer to the nominal 95% level, with mean coverage rates of 0.92, 0.91, 0.87, and 0.91 in Regions 1, 2, 3, and 4, respectively. By contrast, the B-DARMA model shows greater variability, with mean coverages ranging from 0.82 in Region 1 down to 0.55 in Region 3, and 0.82 in Region 2 and 0.84 in Region 4 B-tVARMA tends to undercover, with mean rates of 0.63, 0.86, 0.69, and 0.85 across Regions 1–4. These results suggest that although B-tVARMA can occasionally yield competitive point forecasts, its assumption of time-invariant volatility often leads to overconfident interval estimates. In contrast, the dynamic volatility specification in B-DARCH not only improves point forecast accuracy but also delivers more reliable interval estimates.
Region Currency FMAE FRSS
B-DARCH B-DARMA B-tVARMA B-DARCH B-DARMA B-tVARMA
1 CAD 0.55 0.87 1.21 0.49 1.07 1.75
1 EUR 0.61 2.19 0.55 0.53 5.23 0.55
1 GBP 0.18 0.43 0.21 0.05 0.22 0.06
1 JPY 0.09 0.18 0.06 0.01 0.04 0.01
1 USD 1.08 3.13 2.52 1.91 11.70 7.65
1 other 0.33 0.85 0.66 0.17 0.85 0.60
1 total 2.84 7.65 5.20 3.15 19.11 10.62
1 average 0.47 1.27 0.87 0.53 3.19 1.77
2 CAD 0.35 0.38 0.37 0.22 0.24 0.27
2 CHF 0.83 0.89 0.96 0.97 1.14 1.36
2 EUR 4.51 4.85 5.51 29.70 34.50 42.20
2 GBP 4.06 4.13 5.32 23.00 24.00 35.30
2 USD 1.28 1.68 1.68 3.04 4.59 4.60
2 other 2.52 2.78 3.29 17.00 17.40 20.70
2 total 13.55 14.71 17.13 73.93 81.87 104.43
2 average 2.26 2.45 2.86 12.32 13.65 17.41
3 AUD 6.29 8.24 7.78 57.50 88.50 80.80
3 EUR 0.63 0.93 0.87 0.52 1.09 0.99
3 JPY 2.20 2.17 2.34 23.70 24.10 22.70
3 NZD 3.57 2.91 3.57 14.60 9.96 14.70
3 USD 0.82 0.79 1.02 1.00 0.98 1.52
3 other 2.69 5.09 3.06 13.30 34.80 15.30
3 total 16.20 20.12 18.64 110.62 159.43 136.01
3 average 2.70 3.35 3.11 18.44 26.57 22.67
4 BRL 2.84 4.36 6.38 10.80 23.60 46.20
4 CLP 2.32 2.35 3.09 7.76 8.45 13.00
4 EUR 1.29 0.58 0.64 2.08 0.51 0.60
4 MXN 1.27 2.85 1.54 2.36 8.71 2.98
4 USD 2.70 4.14 2.94 8.68 19.60 9.93
4 other 0.79 0.97 1.94 0.87 1.46 4.93
4 total 11.21 15.25 16.53 32.55 62.33 77.64
4 average 1.87 2.54 2.76 5.43 10.39 12.94
Table 2: Airbnb data analysis: Forecast Mean Absolute Error (FMAE) and Forecast Residual Sum of Squares (FRSS) for Regions 1 to 4 and top currencies. FMAE and FRSS values are multiplied by $10^{2}$.
Figure 3: Airbnb data analysis - Region 1: 92-day forecasts (blue) with 95% confidence intervals, represented by the shaded regions around the forecast lines, for each of the six currencies, from October 1, 2020, to December 31, 2020, compared to the actual values (red) observed over the preceding six months. The models compared include B-DARCH, B-DARMA, and B-tVARMA.
Figure 4: Airbnb data analysis - Region 2: 92-day forecasts (blue) with 95% confidence intervals, represented by the shaded regions around the forecast lines, for each of the six currencies, from October 1, 2020, to December 31, 2020, compared to the actual values (red) observed over the preceding six months. The models compared include B-DARCH, B-DARMA, and B-tVARMA.
Figure 5: Airbnb data analysis - Region 3: 92-day forecasts (blue) with 95% confidence intervals, represented by the shaded regions around the forecast lines, for each of the six currencies, from October 1, 2020, to December 31, 2020, compared to the actual values (red) observed over the preceding six months. The models compared include B-DARCH, B-DARMA, and B-tVARMA.
Figure 6: Airbnb data analysis - Region 4: 92-day forecasts (blue) with 95% confidence intervals, represented by the shaded regions around the forecast lines, for each of the six currencies, from October 1, 2020, to December 31, 2020, compared to the actual values (red) observed over the preceding six months. The models compared include B-DARCH, B-DARMA, and B-tVARMA.
4.7.3 Residual Analysis
The partial autocorrelation function (PACF) of the sum of standardized squared residuals across all components on the test set was computed for each region and model (Supplementary Figures S9S12). All models exhibit a large peak at lag 1, reflecting immediate persistence in the residuals. However, the magnitude of this spike is noticeably smaller for the B-DARCH model. We consider a PACF value to be significant if it exceeds a threshold of 0.2; any autocorrelation values above this level indicate meaningful residual dependence.
Beyond lag 1, the B-DARCH model shows only one PACF value exceeding the 0.2 significance threshold, suggesting effective capture of short-term dependencies and minimal residual autocorrelation at higher lags. In contrast, both B-DARMA and B-tVARMA have PACF values surpassing the 0.2 threshold at multiple lags beyond lag 1, particularly between lags 2 and 10, indicating that residual dependencies persist across a broader range of lags in these models.
These suggest that the B-DARCH model achieves better residual independence and stability, with remaining residual structure potentially explained by unmodeled market-specific holidays, events, and policy implementations.
4.7.4 Parameter Estimates.
A complete discussion of the posterior densities for the DARCH parameters (auto-regressive $\alpha$ and moving average $\tau$ coefficients) in each region appears in the Supplementary Material (Section Parameter Estimates). Briefly, Regions 2 and 3 exhibit higher $\alpha$ values indicating more persistent volatility, while Region 1 shows moderate persistence ($0.750.25<\alpha<$ 0.75), and Region 4 displays a more complex multi-lag structure. These region-specific patterns underscore the flexibility of our B-DARMA-DARCH approach to accommodate various volatility regimes.
5 Discussion
We introduced the Bayesian Dirichlet Auto-Regressive Moving Average with Dirichlet Auto-Regressive Conditional Heteroskedasticity (B-DARMA-DARCH) model, a novel approach for modeling and forecasting dynamic compositional time series with time-varying volatility. By incorporating a DARCH component into the B-DARMA framework, the model effectively captures both mean dynamics and heteroskedasticity inherent in compositional data.
Simulation studies demonstrated the superior performance of the B-DARMA-DARCH model across six scenarios, including data generated from DARMA, DARCH, and tVARMA processes with random shocks simulating misreported observations and regime changes. The B-DARMA-DARCH model consistently achieved the lowest forecast error metrics and exhibited minimal residual autocorrelation, indicating its robustness in capturing complex volatility dynamics, especially in the presence of heteroskedasticity.
Applying the B-DARMA-DARCH model to Airbnb’s currency fee proportions across different regions, we showed that it effectively captures temporal dynamics and volatility patterns in real-world data. The model generally outperformed the standard B-DARMA and B-tVARMA models in terms of forecast accuracy and residual diagnostics. Notably, the inclusion of the DARCH component allowed the model to account for significant disruptions, such as those caused by the COVID-19 pandemic, highlighting its robustness in the face of structural breaks and extreme events.
While the B-DARMA-DARCH model introduces additional complexity due to the dynamic modeling of the precision parameter, this complexity is justified by the improved model fit, empirical coverage rates, and forecasting accuracy in the presence of heteroskedasticity. The model offers a comprehensive framework that accounts for both mean dynamics and volatility patterns, which is particularly important in financial applications where volatility significantly influences decision-making and risk assessment.
However, the computational demands are higher, necessitating careful tuning of sampling parameters to ensure convergence and efficiency. Future research could focus on optimizing computational strategies or exploring approximate inference methods to mitigate these demands.
Several limitations warrant consideration. The current model models each region independently, even though some currencies are shared across multiple regions. This assumption of independence may not hold in practice, as economic events affecting a shared currency could simultaneously impact multiple regions. Incorporating hierarchical structures or accounting for inter-regional dependencies could enhance model performance by capturing these cross-regional correlations, albeit at the cost of increased computational complexity. Additionally, extending the model to handle zero-inflated compositional data or integrating external covariates that influence both the mean and volatility components could broaden its applicability.
Acknowledgement
The authors thank Sean Wilson, Jess Needleman, Jenny Cheng, Erica Savage, Liz Medina, and Yuling Kuo for helpful discussions, Ellie Mertz, Adam Liss, and Peter Coles for championing the research, and Lauren Mackevich for their indispensable operational support.
Supplementary Materials
Alternative Log-Ratio Transformations
Although we have used the additive log-ratio (ALR) link in equations (3), (4), and (5), other transformations to and from the simplex may be employed. A common choice is the centered log-ratio (clr), defined as
${\textnormal{clr}}(\mathbf{y})\;=\;\Bigl{(}\,\ln\bigl{(}\tfrac{y_{1}}{g(\mathbf{y})}\bigr{)},\;\ln\bigl{(}\tfrac{y_{2}}{g(\mathbf{y})}\bigr{)},\;\dots,\;\ln\bigl{(}\tfrac{y_{J}}{g(\mathbf{y})}\bigr{)}\Bigr{)},$
where $g(\mathbf{y})=(\,y_{1}\,y_{2}\,\cdots\,y_{J})^{1/J}$ is the geometric mean of $\mathbf{y}$. The clr mapping places all $J$ components in a transformed $\mathbb{R}^{J}$ space but introduces the constraint that their sum is zero.
Another option is to use the Isometric Log-Ratio (ilr) transformation. For each index $j$, the ilr-transformed coordinate can be expressed as
$z_{j}=\sqrt{\frac{r_{j}}{r_{j}+1}}\,\ln\biggl{(}\frac{g_{j}(\mathbf{y)}}{\bigl{(}\prod_{i\in H_{j}}y_{i}\bigr{)}^{1/r_{j}}}\biggr{)},$
where $g_{j}(\mathbf{y})$ is defined as the geometric mean of a chosen subset $S_{j}\subseteq\mathbf{y}$, and $H_{j}$ is the complement of $S_{j}$ in $\mathbf{y}$. The value $r_{j}$ is the number of elements in $H_{j}$. The particular way in which $S_{j}$ and $H_{j}$ are selected depends on the specific application and how one wishes to interpret the ilr coordinates.
These mappings are mutually related by linear transformations (Egozcue et al., 2003) and each provides a bijection from the $(J-1)$-dimensional simplex onto a $(J-1)$-dimensional subset of $\mathbb{R}^{J}$ or $\mathbb{R}^{J-1}$. Therefore, if the same priors and model assumptions are consistently adapted to each log-ratio space, the resulting compositional models remain equivalent—differing only in how one embeds the data in a Euclidean domain.
Definition of the Partial Autocorrelation Function (PACF)
The PACF at lag $k$ measures the partial correlation between the SSR at time $t$ and the SSR at time $t-k$, controlling for the effects of intermediate lags $1$ to $k-1$. It is defined as the last coefficient $\phi_{kk}$ in the autoregressive model of order $k$
$\text{SSR}_{t}=\phi_{k1}\text{SSR}_{t-1}+\phi_{k2}\text{SSR}_{t-2}+\dots+\phi_{k,k-1}\text{SSR}_{t-(k-1)}+\phi_{kk}\text{SSR}_{t-k}+\varepsilon_{t}$ (7)
where $\varepsilon_{t}$ is the white noise error term at time $t$. The coefficient $\phi_{kk}$ represents the PACF at lag $k$ and can be calculated recursively using the Yule-Walker equations
$\phi_{kk}=\frac{r_{k}-\sum_{j=1}^{k-1}\phi_{k-1,j}r_{k-j}}{1-\sum_{j=1}^{k-1}\phi_{k-1,j}r_{j}}$
Here, $r_{k}$ is the autocorrelation at lag $k$, and $\phi_{k-1,j}$ are the PACF coefficients from the previous lag $k-1$.
By examining the PACF of the SSRs, we can identify the presence of residual autocorrelation at various lags, which indicates whether the model has adequately captured the temporal dependencies in the data.
Standardized Residuals
For the B-DARMA and B-DARCH models, which are based on the Dirichlet distribution as defined in Section 2, we compute standardized residuals using the standard deviations derived from the Dirichlet parameters.
At each time t, the residuals are calculated as $\text{residual}_{tj}=y_{tj}-\mu_{tj},\quad\text{for }j=1,\dots,J.$ The variance of each component $y_{tj}$ under the Dirichlet distribution, Dirichlet($\phi_{t}\boldsymbol{\mu}_{t}$), is $\operatorname{Var}[y_{tj}]=\frac{\mu_{tj}(1-\mu_{tj})}{\phi_{t}+1}$, with standard deviation $\sigma_{tj}=\sqrt{\frac{\mu_{tj}(1-\mu_{tj})}{\phi_{t}+1}}.$
[…]
By examining the PACF of $\{\text{SSR}\}_{t=1}^{T}$ we assess the presence of autocorrelation in the residuals. A well-specified model should produce residuals with minimal autocorrelation, indicating that temporal dependencies have been adequately captured.
Parameter Estimates
Supplementary Figure S13 shows the posterior densities of the auto-regressive ($\alpha$) and moving average ($\tau$) coefficients for each region. In Region 1, the posterior of $\alpha$ is concentrated between 0.25 and 0.75, peaking around 0.45, indicating moderate volatility persistence; shocks have a noticeable but not prolonged effect on volatility. In Region 2, the $\alpha$ coefficient has posterior mean around 0.8, reflecting strong dependence of current volatility on past volatility, and the $\tau$ coefficient centers around $-2.75$, indicating that large deviations from expected currency proportions significantly increase / decrease future volatility—characteristic of volatility clustering. Region 3 shows $\alpha$ tightly concentrated between 0.85 and 0.90 and $\tau$ centered around $-1.2$ suggesting very strong volatility persistence. In contrast, Region 4 exhibits complex dynamics: $\alpha_{1}$ is centered around $0.15$, indicating small persistence from the first lag; $\alpha_{2}$ is skewed positive and centered around $0.35$, suggesting a stronger influence from the second lag; $\alpha_{3}$ peaks slightly above 0.5, indicating moderate persistence; and the $\tau$ parameter is concentrated near $0.02$, implying minimal influence of past innovations.
Overall, the higher $\alpha$ values in Regions 2 and 3 indicate that volatility in currency fee proportions is more persistent in these regions. The negative $\tau$ coefficients suggest that large deviations—such as those caused by abrupt travel restrictions, policy changes, and fluctuating foreign exchange (FX) rates lead to increased future volatility. This implies that Regions 2 and 3 are more susceptible to external shocks affecting travel and currency markets, resulting in prolonged volatility in currency fee proportions. In contrast, Region 1, with moderate volatility persistence, tends to return to stability more rapidly after shocks, possibly due a more stable currency environment or a larger share of domestic travel. Region 4’s dynamics, characterized by less emphasis on immediate past volatility and innovations, may be attributed to regional factors such as diversified travel sources, differing policy responses, or economic structures.
Figure S1: Average PACF of the sum of squared standardized residual values across all 50 simulations for studies 1-3. B-DARMA-DARCH consistently exhibits lower PACF values beyond lag 1, indicating reduced residual autocorrelation and suggesting that its dynamic precision component better captures volatility shocks in these simulated compositional time series.
Figure S2: Average PACF of the sum of squared standardized residual values across all 50 simulations for studies 4-6. The B-DARMA-DARCH model (red) again shows smaller autocorrelation at higher lags, whereas B-DARMA (green) and B-tVARMA (red) leave more persistent structures in the residuals. This result underlines the advantage of modeling time-varying volatility when faced with abrupt changes in compositional data.
Figure S3: Airbnb data analysis- Proportion of fees by currency for four regions: weekly seasonal behavior. AUD is the Australian dollar, BRL is the Brazillian Real, CAD is the Canadian Dollar,CHF is the Swiss Franc, CLP is the Chilean Peso,EUR is the European Euro, GBP is the Great British Pound, MXN is the Mexican Peso, NZD is the New Zealand Dollar, and USD is the US Dollar. Strong weekly cycles appear in many currencies, motivating the inclusion of weekly seasonal terms in our model’s mean structure.
Figure S4: Airbnb data analysis - proportion of fees by billing currency for four regions from Jan 1, 2020 to June 30, 2020. AUD is the Australian dollar, BRL is the Brazillian Real, CAD is the Canadian Dollar,CHF is the Swiss Franc, CLP is the Chilean Peso,EUR is the European Euro, GBP is the Great British Pound, MXN is the Mexican Peso, NZD is the New Zealand Dollar, and USD is the US Dollar. Distinct spikes and dips across multiple currencies (e.g., sharp increase in BRL or EUR) reflect significant disruptions.
Figure S5: Airbnb data analysis - proportion of fees by currency- 30-day rolling ALR means for four regions from Jan 1, 2020 to June 30, 2020. USD is the reference component.
Figure S6: Airbnb data analysis - proportion of fees by currency- 30-day rolling ALR means for four regions from Jan 1, 2017 to June 30, 2020. USD is the reference component. Apart from seasonal patterns, we see notable multi-year trends and cyclical behaviors (e.g., increasing share of EUR in Region 2, upward drift in AUD in Region 3), underlining the presence of both trend and seasonality in compositional dynamics.
Figure S7: Airbnb data analysis - proportion of fees by currency- 30-day rolling ALR variance for four regions from Jan 1, 2017 to June 30, 2020. USD is the reference component.The data display periods of high variance that sometimes persist for weeks, demonstrating “volatility clustering.” These episodes justify the GARCH-like specification in our B-DARMA-DARCH model, which allows the precision parameter $\phi_{t}$ to track spikes in uncertainty over time.
Figure S8: Airbnb data analysis- correlation between ALR and lagged ALR values for each currency in the four regions. USD is the reference component. High positive autocorrelations in many currencies (e.g., EUR, GBP) highlight the strong temporal persistence. Negative cross-lag correlations for some pairs (e.g., EUR vs. BRL in Region 4) indicate that certain currencies move in opposite directions. Both effects reinforce the need for a multi-component time-series framework with auto- and cross-dependence.
Figure S9: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 1.
Figure S10: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 2.
Figure S11: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 3.
Figure S12: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 4.
Figure S13: Airbnb data analysis: Densities of DARCH model coefficients for all 4 regions. Region 1 was modeled with a DARCH(1,0), Regions 2-3 modeled with a DARCH(1,1), and Region 4 modeled with a DARCH(3,1). Each plot shows the estimated density of a specific $\alpha_{l}$ (ar_phi) or $\tau_{k}$ (ma_phi) parameter. Regions 2 and 3 exhibit high $\alpha$ indicating persistent volatility once spikes occur, whereas Region 4’s coefficients suggest moderate or delayed response
Supplementary Figures and Tables
P Q Fourier FMAE FRSS
1 0 6 0.38 1.75
1 0 8 0.36 1.73
1 0 10 0.45 2.65
1 0 12 6.40 442.25
1 0 14 18.66 3662.84
1 0 16 12.42 1638.80
1 0 18 18.87 3809.00
0 1 8 5.81 404.83
1 1 8 1.05 14.60
1 2 8 0.98 15.02
1 3 8 1.03 16.12
2 1 8 0.40 3.96
2 2 8 6.57 716.43
2 3 8 2.09 60.38
3 1 8 8.17 984.45
3 2 8 8.91 1209.30
3 3 8 1.17 16.30
(a) Region 1
P Q Fourier FMAE FRSS
1 0 6 2.15 49.95
1 0 8 2.10 48.46
1 0 10 1.76 38.08
1 0 12 6.46 343.58
1 0 14 13.31 1621.76
1 0 16 9.23 832.27
1 0 18 13.97 1892.17
0 1 10 4.74 193.91
1 1 10 1.27 20.18
1 2 10 1.28 23.08
1 3 10 2.81 48.12
2 1 10 3.08 263.17
2 2 10 2.43 91.08
2 3 10 2.38 73.18
3 1 10 5.84 344.45
3 2 10 2.67 106.77
3 3 10 5.88 380.70
(b) Region 2
P Q Fourier FMAE FRSS
1 0 6 3.55 149.23
1 0 8 3.49 144.21
1 0 10 3.40 136.33
1 0 12 5.78 360.13
1 0 14 11.49 1281.57
1 0 16 9.37 862.47
1 0 18 12.80 1541.94
0 1 10 10.79 982.42
1 1 10 2.83 113.95
1 2 10 3.22 153.84
1 3 10 3.19 150.57
2 1 10 2.89 226.67
2 2 10 3.12 148.21
2 3 10 2.98 131.34
3 1 10 3.94 222.45
3 2 10 3.55 190.52
3 3 10 3.42 166.84
(c) Region 3
P Q Fourier FMAE FRSS
1 0 6 4.72 216.17
1 0 8 4.64 212.57
1 0 10 4.92 236.31
1 0 12 8.00 490.97
1 0 14 12.05 1058.80
1 0 16 9.57 750.15
1 0 18 12.15 1148.46
0 1 8 9.60 693.31
1 1 8 2.66 76.77
1 2 8 1.63 31.03
1 3 8 1.86 38.07
2 1 8 1.43 47.14
2 2 8 1.50 26.37
2 3 8 1.49 24.93
3 1 8 1.33 20.92
3 2 8 1.16 15.53
3 3 8 1.90 40.07
(d) Region 4
Table S1: Airbnb data analysis: B-DARCH model performance across 4 regions – average Forecast Mean Absolute Error (FMAE) and total Forecast Residual Sum of Squares (FRSS) for the top 6 currencies for the validation set from July 1, 2020 to Sep 30, 2020. FMAE and FRSS values are multiplied by $10^{2}$. The table identifies the optimal $(P,Q)$ and number of Fourier terms for each region, showing how model complexity affects forecasting accuracy. Lower FMAE/FRSS indicate better predictive fit, guiding final model selection for each region.
Table S2: Coverage rates (%) of 95% credible intervals for each region and currency under B-DARCH, B-DARMA, and B-tVARMA. The mean coverage is computed across currencies in each region. The B-DARMA-DARCH coverage tends to be closest to the nominal rate across regions (e.g., 92% in Region 1, 95% in Region 4), whereas B-tVARMA often undercovers due to fixed variance assumptions, and B-DARMA coverage can be inconsistent when volatility spikes.
Region Currency B-DARCH B-DARMA B-tVARMA
1 CAD 0.85 0.88 0.72
EUR 0.83 0.60 0.65
GBP 0.96 0.69 0.56
JPY 1.00 0.94 0.82
USD 0.90 0.94 0.44
other 0.96 0.84 0.57
Mean 0.92 0.82 0.63
2 CAD 0.913 0.913 0.783
CHF 0.967 0.891 0.924
EUR 0.957 0.913 0.967
GBP 0.870 0.772 0.924
USD 0.859 0.641 0.663
other 0.880 0.761 0.870
Mean 0.91 0.82 0.86
3 AUD 0.81 0.29 0.43
EUR 0.93 0.77 0.79
JPY 0.71 0.73 0.78
NZD 0.83 0.33 0.43
USD 0.95 0.81 0.81
other 0.97 0.39 0.89
Mean 0.87 0.55 0.69
4 BRL 0.97 0.87 0.97
CLP 0.69 0.54 0.30
EUR 0.97 1.00 1.00
MXN 0.98 0.90 0.99
USD 0.89 0.74 0.96
other 1.00 0.99 0.90
Mean 0.91 0.84 0.85

### v2

@@
-A Bayesian Dirichlet Auto-Regressive Conditional Heteroskedasticity Model for Compositional Time Series
+A Bayesian Dirichlet Auto-Regressive Conditional Heteroskedasticity Model for Forecasting Currency Shares
 Abstract
-We analyze a high-frequency, real-world compositional dataset: Airbnb’s daily currency-fee proportions across multiple regions.
-These proportions sum to one and exhibit changing levels of volatility, particularly following disruptions such as the COVID-19 pandemic.
-Standard compositional models often assume constant variance, but our data reveal episodes of elevated uncertainty and clustering in volatility.
-Motivated by these empirical patterns, we propose a Bayesian Dirichlet Auto-Regressive Moving Average framework extended with a Dirichlet Auto-Regressive Conditional Heteroskedasticity (DARCH) component—the B-DARMA-DARCH model.
-This novel approach retains the Dirichlet likelihood, ensuring valid predictions on the simplex, and introduces a time-varying precision parameter via an ARMA recursion.
-As a result, B-DARMA-DARCH captures both mean dynamics and stochastic volatility in compositional time series.
-Simulations and empirical tests on the Airbnb currency data show that B-DARMA-DARCH outperforms simpler Dirichlet ARMA and Bayesian log-ratio VARMA models in forecasting accuracy, residual diagnostics, and interval coverage.
-By accommodating volatility clustering and structural breaks, B-DARMA-DARCH provides a robust framework for compositional data analysis in finance and beyond, supporting improved risk management and uncertainty quantification in settings where proportions and their volatility are of primary interest.
+In marketplace finance, the daily mix of billing currencies is compositional data that drives forecasting, reporting, and treasury risk.
+We study Airbnb’s 2017–2020 currency‑fee shares across four regions and present a Bayesian Dirichlet ARMA model with a Dirichlet ARCH volatility component, B‑DARMA‑DARCH.
+The model keeps predictions on the simplex, models mean dynamics with VARMA terms on the additive log‑ratio scale, and lets precision vary over time so volatility can spike during disruptions and settle as conditions normalize.
+We include weekly and yearly seasonality and a trend, and we evaluate against Dirichlet ARMA and transformed‑Gaussian VARMA across simulation designs with misreported observations and temporary regime shifts, then validate on a held‑out window and test in 2020 Q4.
+Across simulations and the application, B‑DARMA‑DARCH delivers the most accurate aggregate forecasts, better‑calibrated intervals, and weaker residual persistence at medium lags.
+Modeling precision as a dynamic process provides a practical, interpretable way to forecast proportions and quantify uncertainty when the noise itself moves.
 Keywords: Airbnb, Additive Log Ratio, Bayesian multivariate time series, Compositional time series, Currency Volatility, Data Science, Dirichlet distribution, Finance, GARCH, Generalized ARMA model, Heteroskedasticity, Hospitality Industry, Risk Management, Simplex, Vector ARMA model, Volatility Clustering
@@
 Transformed additive log-ratio VARMA (tVARMA) models have been fruitfully applied to compositional data (Brunsdon and Smith, 1998; Mills, 2009, 2010).
+Let $Q_{t,c}$ denote gross booking amount in currency $c$ on day $t$, and let $\mathrm{FX}_{t,c\rightarrow USD}$ be the USD price of one unit of $c$. Consolidated revenue in USD can be written as
+$R_{t}^{USD}\;=\;\sum_{c=1}^{J}Q_{t,c}\,\mathrm{FX}_{t,c\rightarrow USD}\;=\;V_{t}\sum_{c=1}^{J}S_{t,c}\,\mathrm{FX}_{t,c\rightarrow USD},\quad S_{t,c}=\frac{Q_{t,c}}{V_{t}},\;V_{t}=\sum_{c=1}^{J}Q_{t,c}.$ (1)
+Thus a USD forecast requires both a scale forecast $V_{t}$ (total nominal volume) and a mix forecast $S_{t,\cdot}$ (currency shares).
+One can pursue a bottom‑up route (model each $Q_{t,c}$ and normalize to obtain $S_{t,\cdot}$) or a top‑down route (model $S_{t,\cdot}$ directly on the simplex).
+Our focus is the latter, because the decision‑relevant object for reporting and hedging is the composition itself: even small swings in $S_{t,\cdot}$ materially change $R_{t}^{USD}$ when multiplied by large $V_{t}$ and moving FX rates.
+Empirically, the share series exhibit recurrent bursts of dispersion (see Figure 1), which motivates an observation‑driven volatility mechanism on the simplex.
 In contrast, the Dirichlet distribution models compositional data directly on the simplex.
@@
 (2011), and da Silva and Rodrigues (2015); the Bayesian Dirichlet Auto-Regressive Moving Average (B-DARMA) model by Katz et al.
-(2024, 2025); and frequentist Dirichlet ARMA models by Zheng and Chen (2017).
+(2024, 2025); Katz and Maierhofer (2025); Katz (2025); and frequentist Dirichlet ARMA models by Zheng and Chen (2017).
 Seminal works in Bayesian time series analysis include West (1996), Prado and West (2010), Barber et al.
@@
 (2006) surveyed advancements in multivariate GARCH models.
-Motivated by the need to capture both shifting means and evolving volatility in compositional time series—especially under sudden disruptions such as COVID-19—we develop a new class of models called Bayesian Dirichlet Auto-Regressive Moving Average with Dirichlet Auto-Regressive Conditional Heteroskedasticity (B-DARMA-DARCH).
+Motivated by the need to capture both shifting means and evolving volatility in compositional time series—especially under sudden disruptions such as COVID-19—we develop a new class of models called Bayesian Dirichlet Auto-Regressive Moving Average with Dirichlet Auto-Regressive Conditional Heteroskedasticity (B-DARCH).
 Unlike existing Dirichlet ARMA methods that assume a fixed or purely deterministic precision parameter, our framework introduces a Dirichlet Auto-Regressive Conditional Heteroskedasticity (DARCH) component.
@@
   1. 1.
-    Model volatility clustering by letting $\phi_{t}$ depend stochastically on recent innovations and prior volatility, thereby reflecting abrupt regime changes and prolonged uncertainty.
+    Model volatility clustering by letting $\phi_{t}$ depend dynamically on recent innovations and prior volatility, thereby reflecting abrupt regime changes and prolonged uncertainty.
   2. 2.
@@
     Leverage a full Bayesian framework to deliver predictive distributions (and credible intervals) that adapt automatically to time-varying volatility, resulting in more robust risk assessments and interval coverage.
-These innovations allow B-DARMA-DARCH to simultaneously handle both systematic and stochastic changes in compositional data, providing a flexible, unified approach to dynamic heteroskedasticity on the simplex—a capability hitherto missing in Bayesian compositional time-series models.
-We apply the B-DARMA-DARCH model to Airbnb’s currency fee proportions across different regions, demonstrating its effectiveness in capturing complex temporal dynamics and volatility patterns in real-world compositional data.
-We conduct simulation studies comparing the B-DARMA-DARCH model with the standard B-DARMA and Bayesian transformed VARMA (B-tVARMA) models.
-The B-DARMA-DARCH model consistently outperforms the other models in forecast accuracy and residual diagnostics, highlighting its robustness and applicability in scenarios with heteroskedasticity, structural breaks, and regime changes.
-By directly modeling compositional data within the simplex and incorporating time-varying volatility, our proposed B-DARMA-DARCH model offers a powerful tool for analyzing dynamic compositional time series.
-The next section presents the B-DARMA-DARCH model.
-Section 3 presents simulation studies comparing the B-DARMA-DARCH to both the B-DARMA data model and a Bayesian transformed data normal VARMA model.
-Section 4 presents analysis of the Airbnb data.
-The paper closes with a short discussion.
+These innovations allow B-DARCH to simultaneously handle both systematic and dynamic changes in compositional data, providing a flexible, unified approach to dynamic heteroskedasticity on the simplex—a capability hitherto missing in Bayesian compositional time-series models.
+We apply the B-DARCH model to Airbnb’s currency fee proportions across different regions, demonstrating its effectiveness in capturing complex temporal dynamics and volatility patterns in real-world compositional data.
+We conduct simulation studies comparing the B-DARCH model with the standard B-DARMA and Bayesian transformed VARMA (B-tVARMA) models.
+The B-DARCH model consistently outperforms the other models in forecast accuracy and residual diagnostics, highlighting its robustness and applicability in scenarios with heteroskedasticity, structural breaks, and regime changes.
+By directly modeling compositional data within the simplex and incorporating time-varying volatility, our proposed B-DARCH model offers a powerful tool for analyzing dynamic compositional time series.
+The remainder of the paper is organized as follows.
+The next section introduces the B-DARCH model.
+Section 3 reports simulation studies comparing B-DARCH with three alternatives: (i) a B-DARMA data model; (ii) a Bayesian transformed-data normal VARMA model; and (iii) a time-varying-parameter comparator in ALR space (B-TVP-tVARMA).
+Section 4 analyzes the Airbnb data.
+The paper concludes with a brief discussion.
 2 A Bayesian Dirichlet Auto-Regressive Moving Average - Dirichlet Auto-Regressive Conditional Heteroskedasticity Model
@@
 We model $\mathbf{y}_{t}$ using the Dirichlet distribution with mean vector $\boldsymbol{\mu}_{t}$ and precision parameter $\phi_{t}>0$
-$\mathbf{y}_{t}\mid\boldsymbol{\mu}_{t},\phi_{t}\sim\operatorname{Dirichlet}\left(\phi_{t}\boldsymbol{\mu}_{t}\right),$ (1)
+$\mathbf{y}_{t}\mid\boldsymbol{\mu}_{t},\phi_{t}\sim\operatorname{Dirichlet}\left(\phi_{t},\boldsymbol{\mu}_{t}\right),$ (2)
 where $\boldsymbol{\mu}_{t}=(\mu_{t1},\ldots,\mu_{tJ})^{\top}\in\mathcal{S}^{J-1}$ with probability density function (pdf)
@@
 $\boldsymbol{\eta}_{t}\equiv{\textnormal{alr}}(\boldsymbol{\mu}_{t})=\left(\log\left(\frac{\mu_{t1}}{\mu_{tj^{*}}}\right),\ldots,\log\left(\frac{\mu_{tJ}}{\mu_{tj^{*}}}\right)\right)$ (4)
-where $\boldsymbol{\eta}_{t}$ is the linear predictor, a $J-1$-vector taking values in $\mathbb{R}^{J-1}$, $j^{*}$ is a chosen reference component $1\leq j^{*}\leq J$, and the element of $\boldsymbol{\eta}_{t}$ that would correspond to the $j^{*}$th element $\log(\mu_{j^{*}}/\mu_{j^{*}})=0$ is omitted.
-Given $\boldsymbol{\eta}_{t}$, $\boldsymbol{\mu}_{t}$ is defined by the inverse of equation (3) where $\mu_{tj}=\exp(\eta_{tj})/(1+\sum_{j=1}^{J-1}\exp(\eta_{tj}))$ for $j=1,\dots,J$, $j\neq j^{*}$ and for $j=j*$, $\mu_{tj*}=1/(1+\sum_{j=1}^{J-1}\exp(\eta_{tj}))$.
-The alr-transformed mean vector $\mathbf{\eta}_{t}$ is modeled as a Vector Auto-Regressive Moving Average (VARMA) process
+where $\boldsymbol{\eta}_{t}$ is the linear predictor, a J1J-1-vector taking values in $\mathbb{R}^{J-1}$, $j^{*}$ is a chosen reference component $1\leq j^{*}\leq J$, and the element of $\boldsymbol{\eta}_{t}$ that would correspond to the $j^{*}$th element $\log(\mu_{j^{*}}/\mu_{j^{*}})=0$ is omitted.
+Given $\boldsymbol{\eta}_{t}$, $\boldsymbol{\mu}_{t}$ is defined by the inverse of equation (4) where $\mu_{tj}=\exp(\eta_{tj})/(1+\sum_{j=1}^{J-1}\exp(\eta_{tj}))$ for $1,\dots,J, jj$$\neq j^{*}$ and for j=jj=j*, $\mu_{tj*}=1/(1+\sum_{j=1}^{J-1}\exp(\eta_{tj}))$.
+The ALR-transformed mean vector $\boldsymbol{\eta}_{t}$ is modeled with an observation-driven VARMA-style recursion that uses autoregressive and moving-average terms built from past ${\textnormal{alr}}(\mathbf{y})$ and past linear predictors:
 $\mathbf{\eta}_{t}=\sum_{p=1}^{P}\boldsymbol{A}_{p}(\text{alr}(\mathbf{y}_{t-p})-\mathbf{X}_{t-p}\boldsymbol{\beta})+\sum_{q=1}^{Q}\boldsymbol{B}_{q}(\text{alr}(\mathbf{y}_{t-q})-\mathbf{\eta}_{t-q})+\mathbf{X}_{t}\boldsymbol{\beta},$ (5)
 where $P$ is the VAR lag, $Q$ is the VMA lag, and $t=m+1,\dots,T$ where $m=\max(P,Q)$, $\boldsymbol{A}_{p}$ and $\boldsymbol{B}_{q}$ are $(J-1)\times(J-1)$, $\mathbf{X}_{t}\in\mathbb{R}^{J-1\times r_{\beta}}$ is a matrix of deterministic covariates, and $\boldsymbol{\beta}\in\mathbb{R}^{r_{\beta}}$ is a vector of regression coefficients.
+Interpretation.
+Equation (5) defines $\boldsymbol{\eta}_{t}$ deterministically conditional on past data and past linear predictors, without introducing a separate innovation term for $\boldsymbol{\eta}_{t}$.
+Hence it is an observation-driven specification in the sense of GARCH-type recursions rather than a parameter-driven state-space evolution.
+A natural alternative would be to endow $\boldsymbol{\eta}_{t}$ with a stochastic transition equation (for example, a linear or nonlinear state-space model with evolution noise) and integrate over the latent path.
+In our Dirichlet likelihood with the DARCH precision recursion, this approach becomes computationally onerous due to the nonlinearity of the simplex mapping and the volatility recursion.
+The observation-driven form in (5) keeps inference tractable while still allowing rich short-run dynamics through the AR terms in ${\textnormal{alr}}(\mathbf{y})$ and the MA terms in $\boldsymbol{\eta}$.
+We complement this with a transformed-Gaussian comparator that does include time-variation on the mean side via time-varying coefficients (B-TVP-tVARMA), thereby illustrating the trade-off between state evolution and observation-driven recursions within our framework.
 2.1.2 Precision Parameter $\phi_{t}$
-In the B-DARMA model, precision parameter $\phi_{t}$ is modeled with log link as a function of an $r_{\gamma}$-vector of covariates $\boldsymbol{z}_{t}$,
-$\phi_{t}=\exp(\boldsymbol{z}_{t}\boldsymbol{\gamma}),$ (5)
+In the B-DARMA model, precision parameter $\phi_{t}$ is modeled with log link as a function of an $r_{\gamma}$-vector of covariates $\mathbf{z}_{t}$,
+$\phi_{t}=\exp(\mathbf{z}_{t}\boldsymbol{\gamma}),$ (6)
 where $\boldsymbol{\gamma}$ is an $r_{\gamma}$-vector of coefficients. If there are no covariates, $\log\phi_{t}=\gamma$ for all $t$.
@@
 where $\alpha_{l}\in\mathbb{R}$ are auto-regressive coefficients for the log-precision, $\tau_{k}\in\mathbb{R}$ are coefficients associated with past innovations, and $\mathbf{z}_{t}\in\mathbb{R}^{r_{\gamma}}$ are known covariates with coefficients $\boldsymbol{\gamma}\in\mathbb{R}^{r_{\gamma}}$.
-For ease of notation, we set $P=L$ and $Q=K$, but this assumption is easily relaxed as shown in section 3.
-The DARCH component models the volatility in the precision parameter $\phi_{t}$, addressing non-constant variability over time.
-By capturing volatility clustering and external influences, the DARCH process enhances the model’s ability to understand compositional data dynamics, improving forecasting and risk management.
+For compactness, we write B-DARCH(P,Q)(P,Q) to denote a B-DARMA mean model of order $(P,Q)$ coupled with a DARCH precision process for $\phi_{t}$ of order $(L,K)$.
+In general, $(L,K)$ need not equal $(P,Q)$; when they differ, we state both explicitly as B-DARMA(P,Q)(P,Q)–DARCH(L,K)(L,K) (e.g., B-DARMA(1,0)(1,0)–DARCH(1,1)(1,1)).
+The DARCH component governs time-variation in the precision $\phi_{t}$, accommodating volatility clustering and covariate effects; see Section 3 for cases with $(L,K)\neq(P,Q)$.
 Alternative log-ratio transformations, such as the centered log-ratio (CLR) or the isometric log-ratio (ILR), can also be used.
-A detailed discussion of these transformations is provided in the Supplementary Material (Section Alternative Log-Ratio Transformations).
+A detailed discussion of these transformations is provided in the Supplementary Material (Section 6).
 2.3 Joint Predictive Distribution
@@
 In contrast, in (5) the VAR terms and $\boldsymbol{X}_{t}\boldsymbol{\beta}$ and in (7) the AR terms and $\mathbf{z}_{t}^{\prime}\boldsymbol{\gamma}$ are well defined for $t=1,\ldots,m$.
-Define the $C$-vector $\boldsymbol{\theta}$ of all unknown parameters $\boldsymbol{\theta}=(\boldsymbol{A}_{prs},\boldsymbol{B}_{qrs},\boldsymbol{\beta}^{\prime},\boldsymbol{\gamma}^{\prime},\tau_{j},\alpha_{i})^{\prime}$, where $r,s=1,\ldots,J-1$ index all elements of matrices $\boldsymbol{A}_{p}$ and $\boldsymbol{B}_{q}$, $p=1,\ldots,P$, $q=1,\ldots,Q$, $i=1,\dots,L$, $j=1,\dots,K$ and $C=(P+Q)*(J-1)^{2}+L+K+r_{\beta}+r_{\gamma}$.
+Define the $C$-vector $\boldsymbol{\theta}$ of all unknown parameters $\boldsymbol{\theta}=(\mathbf{A}_{prs},\mathbf{B}_{qrs},\boldsymbol{\beta}^{\prime},\boldsymbol{\gamma}^{\prime},\tau_{j},\alpha_{i})^{\prime}$, where $r,s=1,\ldots,J-1$ index all elements of matrices $\mathbf{A}_{p}$ and $\mathbf{B}_{q}$, $p=1,\ldots,P$, $q=1,\ldots,Q$, $i=1,\dots,L$, $j=1,\dots,K$ and $C=(P+Q)*(J-1)^{2}+L+K+r_{\beta}+r_{\gamma}$.
 Our prior beliefs about $\boldsymbol{\theta}$ ($p(\boldsymbol{\theta})$) are updated by Bayes’ theorem to give the posterior
 $p(\boldsymbol{\theta}|\mathbf{y}_{1:T})=\frac{p(\boldsymbol{\theta})p(\mathbf{y}_{(m+1):T}|\boldsymbol{\theta},\mathbf{y}_{1:m})}{p(\mathbf{y}_{(m+1):T}|\mathbf{y}_{1:m})},$
-[…]
-[…]
+where $p(\mathbf{y}_{(m+1):T}|\boldsymbol{\theta},\mathbf{y}_{1:m})=\prod_{t=m+1}^{T}p(\mathbf{y}_{t}|\boldsymbol{\theta},\mathbf{y}_{(t-m):(t-1)})$, $p(\mathbf{y}_{t}|\boldsymbol{\theta},\mathbf{y}_{(t-m):(t-1)})$ is the density of the Dirichlet in (3), and the normalizing constant $p(\mathbf{y}_{(m+1):T}|\mathbf{y}_{1:m})=\int p(\boldsymbol{\theta})p(\mathbf{y}_{(m+1):T}|\boldsymbol{\theta},\mathbf{y}_{1:m})d\boldsymbol{\theta}$.
+Our objective is to forecast the upcoming $S$ observations, denoted as $\mathbf{y}_{(T+1):(T+S)}$.
+The joint predictive distribution for these future observations is
+$p(\mathbf{y}_{(T+1):(T+S)}\mid\mathbf{y}_{1:T})=\int_{\boldsymbol{\theta}}p(\mathbf{y}_{(T+1):(T+S)}\mid\boldsymbol{\theta})\,p(\boldsymbol{\theta}\mid\mathbf{y}_{1:T})\,d\boldsymbol{\theta}.$
 2.4 Differences between B-DARMA, B-tVARMA, and B-DARCH
-The differences among the Bayesian-ALR-transformed VARMA (B-tVARMA), B-DARMA, and B-DARCH models lie in their handling of compositional constraints and volatility.
-The B-tVARMA model transforms compositional data to an unconstrained space, which may introduce distortions if model assumptions like normality are violated.
-In contrast, the B-DARMA and B-DARCH models directly model data within the simplex using the Dirichlet distribution.
-For volatility, the B-tVARMA model assumes constant variance unless extended to include heteroskedasticity.
-The B-DARMA model allows a time-varying precision parameter $\phi_{t}$ modeled by exogenous factors but lacks a stochastic structure, potentially missing stochastic volatility patterns.
-The B-DARCH model introduces a dynamic $\phi_{t}$ with an ARMA structure, capturing both systematic and stochastic changes in volatility, such as clustering and sudden shifts.
-While the B-tVARMA model is simpler, it may inadequately represent compositional data and volatility dynamics.
-The B-DARMA model offers a balance but may miss important volatility features.
-The B-DARCH model, though more complex, comprehensively accounts for mean dynamics and volatility patterns, enhancing model fit and forecasting accuracy when heteroskedasticity is present.
+The three models differ in how they respect the simplex and how they treat time‑varying volatility.
+B‑tVARMA moves the composition to the ALR space and fits a Gaussian VARMA with constant observation covariance unless explicitly extended.
+This keeps estimation simple but can misalign with the simplex after back‑transformation if the Gaussian–ALR assumptions are strained.
+B‑DARMA keeps the Dirichlet likelihood on the simplex; its precision $\phi_{t}$ is a deterministic regression in covariates, so dispersion may change with seasonality or trend but does not react to forecast errors.
+B‑DARCH also retains the Dirichlet likelihood but replaces the fixed precision with a GARCH‑type, observation‑driven recursion for log $\phi_{t}$.
+Volatility is updated dynamically from past residuals and past precision, allowing clustering and short‑run bursts in dispersion while preserving the compositional support.
+This is time‑varying volatility in an observation‑driven sense, not a parameter‑driven stochastic‑volatility state‑space model.
+In practice, B‑tVARMA is parsimonious but may under‑represent compositional and volatility features; B‑DARMA captures the simplex with covariate‑driven variance; and B‑DARCH adds an adaptive volatility mechanism that improves fit and forecasting when conditional heteroskedasticity matters.
+Assumption sensitivity and alternatives.
+Both B-tVARMA and the Dirichlet-based B-DARMA/B-DARCH are parametric choices that can distort if their distributional assumptions are violated.
+On the ALR scale, B-tVARMA assumes Gaussian residuals; heavy tails or skewness can lead to bias on back-transformation to the simplex.
+For B-DARMA and B-DARCH, the Dirichlet likelihood imposes a specific mean–variance relationship and a restricted covariance structure ($\operatorname{Cov}(y_{i},y_{j})=-\mu_{i}\mu_{j}/(\phi_{t}+1)$ for $i\neq j$), which rules out positive contemporaneous correlations among components.
+When the empirical dispersion or dependence deviates from these constraints, Dirichlet-based models can also misrepresent uncertainty or cross-component co-movement.
+Other families for proportions are available, including logistic-normal models on a log-ratio scale, generalized Dirichlet specifications, mixtures (e.g., Dirichlet mixtures), and zero-augmented variants for sparse compositions.
+We focus on the Dirichlet class because it guarantees simplex-valid predictions and enables a tractable, observation-driven volatility mechanism via the DARCH recursion.
+The transformed-Gaussian comparator (B-tVARMA) provides a complementary baseline with a richer covariance structure in ALR space.
+In applications where positive cross-correlations or more flexible dispersion are essential, generalized Dirichlet or logistic-normal formulations are natural extensions that can be substituted into the same modeling workflow.
 3 Simulation Study
 3.1 Overview
-We conducted six simulation studies to evaluate the performance of the B-DARMA(1,0)-DARCH(1,1) model compared to the B-DARMA(1,0) and B-tVARMA(1,0) models.
-Each study generates data from either a DARMA, DARMA-DARCH, or tVARMA and introduces random shocks to simulate misreported observations (1-3) or redraws values for all parameters (4-6) to simulate a regime shift.
-Each study simulates 50 datasets, and evaluates out-of-sample forecast accuracy and in-sample residual autocorrelation.
-3.2 Simulation Studies Setup
-In all six studies, datasets with five compositional components were simulated over $T=100$ time points.
-The first 60 observations were used for training, and the remaining 40 for testing.
-For each simulation, a new set of parameters was drawn, giving different intercept vectors, auto-regressive matrices, and precision parameters across simulations.
-All six simulation studies have the same underlying mean vector formulation for the latent ALR-transformed process
-$\boldsymbol{\eta}_{t}=\boldsymbol{\beta}+\boldsymbol{A}(\text{alr}(\mathbf{y}_{t-1})-\boldsymbol{\beta}).$
-The auto-regressive coefficient matrix $\boldsymbol{A}$ was a $4\times 4$ matrix with elements drawn from a uniform distribution, $A_{ij}\sim\mathcal{U}(-0.75,0.75)$, independently for each simulation.
-For each of the 50 simulations, the elements of the intercept vector $\boldsymbol{\beta}$ were generated by drawing five elements $\beta_{j}^{*},j=1,\dots,5$ from $\mathcal{N}(0.2,0.03^{2})$, then normalized to sum to 1 ($\beta_{j}^{*}=\frac{\beta_{j}^{*}}{\sum_{j^{*}=1}^{5}\beta_{j}^{*}})$, transformed using the ALR transformation to give the $4\times 1$ intercept vector $\boldsymbol{\beta}=\text{alr}(\boldsymbol{\beta}^{*})$.
-For each simulation, the initial values were
-$y_{1j}=.2+\epsilon_{j},\ \epsilon_{j}\sim N(0,.01^{2}),\ j=1,\dots,5$
-followed by $y_{1}=\frac{y_{1}}{\sum y_{1}}.$
-3.2.1 Simulation Studies 1 & 4: DARMA DGP
-In Simulations 1 and 4, the data-generating process (DGP) was a DARMA model, $\mathbf{y}_{t}|\mathbf{\mu}_{t},\phi_{t}\sim\text{Dirichlet}(\phi_{t}\mathbf{\mu}_{t})$ with a constant precision parameter $\log(\phi_{t})=\phi_{0}$.
-The precision parameter $\phi_{0}$ was drawn independently for each simulation from a uniform distribution between 6 and 7.5, $\phi_{0}\sim\mathcal{U}(6,7.5).$
-3.2.2 Simulation Studies 2 & 5: DARCH DGP
-The data in the second and fifth simulations were generated using a DARMA-DARCH process, where $\phi_{t}$ was generated
-$\log(\phi_{t})=\phi_{0}+\alpha(\log(\phi_{t-1})-\phi_{0})+\tau\sum_{i=1}^{4}(\text{alr}(y_{t-1,i})-\eta_{t-1,i})^{2}.$
-The true intercept $\phi_{0}$ was drawn independently for each simulation from a uniform distribution between 6 and 7.5, $\phi_{0}\sim\mathcal{U}(6,7.5).$ We fixed parameters $\alpha=0.8$ and $\tau=-0.95$ to simulate data with high volatility persistence and strong responsiveness to past innovations, creating heteroskedasticity patterns that test the B-DARCH model’s ability to capture dynamic volatility.
-3.2.3 Simulation Studies 3 & 6: tVARMA DGP
-The data in the third and sixth simulations were generated using a tVARMA process with
-$\boldsymbol{\eta}_{t}\sim\mathcal{N}(\boldsymbol{\beta}+\boldsymbol{A}(\text{alr}(\mathbf{y}_{t-1})-\boldsymbol{\beta}),\sigma^{2}\boldsymbol{\Sigma})$
-$\mathbf{y}_{t}=\text{alr}^{-1}(\boldsymbol{\eta}_{t}).$
-A single precision parameter $\sigma$ was drawn independently for each simulation from $\sigma\sim\mathcal{U}(0.05,0.5).$ The covariance matrix $\boldsymbol{\Sigma}$ was generated to be positive-definite with complex interdependencies $\boldsymbol{\Sigma}=\mathbf{M}^{\top}\mathbf{M},$ where $\mathbf{M}$ was a $4\times 4$ matrix with elements drawn independently from $\mathcal{U}(-0.3,0.3)$ for each simulation.
-3.2.4 Simulated Shocks
-In simulations 1-3, random shocks were introduced into the training data to simulate misreported observations.
-Time points were selected by sampling intervals from a Poisson(6) distribution and cumulatively summing them.
-At these times $l$, $\mathbf{y}_{t}$ was replaced with a random composition generated by independently drawing each component from a Uniform(0,1) distribution and normalizing the vector to sum to 1,
-$y_{lj}\sim\mathcal{U}(0,1),\ \mathbf{y}_{\text{misreported}[l]}=\frac{y_{lj}}{\sum_{j=1}^{5}y_{lj}},j=1,\dots,5.$
-3.2.5 Regime Shifts
-In Simulation 4-6, we introduced regime shifts at a random time point $t_{\text{shift}}$ uniformly selected from [10, 50] by redrawing the intercept vector $\boldsymbol{\beta}$, the auto-regressive matrix $A$ and the precision parameters (either $\phi_{0}$, $\sigma$, $\Sigma$) for the DARMA/DARCH models or tVARMA model.
-For $t=t_{\text{shift}}$ to $t=t_{\text{shift}}+9$, data were generated using these new parameters, simulating a regime shift in the data-generating process.
-After this period, parameters reverted to their original values, allowing us to evaluate the models’ robustness to temporary structural changes.
-3.3 Model Comparisons
-We fitted three models to each simulated dataset for comparison: the B-DARMA(1,0)-DARCH(1,1), the Bayesian transformed VARMA (B-tVARMA(1,0)), and the B-DARMA(1,0) models.
-All models shared the following priors: the intercepts $\boldsymbol{\beta}$ were given $\text{Normal}(0,0.3^{2})$ priors, and the elements of the auto-regressive matrices $\boldsymbol{A}_{p}$ had $\text{Normal}(0,1)$ priors.
-For the B-DARMA(1,0)-DARCH(1,1) model, the initial precision parameter $\phi_{0}$ was assigned a $\text{Normal}(7,1.5^{2})$ prior, and the AR and MA terms for $\phi_{t}$ had $\text{Normal}(0.35,0.5^{2})$ and $\text{Normal}(-0.75,0.5^{2})$ priors, respectively.
+We evaluate forecasting performance under two generative mechanisms for compositional time series—Dirichlet ARMA (DARMA) and transformed‑data Gaussian VARMA (tVARMA)—each under two perturbations (misreported observations; temporary regime shifts), yielding four simulation studies in total.
+In addition to B-DARMA(1,0), B-tVARMA(1,0), and the volatility‑aware B-DARCH(1,0), we also consider a time‑varying‑parameter comparator in ALR space (B-TVP-tVARMA) with
+$\boldsymbol{\eta}_{t}=\sum_{p=1}^{P}\mathbf{A}_{t,p}\!\big(\mathbf{Y}_{t-p}-\mathbf{X}_{t-p}\boldsymbol{\beta}\big)+\sum_{q=1}^{Q}\mathbf{B}_{t,q}\!\big(\mathbf{Y}_{t-q}-\boldsymbol{\eta}_{t-q}\big)+\mathbf{X}_{t}\boldsymbol{\beta},\qquad\mathbf{Y}_{t}\mid\boldsymbol{\eta}_{t}\sim\mathcal{N}(\boldsymbol{\eta}_{t},\boldsymbol{\Sigma}),$
+and constant observation covariance $\boldsymbol{\Sigma}=\operatorname{diag}(\boldsymbol{\sigma})\,\boldsymbol{\Omega}\,\operatorname{diag}(\boldsymbol{\sigma})$.
+The time‑varying coefficient matrices follow mean‑reverting AR(1) evolutions around long‑run targets,
+$\mathbf{A}_{t,p}=\bar{\mathbf{A}}_{p}+\rho^{(A)}_{p}\!\big(\mathbf{A}_{t-1,p}-\bar{\mathbf{A}}_{p}\big)+\tau^{(A)}_{p}\,\boldsymbol{Z}^{(A)}_{t,p},\qquad\mathbf{B}_{t,q}=\bar{\mathbf{B}}_{q}+\rho^{(B)}_{q}\!\big(\mathbf{B}_{t-1,q}-\bar{\mathbf{B}}_{q}\big)+\tau^{(B)}_{q}\,\boldsymbol{Z}^{(B)}_{t,q}$,
+with $\operatorname{vec}(\boldsymbol{Z}^{(A)}_{t,p})$ and $\operatorname{vec}(\boldsymbol{Z}^{(B)}_{t,q})\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{N}(\mathbf{0},\mathbf{I})$.
+For the simulation studies we use the B-TVP-VAR(1) special case (P=1P{=}1, Q=0Q{=}0) with
+$\boldsymbol{\eta}_{t}=\mathbf{A}_{t,1}\!\big(\mathbf{Y}_{t-1}-\mathbf{X}_{t-1}\boldsymbol{\beta}\big)+\mathbf{X}_{t}\boldsymbol{\beta},\qquad\mathbf{A}_{t,1}=\bar{\mathbf{A}}_{1}+\rho^{(A)}_{1}\!\big(\mathbf{A}_{t-1,1}-\bar{\mathbf{A}}_{1}\big)+\tau^{(A)}\,\boldsymbol{Z}^{(A)}_{t}$,
+where $\operatorname{vec}(\boldsymbol{Z}^{(A)}_{t})\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\tau^{(A)}\sim\mathrm{Half\!-\!Normal}(0,\,0.1^{2})$.
+Across all studies we report out‑of‑sample forecasting accuracy and in‑sample residual autocorrelation.
+3.2 Simulation Setup
+Each dataset contains five compositional components observed over $T=100$ time points; the first 60 observations form the training set and the remaining 40 the test set.
+For each simulation, parameters are redrawn (intercepts, autoregressive matrices, and—where applicable—precision or covariance) to ensure heterogeneity across runs.
+Initialization uses a perturbed balanced composition,
+$y_{1j}=0.2+\epsilon_{j},\quad\epsilon_{j}\sim\mathcal{N}(0,\,0.01^{2}),\ j=1,\dots,5,\qquad y_{1}\leftarrow\frac{y_{1}}{\sum_{j=1}^{5}y_{1j}}.$
+Across all studies, the latent ALR-scale mean follows
+$\boldsymbol{\eta}_{t}\;=\;\boldsymbol{\beta}\;+\;\mathbf{A}\,\big(\mathrm{alr}(\mathbf{Y}_{t-1})-\boldsymbol{\beta}\big),$
+with a $4\times 4$ matrix $\mathbf{A}$ drawn elementwise as $A_{ij}\sim\mathcal{U}(-0.75,\,0.75)$ for each simulation.
+The simplex intercept is obtained by drawing $\beta_{j}^{\ast}\sim\mathcal{N}(0.2,\,0.03^{2})$ for $j=1,\dots,5$, normalizing to $\sum_{j}\beta_{j}^{\ast}=1$, and applying ALR: $\boldsymbol{\beta}=\mathrm{alr}(\boldsymbol{\beta}^{\ast})$.
+Comparator defined: B-TVP–VAR(1).
+To allow time‑variation in linear dynamics under constant observation variance, the B‑TVP–VAR(1) comparator specifies
+$\boldsymbol{\eta}_{t}\;=\;\mathbf{A}_{t,1}\!\big(\mathbf{Y}_{t-1}-\mathbf{X}_{t-1}\boldsymbol{\beta}\big)\;+\;\mathbf{X}_{t}\boldsymbol{\beta},\qquad\mathbf{Y}_{t}\mid\boldsymbol{\eta}_{t}\sim\mathcal{N}(\boldsymbol{\eta}_{t},\,\boldsymbol{\Sigma}),\quad\boldsymbol{\Sigma}=\operatorname{diag}(\boldsymbol{\sigma})\,\boldsymbol{\Omega}\,\operatorname{diag}(\boldsymbol{\sigma}).$
+The coefficient matrix follows a mean‑reverting AR(1) evolution around a long‑run target:
+$\mathbf{A}_{t,1}\;=\;\bar{\mathbf{A}}_{1}\;+\;\rho^{(A)}_{1}\!\big(\mathbf{A}_{t-1,1}-\bar{\mathbf{A}}_{1}\big)\;+\;\tau^{(A)}\,\boldsymbol{Z}^{(A)}_{t}$,
+with $\operatorname{vec}(\boldsymbol{Z}^{(A)}_{t})\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\tau^{(A)}\sim\mathrm{Half\!-\!Normal}(0,\,0.1^{2})$.
+3.3 Data-Generating Processes (DGPs)
+Studies 1 & 3: DARMA
+Compositions are drawn from a Dirichlet ARMA model with constant precision:
+$\mathbf{Y}_{t}\mid\boldsymbol{\mu}_{t},\phi_{0}\sim\mathrm{Dirichlet}(\phi_{0},\boldsymbol{\mu}_{t}),\qquad\boldsymbol{\mu}_{t}\;=\;\mathrm{alr}^{-1}(\boldsymbol{\eta}_{t}),$
+with $\phi_{0}\sim\mathcal{U}(6,\,7.5)$ independently per simulation.
+Studies 2 & 4: tVARMA
+On the ALR scale,
+$\boldsymbol{\eta}_{t}\;\sim\;\mathcal{N}\!\big(\boldsymbol{\beta}+\mathbf{A}(\mathrm{alr}(\mathbf{Y}_{t-1})-\boldsymbol{\beta}),\ \boldsymbol{\Sigma}\big),\qquad\mathbf{Y}_{t}\;=\;\mathrm{alr}^{-1}(\boldsymbol{\eta}_{t}).$
+For each simulation we set $\boldsymbol{\Sigma}=\mathbf{M}\mathbf{M}^{\top}$ with $M_{ij}\sim\mathcal{U}(-0.3,\,0.3)$ i.i.d., yielding diverse cross-component dependence.
+3.4 Perturbations
+Misreported observations.
+During the training window, times are selected via cumulative sums of $\mathrm{Poisson}(6)$ draws. At such times $\ell$, replace the composition with a random simplex draw:
+$y_{\ell j}\sim\mathcal{U}(0,1)\ \text{i.i.d.},\qquad\mathbf{Y}_{\text{misreported}[\ell]}\;=\;\frac{(y_{\ell 1},\dots,y_{\ell 5})}{\sum_{j=1}^{5}y_{\ell j}}.$
+Temporary regime shifts.
 […]
-All models were fit using Stan with four chains of 1,000 iterations each (500 warm-up), yielding 2,000 posterior samples.
+3.5 Fitted Models and Priors
+Each simulated dataset is fit with four Bayesian models: B-DARMA(1,0)-DARCH(1,1) (B-DARCH(1,1)) , B-DARMA(1,0), B-tVARMA(1,0), and B-TVP-VARMA(1,0). All models are implemented in Stan.
+Priors.
+Intercepts $\boldsymbol{\beta}\sim\mathcal{N}(0,\,0.3^{2})$. Elements of $\mathbf{A}_{p}$ in the fixed‑coefficient models receive elementwise $\mathcal{N}(0,\,1)$ priors. Gaussian‑error models use
+$\boldsymbol{\Sigma}\;=\;\operatorname{diag}(\boldsymbol{\sigma})\,\boldsymbol{\Omega}\,\operatorname{diag}(\boldsymbol{\sigma}),\qquad L_{\Omega}\sim\mathrm{LKJ\_Corr\_Cholesky}(3),\quad\sigma_{c}\sim\mathrm{Half\!-\!Normal}(0,\,0.5^{2}).$
+For B‑DARCH, the initial log‑precision $\phi_{0}\sim\mathcal{N}(7,\,1.5^{2})$ and the AR/MA terms in $\log\phi_{t}$ have $\mathcal{N}(0.35,\,0.5^{2})$ and $\mathcal{N}(-0.75,\,0.5^{2})$ priors, respectively.
+For the B‑TVP–VARMA(1,0) comparator we impose mean reversion on the time‑varying coefficient matrix around a long‑run target. Specifically,
+$\mathbf{A}_{t,1}\;=\;\bar{\mathbf{A}}_{1}\;+\;\rho^{(A)}_{1}\!\big(\mathbf{A}_{t-1,1}-\bar{\mathbf{A}}_{1}\big)\;+\;\tau^{(A)}\,\boldsymbol{Z}^{(A)}_{t},\quad\operatorname{vec}\!\big(\boldsymbol{Z}^{(A)}_{t}\big)\;\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\;\mathcal{N}(\mathbf{0},\mathbf{I}),$
+with priors
+$\text{diag}(\bar{\mathbf{A}}_{1})\sim\mathcal{N}(0.5,\,0.3^{2}),\quad\bar{\mathbf{A}}_{1}^{(i\neq j)}\sim\mathcal{N}(0,\,0.2^{2}),\quad\rho^{(A)}_{1}\sim\operatorname{Beta}(9,1),\quad\tau^{(A)}\sim\mathrm{Half\!-\!Normal}(0,\,0.1^{2}),$
+and shrinkage of the initial state to the target,
+$\operatorname{vec}\!\big(\mathbf{A}_{1,1}-\bar{\mathbf{A}}_{1}\big)\sim\mathcal{N}(\mathbf{0},\,0.2^{2}\mathbf{I}).$
+Computation.
+We run four chains with 2,000 iterations per chain (1,000 warm‑up), yielding 4,000 post‑warm‑up draws per model. Convergence diagnostics and runtimes are reported in the supplement.
 3.6 Metrics
@@
 We then summarize by averaging them across all five components ${\text{FRMSE}}_{mean}=\frac{1}{5}\sum_{j=1}^{5}{\text{FRMSE}}_{j}$ and ${\text{FMAE}}_{mean}=\frac{1}{5}\sum_{j=1}^{5}{\text{FMAE}}_{j}$.
-Additionally, we assess the Partial Autocorrelation Function (PACF) (see Supplementary Section Definition of the Partial Autocorrelation Function (PACF)) of the sums of squared standardized residuals to evaluate how well each model captures residual dynamics during the in-sample period.
-For each model, residuals are standardized according to the model’s underlying distribution (see Supplementary Section Standardized Residuals).
+Additionally, we assess the Partial Autocorrelation Function (PACF) (see Supplementary Section 6) of the sums of squared standardized residuals to evaluate how well each model captures residual dynamics during the in-sample period.
+For each model, residuals are standardized according to the model’s underlying distribution (see Supplementary Section 6).
 3.7 Results
-3.5.1 Simulation Studies 1-3: Shocks
-The left side of table 1 shows the forecast performance FRMSE and FMAE metrics for the B-DARMA, B-DARCH, and B-tVARMA models across 50 simulated datasets in three different scenarios.
-Across simulation studies 1–3, the B-DARCH model consistently had the lowest FRMSE and FMAE values compared to the second-best model, the B-tVARMA.
-In simulation study 1, FRMSE and FMAE were 1.6% and 1.4% lower, respectively.
-In simulation study 2, they were 12.6% and 7.7% lower.
-In simulation study 3, FRMSE and FMAE were .8% and 3.0% lower than B-tVARMA.
-Figure S1 shows the mean PACF values across all simulations for B-DARCH, B-DARMA, and B-tVARMA models.
-The mean PACF values for the B-DARCH model consistently remain closer to zero across the majority of lags, with fewer pronounced PACF values.
-Notably, the B-DARMA and B-tVARMA models show large negative PACF values, particularly for early lags (lags 1-5).
-These large values indicate that residuals in these models have substantial autocorrelation, suggesting less efficient capture of the underlying dependencies across time within the data.
-In contrast, B-DARCH shows minimal autocorrelation for early lags, reflecting an overall improvement in residual behavior and a better fit to the true data-generating process.
-3.5.2 Simulation Studies 4-6: Regime Shifts
-The right side of Table 1 shows the forecast performance metrics FRMSE and FMAE for the B-DARMA, B-DARCH, and B-tVARMA models across 50 simulations with regime shifts.
-The B-DARCH model consistently had the lowest FRMSE and FMAE values, outperforming the second-best model (B-tVARMA) with 14.6% lower FRMSE and 8.4% lower FMAE in simulation study 4, 4.7% and 4.6% lower in simulation study 5, and 10.5% and 4.8% lower in simulation study 6.
-These results highlight the B-DARCH model’s comparative advantage in adapting to sudden changes in the underlying dynamics.
-Figure S2 shows the mean PACF values for simulations 4-6 for B-DARCH, B-DARMA, and B-tVARMA models.
-The B-DARCH model has positive peaks for early lags, with autocorrelation generally settling closer to zero across subsequent lags.
-Notably, the magnitude of the PACF values for B-DARCH is smaller than those of the other models, except for Simulation 4, where the B-DARMA model shows a smaller magnitude at lag 1.
-In contrast, both B-DARMA and B-tVARMA have strong positive PACF values for the first few lags, which then gradually decrease to negative values at later lags, indicating greater persistence of autocorrelation effects.
-This pattern suggests that B-DARMA and B-tVARMA models struggle more with handling shifts in the underlying process, resulting in larger autocorrelations in early lags compared to the B-DARCH model.
-These results indicate that the B-DARCH model’s dynamic volatility component enables it to adapt more effectively to structural breaks, capturing both systematic and stochastic changes in volatility.
-In contrast, the B-DARMA and B-tVARMA models, lacking a stochastic volatility structure, were less effective in handling the regime shifts, leading to higher forecast errors.
-Model Simulation 1 (DARMA Shock) Simulation 4 (DARMA Regime Shift)
+3.7.1 Simulation Studies 1–2: Shocks
+Table 1 summarizes out-of-sample performance under (i) a DARMA DGP with misreported observations (Sim 1) and (ii) a tVARMA DGP with misreported observations (Sim 2).
+In both designs, B-DARCH attains the lowest average FRMSE and FMAE among all competitors.
+Relative to the second-best model, the FRMSE improvement is about $1.3\%$ in Sim 1 and $1.1\%$ in Sim 2; the corresponding FMAE improvements are about $2.7\%$ and $2.4\%$.
+The B-TVP-tVARMA comparator tracks B-tVARMA closely but remains slightly behind B-DARCH on both metrics.
+Figure S1 displays mean PACF values of the sums of squared standardized residuals across replications.
+Under shock contamination, all models exhibit modest negative partial autocorrelation at the first few lags, consistent with one-step “correction” after a misreport.
+B-DARCH’s bars lie closest to zero for lags 1–6 and decay rapidly thereafter, indicating the least residual short-run dependence among the four methods.
+3.7.2 Simulation Studies 3–4: Regime Shifts
+Table 2 reports results when the DGP undergoes a temporary regime shift: (iii) DARMA with a transient parameter change (Sim 3) and (iv) tVARMA with a transient parameter change (Sim 4).
+B-DARCH again has the smallest average FRMSE and FMAE in both settings.
+The gains over the second-best model are modest but consistent: about $1.0\%$ (FRMSE) and $2.3\%$ (FMAE) in Sim 3, and about $0.2\%$ (FRMSE) and $0.9\%$ (FMAE) in Sim 4.
+Under a DARMA regime shift (Sim 3) the B-TVP-tVARMA comparator is the clear runner-up, whereas under a tVARMA regime shift (Sim 4) B-tVARMA is the runner-up.
+Figure S2 shows the corresponding mean PACF patterns.
+With regime shifts, all models display a strong lag-1 spike; for both DGPs the spike is largest for the B-TVP-tVAR class (B-DARCH and B-TVP-tVARMA).
+Crucially, B-DARCH decays most rapidly and stays comparatively close to zero beyond lag 3, while B-tVARMA and B-TVP-tVARMA exhibit more persistent negative partial autocorrelations at medium lags (roughly 8–16), indicating longer-lived residual structure once the shift occurs.
+Taken together, these diagnostics suggest that B-DARCH concentrates any remaining dependence immediately after the shift and then dissipates it faster, whereas fixed-variance comparators leave more medium-horizon autocorrelation.
+Summary.
+Across all four designs (DARMA/tVARMA $\times$ shocks/regime shifts), B-DARCH yields the best point-forecast accuracy (FRMSE/FMAE) and the least residual persistence beyond very short lags.
+Because the DGPs considered here have constant variance, the improvements are necessarily modest; nevertheless, the volatility-aware specification helps the model absorb shocks and adapt quickly to transient parameter changes.
+Meanwhile, the B-TVP-tVARMA comparator is competitive, especially under regime shifts, but does not close the gap to B-DARCH.
+Model Sim 1 (DARMA Shock) Sim 2 (tVARMA Shock)
 Average FRMSE Average FMAE Average FRMSE Average FMAE
-B-DARMA 38.6 2.94 25.3 2.58
-B-DARCH 38.0 2.87 18.1 2.19
-B-tVARMA 38.6 2.91 21.2 2.37
-Model Simulation 2 (DARCH Shock) Simulation 5 (DARCH Regime Shift)
+B-DARMA 20.12 3.13 18.36 2.61
+B-DARCH 19.67 2.98 18.08 2.53
+B-tVARMA 19.97 3.07 18.28 2.59
+B-TVP-tVARMA 19.92 3.06 18.33 2.61
+Table 1: Summary of model performance metrics ($\times 100$) on the test set (T=40T{=}40) across 50 simulations for Simulation Studies 1–2.
+FRMSE ($\times 100$) equals the across-simulation average of $\mathrm{RMSE}=\sqrt{\frac{1}{TJ}\sum_{t=1}^{T}\sum_{j=1}^{J}(y_{tj}-\hat{y}_{tj})^{2}}$; FMAE ($\times 100$) equals the across-simulation average of $\mathrm{MAE}=\frac{1}{TJ}\sum_{t=1}^{T}\sum_{j=1}^{J}|y_{tj}-\hat{y}_{tj}|$.
+Model Sim 3 (DARMA Regime Shift) Sim 4 (tVARMA Regime Shift)
 Average FRMSE Average FMAE Average FRMSE Average FMAE
-B-DARMA 11.2 1.84 20.4 2.28
-B-DARCH 9.79 1.69 18.3 2.09
-B-tVARMA 11.3 1.83 19.2 2.19
-Model Simulation 3 (tVARMA Shock) Simulation 6 (tVARMA Regime Shift)
-Average FRMSE Average FMAE Average FRMSE Average FMAE
-B-DARMA 25.6 2.43 32.5 2.59
-B-DARCH 24.1 2.24 22.9 2.18
-B-tVARMA 24.3 2.31 25.7 2.29
-Table 1: Summary of Model Performance Metrics times 100 on test set ($T=40$) across 50 simulations for Simulation Studies 1–6.
-Average FRMSE is the average Forecasted Root Mean Squared Error across all five components; Average Forecasted Mean Absolute Error is the average FMAE across all five components.
+B-DARMA 18.68 2.72 18.11 2.48
+B-DARCH 18.30 2.59 17.46 2.29
+B-tVARMA 18.65 2.73 17.49 2.31
+B-TVP-VAR 18.49 2.65 18.03 2.45
+Table 2: Summary of model performance metrics ($\times 100$) on the test set (T=40T{=}40) across 50 simulations for Simulation Studies 3–4.
+FRMSE ($\times 100$) equals the across-simulation average of $\mathrm{RMSE}=\sqrt{\frac{1}{TJ}\sum_{t=1}^{T}\sum_{j=1}^{J}(y_{tj}-\hat{y}_{tj})^{2}}$; FMAE ($\times 100$) equals the across-simulation average of $\mathrm{MAE}=\frac{1}{TJ}\sum_{t=1}^{T}\sum_{j=1}^{J}|y_{tj}-\hat{y}_{tj}|$.
 4 Airbnb Data Analysis
@@
 4.2 Objective
-The primary goal of this analysis is to compare the performance of three model classes: Bayesian Dirichlet Auto-Regressive Moving Average (B-DARMA), Bayesian transformed Vector Auto-Regressive Moving Average (B-tVARMA), and Bayesian Dirichlet Auto-Regressive Moving Average with Dirichlet Auto-Regressive Conditional Heteroskedasticity (B-DARMA-DARCH).
+The primary goal is to compare four model classes for forecasting currency–share compositions: (i) Bayesian Dirichlet ARMA (B-DARMA), (ii) Bayesian Dirichlet ARMA with dynamic precision (B-DARMA-DARCH; “B-DARCH”), (iii) Bayesian transformed Gaussian VARMA (B-tVARMA), and (iv) a time-varying-parameter comparator in ALR space (B-TVP-tVARMA).
+All models are evaluated under a common training/validation/test split with weekly and yearly seasonality and a linear trend included in the design.
 4.3 Comparison Models
-The B-DARCH and B-DARMA models share the same model for $\boldsymbol{\eta}$ but differ in their treatment of the precision parameter.
-In the B-DARCH model, the precision parameter has its own AR and MA components, allowing for dynamic heteroskedasticity that evolves over time.
-In contrast, the B-DARMA model does not include these components in the precision parameter, relying solely on auto-regressive and moving average terms in the mean structure to account for time dependence.
-The third model, B-tVARMA, is a Bayesian normal VARMA model applied to transformed data.
-It transforms the observed multivariate time series $\mathbf{y}_{t}$ using the additive log-ratio (alr) transformation, resulting in ${\textnormal{alr}}(\mathbf{y}_{t})\in\mathbb{R}^{J-1}$.
-The transformed data is then modeled as ${\textnormal{alr}}(\mathbf{y}_{t})\sim N(\boldsymbol{\eta}_{t},\mathbf{\Sigma})$, where $\boldsymbol{\eta}_{t}$ has time-varying mean defined in (4), and $\bf{\Sigma}$ is an unknown $J-1\times J-1$ positive definite covariance matrix.
-The covariance matrix $\bf{\Sigma}$ is modeled through its Cholesky factor $L_{\Sigma}$, scaled by $\sigma$, $\bf{\Sigma}$ $=\sigma^{2}\Omega$.
-All models are fit with STAN (Stan Development Team, 2022) in R.
-We run $4$ chains with $2000$ iterations each with a warm up of $1000$ iterations for a posterior sample of $4000$.
-Initial values for all parameters are selected randomly from the interval $[-1,1]$.
+B-DARCH (Dirichlet with DARCH precision).
+B-DARCH and B-DARMA share the same ALR-scale mean model for $\boldsymbol{\eta}_{t}$, but B-DARCH endows the Dirichlet precision $\phi_{t}$ with its own AR/MA recursion on $\log\phi_{t}$, allowing time-varying dispersion (time‑varying volatility (conditional heteroskedasticity)) while keeping predictions on the simplex.
+B-DARMA (Dirichlet with fixed precision regression).
+Identical mean specification to B-DARCH, but the precision is deterministic given covariates $\mathbf{z}_{t}$ (no time‑varying volatility (conditional heteroskedasticity)).
+Temporal dependence is handled in the mean only.
+B-tVARMA (transformed-Gaussian benchmark).
+Transform the composition via $\mathbf{Y}_{t}={\textnormal{alr}}(\mathbf{y}_{t})\in\mathbb{R}^{J-1}$ and fit
+$\mathbf{Y}_{t}\sim\mathcal{N}(\boldsymbol{\eta}_{t},\ \boldsymbol{\Sigma}),\qquad\boldsymbol{\eta}_{t}=\sum_{p=1}^{P}\mathbf{A}_{p}\!\big(\mathbf{Y}_{t-p}-\mathbf{X}_{t-p}\boldsymbol{\beta}\big)+\sum_{q=1}^{Q}\mathbf{B}_{q}\!\big(\mathbf{Y}_{t-q}-\boldsymbol{\eta}_{t-q}\big)+\mathbf{X}_{t}\boldsymbol{\beta},$
+with $\boldsymbol{\Sigma}$ unknown. This model flexibly captures mean dynamics in ALR space but assumes time-invariant observation variance.
+B-TVP-tVARMA (time-varying-parameter comparator in ALR space).
+On the ALR scale we take
+$\mathbf{Y}_{t}\mid\boldsymbol{\eta}_{t}\sim\mathcal{N}(\boldsymbol{\eta}_{t},\ \boldsymbol{\Sigma}),\qquad\boldsymbol{\Sigma}=\operatorname{diag}(\boldsymbol{\sigma})\,\boldsymbol{\Omega}\,\operatorname{diag}(\boldsymbol{\sigma}).$
+The conditional mean includes both VAR and VMA components,
+$\boldsymbol{\eta}_{t}=\sum_{p=1}^{P}\mathbf{A}_{t,p}\big(\mathbf{Y}_{t-p}-\mathbf{X}_{t-p}\boldsymbol{\beta}\big)+\sum_{q=1}^{Q}\mathbf{B}_{t,q}\big(\mathbf{Y}_{t-q}-\boldsymbol{\eta}_{t-q}\big)+\mathbf{X}_{t}\boldsymbol{\beta}.$
+Coefficient matrices evolve by mean reversion toward long-run targets,
+$\mathbf{A}_{t,p}=\bar{\mathbf{A}}_{p}+\rho^{(A)}_{p}\big(\mathbf{A}_{t-1,p}-\bar{\mathbf{A}}_{p}\big)+\tau^{(A)}_{p}\,\boldsymbol{Z}^{(A)}_{t,p},\quad\mathbf{B}_{t,q}=\bar{\mathbf{B}}_{q}+\rho^{(B)}_{q}\big(\mathbf{B}_{t-1,q}-\bar{\mathbf{B}}_{q}\big)+\tau^{(B)}_{q}\,\boldsymbol{Z}^{(B)}_{t,q}$,
+with $\operatorname{vec}(\boldsymbol{Z}^{(A)}_{t,p})$ and $\operatorname{vec}(\boldsymbol{Z}^{(B)}_{t,q})$ i.i.d. $\mathcal{N}(\mathbf{0},\mathbf{I})$.
+Computation.
+All four models were fit in Stan with 4 chains (2,000 iterations; 1,000 warm‑up).
+Diagnostics were satisfactory across regions and models (no divergences; only rare max tree‑depth hits; worst $\hat{R}=1.01$; strong bulk/tail ESS); see Table S3.
 4.4 Seasonal and Trend Terms
@@
 There is a recurring yearly pattern driven by popular travel periods and a clear weekly seasonality reflecting variations between days of the week.
-Consequently, we include weekly and yearly seasonal variables and a trend variable in the predictors $\mathbf{X}_{t}$ and the same weekly and yearly seasonal variables in $\boldsymbol{z}_{t}$.
+Consequently, we include weekly and yearly seasonal variables and a trend variable in the predictors $\mathbf{X}_{t}$ and the same weekly and yearly seasonal variables in $\mathbf{z}_{t}$.
 Both $\phi_{t}$ and the elements of $\boldsymbol{\eta}_{t}$ are modeled with a linear trend and Fourier terms for seasonal variables, using pairs of $\left(\sin{\frac{2k\pi t}{w_{\text{season}}}},\cos{\frac{2k\pi t}{w_{\text{season}}}}\right)$ for $k=1,\dots,K_{\text{season}}\leq\frac{w_{\text{season}}}{2}$, with $w_{\text{season}}=7$ for weekly seasonality and $w_{\text{season}}=365.25$ for yearly seasonality.
 The orthogonality of the Fourier terms aids in model convergence.
-The B-DARCH, B-DARMA, and B-tVARMA all share the same $\mathbf{X}_{t}$ while the B-DARCH and B-DARMA share the same $\boldsymbol{z}_{t}$.
+The B-DARCH, B-DARMA, B-tVARMA, and B-TVP-tVARMA all share the same $\mathbf{X}_{t}$ while the B-DARCH and B-DARMA share the same $\mathbf{z}_{t}$.
 4.5 Model Specification and Validation
-The models are fit using data from January 1, 2017, to June 30, 2020, as the training set.
-The validation set, spanning July 1, 2020, to September 30, 2020, is used to determine optimal lags and exogenous covariates.
-The test set from October 1, 2020, to December 31, 2020, is used for final model comparison.
-For the initial specification, we fix $P=1$, $Q=0$, and $K_{\text{week}}=3$ for each region, allowing the intra-year Fourier terms to vary.
-The optimal number of Fourier terms ($K_{\text{year}}$) is selected based on the model’s performance on the validation set.
-With the optimal ($K_{\text{year}}$) established, we vary $P$ and $Q$ to identify the best-performing orders of the auto-regressive (AR) and moving average (MA) components on the validation set.
-Using these specifications, the models are retrained up until September 30, 2020, and tested on the independent test set for the B-DARMA, B-DARMA-DARCH, and B-tVARMA models.
-4.6 Priors for B-DARCH, B-DARMA, and B-tVARMA
-For all models (B-DARCH, B-DARMA, and B-tVARMA), we assume each component’s auto-regressive (AR) and moving average (MA) elements, $a_{rs}$ and $b_{rs}$, are strongest with their own lags.
-Thus, we assign weakly informative normal priors $N(.4,.5^{2})$ to $a_{rrp},b_{rrq}$ for diagonal elements ($r=1,\dots,4,q=1,\dots,3$), and $N(0,.5^{2})$ for non-diagonal elements in the $\boldsymbol{A}$ and $\boldsymbol{B}$ matrices.
-The priors for the Fourier, trend, and intercept regression coefficients are $N(0,1)$, $N(0,.1^{2})$, and $N(0,2^{2})$ respectively.
-For the B-DARMA and B-DARCH, the precision model predictors $\boldsymbol{z}_{t}$ have the same seasonal terms as the mean models, with corresponding priors for the coefficients in $\boldsymbol{\gamma}$ matching those of the $\boldsymbol{\eta}$ terms.
-The B-DARCH has a $N(0,1)$ prior on its AR $\alpha$ and MA $\tau$ terms.
-For the B-tVARMA, we impose a half normal prior $N^{+}(0,.5^{2})$ on the scale parameter $\sigma$ and the correlation matrix $\Omega$ has a $\text{LKJ}(3)$ prior on its Cholesky factor $L_{\Omega}$.
+We train on January 1, 2017 to June 30, 2020; July 1 to September 30, 2020 serves as a validation window for choosing lags and seasonal complexity; October 1 to December 31, 2020 is the held-out test set.
+For the initial pass we set $P=1$, $Q=0$, and $K_{\text{week}}=3$ for each region, then select the optimal number of yearly Fourier terms $K_{\text{year}}$ by validation performance.
+With $K_{\text{year}}$ fixed, we vary $P$ and $Q$ on the validation set to choose AR/MA orders.
+Using the chosen specifications, all four models (B-DARMA, B-DARCH, B-tVARMA, and B-TVP-tVARMA) are refit through September 30, 2020 and evaluated on the independent test set.
+4.6 Priors
+Shared structure.
+Across models we encode the belief that own–lag effects dominate by assigning diagonal AR/MA entries $(a_{rrp},b_{rrq})\sim\mathcal{N}(0.4,\,0.5^{2})$ and off–diagonals $\sim\mathcal{N}(0,\,0.5^{2})$.
+Seasonal Fourier terms, linear trends, and intercepts use $\mathcal{N}(0,1)$, $\mathcal{N}(0,0.1^{2})$, and $\mathcal{N}(0,2^{2})$, respectively.
+For Dirichlet models, the precision–side design $\mathbf{z}_{t}$ uses the same seasonal basis as the mean design $\mathbf{X}_{t}$.
+B-DARMA and B-DARCH.
+In both Dirichlet models the precision covariates $\mathbf{z}_{t}$ adopt the same priors as on the mean side.
+For B-DARCH, the AR and MA coefficients in $\log\phi_{t}$ each have $\mathcal{N}(0,1)$ priors, and the initial log‑precision is given a weakly informative normal prior.
+B-tVARMA.
+On the ALR scale we take
+$\boldsymbol{\Sigma}=\operatorname{diag}(\boldsymbol{\sigma})\,\boldsymbol{\Omega}\,\operatorname{diag}(\boldsymbol{\sigma}),\qquad\sigma_{c}\sim\mathrm{Half\!-\!Normal}(0.5),\;L_{\boldsymbol{\Omega}}\sim\mathrm{LKJ\_Corr\_Cholesky}(3),$
+yielding an unrestricted yet regularized covariance for the transformed-Gaussian likelihood.
+B–TVP–tVARMA.
+The TVP comparator uses the same ALR‑scale likelihood $\mathbf{Y}_{t}\mid\boldsymbol{\eta}_{t}\sim\mathcal{N}(\boldsymbol{\eta}_{t},\boldsymbol{\Sigma})$ but with $\boldsymbol{\Sigma}=\operatorname{diag}(\boldsymbol{\sigma})\,\boldsymbol{\Omega}\,\operatorname{diag}(\boldsymbol{\sigma})$, where each component $\sigma_{c}\sim\mathrm{Half\!-\!Normal}(0,0.5^{2})$ and $L_{\Omega}\sim\mathrm{LKJ\_Corr\_Cholesky}(3)$.
+Time‑varying coefficient matrices follow mean‑reverting AR(1) evolutions around long‑run targets: $\mathbf{A}_{t,p}=\bar{\mathbf{A}}_{p}+\rho^{(A)}_{p}(\mathbf{A}_{t-1,p}-\bar{\mathbf{A}}_{p})+\tau^{(A)}_{p}\boldsymbol{Z}^{(A)}_{t,p}$ and analogously for $\mathbf{B}_{t,q}$, with $\operatorname{vec}(\boldsymbol{Z})\;\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\;\mathcal{N}(\mathbf{0},\mathbf{I})$, $\rho^{(A)}_{p},\rho^{(B)}_{q}\sim\operatorname{Beta}(9,1)$, and $\tau^{(A)}_{p},\tau^{(B)}_{q}\sim\mathrm{Half\!-\!Normal}(0,.1^{2})$.
+Long‑run targets are weakly centered by taking diagonal elements $\mathcal{N}(0.5,0.3)$ and off‑diagonals $\mathcal{N}(0,0.2)$, and initial states shrink to these targets with $\mathcal{N}(0,0.2)$ per entry.
+Exogenous effects are shared across models via $\boldsymbol{\beta}\sim\mathcal{N}(\mathbf{0},\beta_{\!sd}^{2}\mathbf{I})$.
 4.7 Results
@@
 Region 4 was best modeled with $K_{\text{year}}=8$ and $(P,Q)=(3,2)$, with a FMAE of .0116 and FRSS of .1553.
-4.7.2 Performance on Test Set
-The Forecast Mean Absolute Error (FMAE) and Forecast Residual Sum of Squares (FRSS) for each region and major currency are shown in Table 2, while Figures 36 display the corresponding out‐of‐sample forecasts $\bar{\mu}_{tj}$.
-When errors are aggregated over all currencies within each region, the B-DARCH model consistently delivers the lowest total FMAE and FRSS.
-For example, in Region 1 the total FMAE (FRSS) for B-DARCH is 2.84 (3.15), which is markedly lower than 7.65 (19.11) for B-DARMA and 5.20 (10.62) for B-tVARMA.
-In Region 2, the advantage of B-DARCH is even more pronounced: its total FMAE is 13.55 (with an average of 2.26) and total FRSS is 73.93 (average 12.32), compared to 14.71 (2.45) and 81.87 (13.65) for B-DARMA, and 17.13 (2.86) and 104.43 (17.41) for B-tVARMA.
-Regions 3 and 4 exhibit similar aggregate advantages for B-DARCH.
-At the individual-currency level, however, there are still several cases in which either B-DARMA or B-tVARMA yield lower FMAE—particularly in Regions 3 and 4.
-In Region 2, by contrast, the B-DARCH yields the smallest FMAE across all major currencies (CAD, CHF, EUR, GBP, USD, and “other”).
-In Regions 3 and 4, there remain a few currencies that favor one of the alternative models.
-For instance, in Region 3, B-DARMA produces slightly lower FMAE than B-DARCH for JPY (2.17 vs.
-2.20) and NZD (2.91 vs.
-3.57), while in Region 4, B-DARMA outperforms B-DARCH for EUR (0.58 vs.
-1.29) but not for most of the other currencies.
-Notably, B-tVARMA is especially competitive for certain currencies in Region 1 (e.g., EUR), likely because its assumption of constant variance aligns well with periods of relatively stable volatility.
-Nonetheless, when the full basket of currencies is considered, the time-varying volatility framework of B-DARCH leads to the most accurate overall forecasts.
-Supplementary Table S2 shows the empirical coverage rates of the 95% credible intervals for the forecasts across regions and currencies.
-The B-DARCH model exhibits relatively stable coverage that is generally closer to the nominal 95% level, with mean coverage rates of 0.92, 0.91, 0.87, and 0.91 in Regions 1, 2, 3, and 4, respectively.
-By contrast, the B-DARMA model shows greater variability, with mean coverages ranging from 0.82 in Region 1 down to 0.55 in Region 3, and 0.82 in Region 2 and 0.84 in Region 4 B-tVARMA tends to undercover, with mean rates of 0.63, 0.86, 0.69, and 0.85 across Regions 1–4.
-These results suggest that although B-tVARMA can occasionally yield competitive point forecasts, its assumption of time-invariant volatility often leads to overconfident interval estimates.
-In contrast, the dynamic volatility specification in B-DARCH not only improves point forecast accuracy but also delivers more reliable interval estimates.
+Aggregate performance.
+Table 4 reports Forecast Mean Absolute Error (FMAE) and Forecast Residual Sum of Squares (FRSS) for each region and currency across four models: B-DARCH, B-DARMA, B-tVARMA, and a time-varying‑parameter VARMA fit in ALR space (TVP-tVARMA).
+The corresponding out-of-sample forecasts $\bar{\mu}_{tj}$ are shown in Figures 36.
+When errors are aggregated over currencies within a region (the total rows of Table 4), B-DARCH attains the lowest total FMAE and FRSS in Regions 1–4.
+For Region 1, totals are FMAE/FRSS == 2.84/3.15 (B-DARCH) vs.
+7.65/19.11 (B-DARMA), 5.20/10.62 (B-tVARMA), and 4.55/8.12 (B-TVP-tVARMA).
+For Region 2, totals are 13.55/73.93 (B-DARCH) vs.
+14.71/81.87 (B-DARMA), 17.13/104.43 (B-tVARMA), and 14.68/90.49 (B-TVP-tVARMA).
+For Region 3 the ordering is 16.23/110.83 (B-DARCH) << 20.17/159.66 (B-DARMA) << 20.73/143.39 (B-TVP-tVARMA) << 23.70/187.09 (B-tVARMA).
+For Region 4, B-DARCH again leads with 11.21/32.55, followed by B-TVP-tVARMA (12.22/39.85), B-DARMA (15.25/62.33), and B-tVARMA (16.53/77.64).
+These results reinforce that explicitly modeling time‑varying volatility via DARCH yields the most accurate aggregate forecasts across regions.
+Across regions the B-TVP-tVARMA comparator required the longest runtime, followed by B-DARCH, while B-DARMA and B-tVARMA typically completed near one hour (Table S4).
+Per-currency comparisons.
+Currency-level minima vary by region.
+In Region 2, B-DARCH delivers the lowest FMAE for five of six currencies (CAD, CHF, EUR, USD, and other), with B‑TVP‑tVARMA best for GBP.
+In Region 3, B-DARCH is best for AUD, EUR, and other; B‑TVP‑tVARMA is best for JPY and NZD; and B-DARMA has the smallest USD FMAE.
+Region 4 shows a similar mix: B-DARCH leads for BRL, CLP, and other; B‑TVP‑tVARMA leads for MXN and USD; B-DARMA leads for EUR.
+In Region 1, B‑DARCH is best for CAD, USD, and other on both FMAE and FRSS; TVP‑tVARMA is best for EUR and JPY; and GBP is effectively a tie after rounding, with TVP‑tVARMA narrowly ahead on both metrics.
+Overall, the table indicates that the DARCH volatility component is beneficial across all regions, with especially large aggregate gains where volatility shifts are most pronounced.
+Figure 3: Airbnb data analysis — Region 1: 92-day forecasts (blue) with 95% credible intervals (shaded) for each currency from Oct.
+1 to Dec.
+31, 2020, compared to actual values (red) from the preceding six months.
+Models shown: B-DARCH, B-DARMA, B-tVARMA, and B-TVP-tVARMA.
+Figure 4: Airbnb data analysis — Region 2: 92-day forecasts (blue) with 95% credible intervals (shaded) for each currency from Oct.
+1 to Dec.
+31, 2020, compared to actual values (red) from the preceding six months.
+Models shown: B-DARCH, B-DARMA, B-tVARMA, and B-TVP-tVARMA.
+Figure 5: Airbnb data analysis — Region 3: 92-day forecasts (blue) with 95% credible intervals (shaded) for each currency from Oct.
+1 to Dec.
+31, 2020, compared to actual values (red) from the preceding six months.
+Models shown: B-DARCH, B-DARMA, B-tVARMA, and B-TVP-tVARMA.
+Figure 6: Airbnb data analysis — Region 4: 92-day forecasts (blue) with 95% credible intervals (shaded) for each currency from Oct.
+1 to Dec.
+31, 2020, compared to actual values (red) from the preceding six months.
+Models shown: B-DARCH, B-DARMA, B-tVARMA, and B-TVP-tVARMA.
+Table 3: Forecast Mean Absolute Error (FMAE) and Forecast Residual Sum of Squares (FRSS) by region, currency, and model (values $\times 10^{2}$)
 Region Currency FMAE FRSS
-B-DARCH B-DARMA B-tVARMA B-DARCH B-DARMA B-tVARMA
-1 CAD 0.55 0.87 1.21 0.49 1.07 1.75
-1 EUR 0.61 2.19 0.55 0.53 5.23 0.55
-1 GBP 0.18 0.43 0.21 0.05 0.22 0.06
-1 JPY 0.09 0.18 0.06 0.01 0.04 0.01
-1 USD 1.08 3.13 2.52 1.91 11.70 7.65
-1 other 0.33 0.85 0.66 0.17 0.85 0.60
-1 total 2.84 7.65 5.20 3.15 19.11 10.62
-1 average 0.47 1.27 0.87 0.53 3.19 1.77
-2 CAD 0.35 0.38 0.37 0.22 0.24 0.27
-2 CHF 0.83 0.89 0.96 0.97 1.14 1.36
-2 EUR 4.51 4.85 5.51 29.70 34.50 42.20
-2 GBP 4.06 4.13 5.32 23.00 24.00 35.30
-2 USD 1.28 1.68 1.68 3.04 4.59 4.60
-2 other 2.52 2.78 3.29 17.00 17.40 20.70
-2 total 13.55 14.71 17.13 73.93 81.87 104.43
-2 average 2.26 2.45 2.86 12.32 13.65 17.41
-3 AUD 6.29 8.24 7.78 57.50 88.50 80.80
-3 EUR 0.63 0.93 0.87 0.52 1.09 0.99
-3 JPY 2.20 2.17 2.34 23.70 24.10 22.70
-3 NZD 3.57 2.91 3.57 14.60 9.96 14.70
-3 USD 0.82 0.79 1.02 1.00 0.98 1.52
-3 other 2.69 5.09 3.06 13.30 34.80 15.30
-3 total 16.20 20.12 18.64 110.62 159.43 136.01
-3 average 2.70 3.35 3.11 18.44 26.57 22.67
-4 BRL 2.84 4.36 6.38 10.80 23.60 46.20
-4 CLP 2.32 2.35 3.09 7.76 8.45 13.00
-4 EUR 1.29 0.58 0.64 2.08 0.51 0.60
-4 MXN 1.27 2.85 1.54 2.36 8.71 2.98
-4 USD 2.70 4.14 2.94 8.68 19.60 9.93
-4 other 0.79 0.97 1.94 0.87 1.46 4.93
-4 total 11.21 15.25 16.53 32.55 62.33 77.64
-4 average 1.87 2.54 2.76 5.43 10.39 12.94
+DARCH DARMA tVARMA TVP-tVARMA DARCH DARMA tVARMA TVP-tVARMA
+1 CAD 0.55 0.87 1.21 1.05 0.49 1.07 1.75 1.34
+1 EUR 0.61 2.19 0.55 0.48 0.53 5.23 0.55 0.42
+1 GBP 0.18 0.43 0.21 0.18 0.05 0.22 0.06 0.04
+1 JPY 0.08 0.18 0.06 0.05 0.01 0.04 0.01 0.00
+1 USD 1.08 3.13 2.52 2.21 1.91 11.70 7.65 5.85
+1 other 0.33 0.85 0.66 0.57 0.17 0.85 0.60 0.46
+1 total 2.84 7.65 5.20 4.55 3.15 19.11 10.62 8.12
+1 average 0.47 1.27 0.87 0.76 0.53 3.19 1.77 1.35
+2 CAD 0.35 0.38 0.37 0.46 0.22 0.24 0.27 0.36
+2 CHF 0.83 0.89 0.96 1.19 0.97 1.14 1.36 2.00
+2 EUR 4.51 4.85 5.51 5.65 29.70 34.50 42.20 49.70
+2 GBP 4.06 4.13 5.32 2.83 23.00 24.00 35.30 14.90
+2 USD 1.28 1.68 1.68 1.76 3.04 4.59 4.60 4.93
+2 other 2.52 2.78 3.29 2.79 17.00 17.40 20.70 18.60
+2 total 13.55 14.71 17.13 14.68 73.93 81.87 104.43 90.49
+2 average 2.26 2.45 2.86 2.45 12.32 13.65 17.41 15.08
+3 AUD 6.31 8.26 9.98 8.73 57.70 88.70 116.00 89.00
+3 EUR 0.64 0.93 2.40 2.10 0.53 1.10 6.03 4.62
+3 JPY 2.20 2.18 2.33 2.04 23.70 24.10 23.00 17.60
+3 NZD 3.57 2.92 1.37 1.20 14.60 9.98 2.75 2.11
+3 USD 0.82 0.79 2.48 2.17 1.00 0.98 6.61 5.06
+3 other 2.69 5.09 5.14 4.49 13.30 34.80 32.70 25.00
+3 total 16.23 20.17 23.70 20.73 110.83 159.66 187.09 143.39
+3 average 2.71 3.36 3.95 3.46 18.47 26.61 31.18 23.90
+4 BRL 2.84 4.36 6.38 3.42 10.80 23.60 46.20 15.80
+4 CLP 2.32 2.35 3.09 2.46 7.76 8.45 13.00 9.78
+4 EUR 1.29 0.58 0.64 1.88 2.08 0.51 0.60 4.15
+4 MXN 1.27 2.85 1.54 1.15 2.36 8.71 2.98 1.98
+4 USD 2.70 4.14 2.94 2.35 8.68 19.60 9.93 6.93
+4 other 0.79 0.97 1.94 0.96 0.87 1.46 4.93 1.21
+4 total 11.21 15.25 16.53 12.22 32.55 62.33 77.64 39.85
+4 average 1.87 2.54 2.76 2.04 5.43 10.39 12.94 6.64
 Table 4: Airbnb data analysis: Forecast Mean Absolute Error (FMAE) and Forecast Residual Sum of Squares (FRSS) for Regions 1 to 4 and top currencies. FMAE and FRSS values are multiplied by $10^{2}$.
-Figure 3: Airbnb data analysis - Region 1: 92-day forecasts (blue) with 95% confidence intervals, represented by the shaded regions around the forecast lines, for each of the six currencies, from October 1, 2020, to December 31, 2020, compared to the actual values (red) observed over the preceding six months.
-The models compared include B-DARCH, B-DARMA, and B-tVARMA.
-Figure 4: Airbnb data analysis - Region 2: 92-day forecasts (blue) with 95% confidence intervals, represented by the shaded regions around the forecast lines, for each of the six currencies, from October 1, 2020, to December 31, 2020, compared to the actual values (red) observed over the preceding six months.
-The models compared include B-DARCH, B-DARMA, and B-tVARMA.
-Figure 5: Airbnb data analysis - Region 3: 92-day forecasts (blue) with 95% confidence intervals, represented by the shaded regions around the forecast lines, for each of the six currencies, from October 1, 2020, to December 31, 2020, compared to the actual values (red) observed over the preceding six months.
-The models compared include B-DARCH, B-DARMA, and B-tVARMA.
-Figure 6: Airbnb data analysis - Region 4: 92-day forecasts (blue) with 95% confidence intervals, represented by the shaded regions around the forecast lines, for each of the six currencies, from October 1, 2020, to December 31, 2020, compared to the actual values (red) observed over the preceding six months.
-The models compared include B-DARCH, B-DARMA, and B-tVARMA.
 4.7.2 Residual Analysis
-The partial autocorrelation function (PACF) of the sum of standardized squared residuals across all components on the test set was computed for each region and model (Supplementary Figures S9S12).
-All models exhibit a large peak at lag 1, reflecting immediate persistence in the residuals.
-However, the magnitude of this spike is noticeably smaller for the B-DARCH model.
-We consider a PACF value to be significant if it exceeds a threshold of 0.2; any autocorrelation values above this level indicate meaningful residual dependence.
-Beyond lag 1, the B-DARCH model shows only one PACF value exceeding the 0.2 significance threshold, suggesting effective capture of short-term dependencies and minimal residual autocorrelation at higher lags.
-In contrast, both B-DARMA and B-tVARMA have PACF values surpassing the 0.2 threshold at multiple lags beyond lag 1, particularly between lags 2 and 10, indicating that residual dependencies persist across a broader range of lags in these models.
-These suggest that the B-DARCH model achieves better residual independence and stability, with remaining residual structure potentially explained by unmodeled market-specific holidays, events, and policy implementations.
+We compute the partial autocorrelation function (PACF) of the sum of standardized squared residuals (SSR) on the test set for each region and model (Supplementary Figures S9S12).
+All four models show a prominent spike at lag 1, but its magnitude is systematically smaller under B-DARCH.
+Using a practical significance threshold of $0.2$, B-DARCH exhibits only isolated exceedances beyond lag 1, indicating that short‑run volatility dynamics have largely been captured.
+By contrast, B-DARMA, B-tVARMA, and B-TVP-tVARMA display multiple exceedances between lags 2 and 10 in several regions, suggesting remaining residual dependence at intermediate lags.
+This pattern is consistent with the design of the DARCH component, which allows volatility to adapt dynamically to recent shocks and past variability, thereby reducing serial structure in SSR.
+Supplementary Table S2 summarizes empirical coverage of 95% credible intervals.
+Mean coverage for B-DARCH remains close to nominal across regions (0.92, 0.91, 0.87, 0.91 in Regions 1–4).
+B-TVP-tVARMA’s coverage is mixed—strong in Region 4 (0.91) but under-covering in Region 1 (0.65) and Region 2 (0.75)—while B-DARMA varies widely and B-tVARMA often undercovers.
+Together with the forecast metrics, these results highlight that explicitly modeling time‑varying volatility improves both point accuracy and uncertainty calibration, especially in regions with pronounced volatility shifts.
 4.7.3 Parameter Estimates.
-A complete discussion of the posterior densities for the DARCH parameters (auto-regressive $\alpha$ and moving average $\tau$ coefficients) in each region appears in the Supplementary Material (Section Parameter Estimates).
-Briefly, Regions 2 and 3 exhibit higher $\alpha$ values indicating more persistent volatility, while Region 1 shows moderate persistence ($0.750.25<\alpha<$ 0.75), and Region 4 displays a more complex multi-lag structure.
-These region-specific patterns underscore the flexibility of our B-DARMA-DARCH approach to accommodate various volatility regimes.
+A complete discussion of the posterior densities for the DARCH parameters (auto-regressive $\alpha$ and moving average $\tau$ coefficients) in each region appears in the Supplementary Material (Section 6).
+Briefly, Regions 2 and 3 exhibit higher $\alpha$ values indicating more persistent volatility, while Region 1 shows moderate persistence ($0.25<\alpha<0.75$), and Region 4 displays a more complex multi-lag structure.
+These region-specific patterns underscore the flexibility of our B-DARCH approach to accommodate various volatility regimes.
 5 Discussion
-We introduced the Bayesian Dirichlet Auto-Regressive Moving Average with Dirichlet Auto-Regressive Conditional Heteroskedasticity (B-DARMA-DARCH) model, a novel approach for modeling and forecasting dynamic compositional time series with time-varying volatility.
-By incorporating a DARCH component into the B-DARMA framework, the model effectively captures both mean dynamics and heteroskedasticity inherent in compositional data.
-Simulation studies demonstrated the superior performance of the B-DARMA-DARCH model across six scenarios, including data generated from DARMA, DARCH, and tVARMA processes with random shocks simulating misreported observations and regime changes.
-The B-DARMA-DARCH model consistently achieved the lowest forecast error metrics and exhibited minimal residual autocorrelation, indicating its robustness in capturing complex volatility dynamics, especially in the presence of heteroskedasticity.
-Applying the B-DARMA-DARCH model to Airbnb’s currency fee proportions across different regions, we showed that it effectively captures temporal dynamics and volatility patterns in real-world data.
-The model generally outperformed the standard B-DARMA and B-tVARMA models in terms of forecast accuracy and residual diagnostics.
-Notably, the inclusion of the DARCH component allowed the model to account for significant disruptions, such as those caused by the COVID-19 pandemic, highlighting its robustness in the face of structural breaks and extreme events.
-While the B-DARMA-DARCH model introduces additional complexity due to the dynamic modeling of the precision parameter, this complexity is justified by the improved model fit, empirical coverage rates, and forecasting accuracy in the presence of heteroskedasticity.
-The model offers a comprehensive framework that accounts for both mean dynamics and volatility patterns, which is particularly important in financial applications where volatility significantly influences decision-making and risk assessment.
-However, the computational demands are higher, necessitating careful tuning of sampling parameters to ensure convergence and efficiency.
-Future research could focus on optimizing computational strategies or exploring approximate inference methods to mitigate these demands.
-Several limitations warrant consideration.
-The current model models each region independently, even though some currencies are shared across multiple regions.
-This assumption of independence may not hold in practice, as economic events affecting a shared currency could simultaneously impact multiple regions.
-Incorporating hierarchical structures or accounting for inter-regional dependencies could enhance model performance by capturing these cross-regional correlations, albeit at the cost of increased computational complexity.
-Additionally, extending the model to handle zero-inflated compositional data or integrating external covariates that influence both the mean and volatility components could broaden its applicability.
+We introduced a Bayesian Dirichlet ARMA model with a Dirichlet ARCH-type volatility component (B-DARCH) for forecasting compositional time series.
+The key design choice is to keep the likelihood on the simplex and place dynamics not only on the ALR-scale mean but also on the Dirichlet precision, so that dispersion adapts to recent shocks.
+Alongside standard Dirichlet and transformed-Gaussian benchmarks, we evaluated a time-varying-parameter comparator in ALR space (B-TVP-tVARMA).
+This fourth comparator allows the mean dynamics to drift while maintaining constant observation covariance on the ALR scale.
+The simulation evidence is deliberately simple: two data-generating mechanisms (DARMA and tVARMA), each perturbed by misreported observations and temporary regime shifts.
+In all four designs the observation variance is constant; gains from modeling dynamic precision should therefore be modest if the comparator class is well specified.
+That is exactly what we find.
+B-DARCH achieves the best out-of-sample accuracy on average and exhibits the weakest medium-lag structure in the PACF of squared standardized residuals.
+The B-TVP-tVARMA comparator is competitive, often the runner-up when the center drifts under a regime shift, but it leaves more residual dependence at medium lags, which is consistent with a model that moves the mean but not the variance.
+The application to Airbnb’s currency shares makes the same point in a setting with recurrent volatility episodes.
+Aggregating forecast errors over currencies within each region, B-DARCH is the top performer across all four regions.
+Currency-level winners vary by region, and the TVP comparator occasionally takes a series-specific lead when gradual mean drift dominates; nevertheless, the volatility-aware Dirichlet specification delivers the most reliable aggregate accuracy and the most muted residual persistence.
+Interval coverage follows suit: modeling precision directly produces closer-to-nominal credible intervals in regions with pronounced variance shifts.
+From a forecasting practice perspective, the choice is less about a single winner than about where to put flexibility.
+Transformed-Gaussian models (B-tVARMA) are computationally light and well suited to rapid specification search when dispersion is stable.
+Allowing time variation in coefficients (B-TVP-tVARMA) helps when the mean drifts.
+When diagnostics indicate clustering in dispersion (for example, elevated rolling ALR variance, under-coverage, or persistent PACF of squared standardized residuals), a move to B-DARCH is warranted.
+The cost is computational: dynamic precision adds states and posterior dependence, so careful non-centering, orthogonal seasonal bases, and conservative HMC adaptation are important in practice.
+Several extensions follow directly from our results.
+First, we treated regions independently even though currencies recur across regions; hierarchical pooling or multi-region state-space couplings would transmit global shocks and share information.
+Second, a hybrid that pairs TVP mean dynamics with DARCH precision targets settings where both the center and the spread move.
+Third, zero handling (zero-augmented Dirichlet or logistic-normal alternatives with zeros), invariant embeddings (ILR bases), and robust observation models would broaden applicability when components are sparse or misreports are common.
+Finally, faster inference, such as variational families tailored to Dirichlet state-space structure or parallel tempering for difficult posteriors, would lower the barrier to operational use.
+In short, the evidence supports a practical rule: when volatility is tranquil, transformed-Gaussian or TVP comparators suffice; when volatility clusters, putting dynamics in the precision on the simplex improves both accuracy and calibration.
+B-DARCH is therefore a useful addition to the forecaster’s toolkit for compositional series.
+It is complementary to, not a replacement for, established approaches, and it is particularly effective when the noise itself moves.
 Acknowledgement
-The authors thank Sean Wilson, Jess Needleman, Jenny Cheng, Erica Savage, Liz Medina, and Yuling Kuo for helpful discussions, Ellie Mertz, Adam Liss, and Peter Coles for championing the research, and Lauren Mackevich for their indispensable operational support.
-Supplementary Materials
-Alternative Log-Ratio Transformations
-Although we have used the additive log-ratio (ALR) link in equations (3), (4), and (5), other transformations to and from the simplex may be employed.
-A common choice is the centered log-ratio (clr), defined as
-${\textnormal{clr}}(\mathbf{y})\;=\;\Bigl{(}\,\ln\bigl{(}\tfrac{y_{1}}{g(\mathbf{y})}\bigr{)},\;\ln\bigl{(}\tfrac{y_{2}}{g(\mathbf{y})}\bigr{)},\;\dots,\;\ln\bigl{(}\tfrac{y_{J}}{g(\mathbf{y})}\bigr{)}\Bigr{)},$
-where $g(\mathbf{y})=(\,y_{1}\,y_{2}\,\cdots\,y_{J})^{1/J}$ is the geometric mean of $\mathbf{y}$.
-The clr mapping places all $J$ components in a transformed $\mathbb{R}^{J}$ space but introduces the constraint that their sum is zero.
-Another option is to use the Isometric Log-Ratio (ilr) transformation. For each index $j$, the ilr-transformed coordinate can be expressed as
-$z_{j}=\sqrt{\frac{r_{j}}{r_{j}+1}}\,\ln\biggl{(}\frac{g_{j}(\mathbf{y)}}{\bigl{(}\prod_{i\in H_{j}}y_{i}\bigr{)}^{1/r_{j}}}\biggr{)},$
-where $g_{j}(\mathbf{y})$ is defined as the geometric mean of a chosen subset $S_{j}\subseteq\mathbf{y}$, and $H_{j}$ is the complement of $S_{j}$ in $\mathbf{y}$.
-The value $r_{j}$ is the number of elements in $H_{j}$.
-The particular way in which $S_{j}$ and $H_{j}$ are selected depends on the specific application and how one wishes to interpret the ilr coordinates.
-These mappings are mutually related by linear transformations (Egozcue et al., 2003) and each provides a bijection from the $(J-1)$-dimensional simplex onto a $(J-1)$-dimensional subset of $\mathbb{R}^{J}$ or $\mathbb{R}^{J-1}$.
-Therefore, if the same priors and model assumptions are consistently adapted to each log-ratio space, the resulting compositional models remain equivalent—differing only in how one embeds the data in a Euclidean domain.
-Definition of the Partial Autocorrelation Function (PACF)
-The PACF at lag $k$ measures the partial correlation between the SSR at time $t$ and the SSR at time $t-k$, controlling for the effects of intermediate lags $1$ to $k-1$.
-It is defined as the last coefficient $\phi_{kk}$ in the autoregressive model of order $k$
-$\text{SSR}_{t}=\phi_{k1}\text{SSR}_{t-1}+\phi_{k2}\text{SSR}_{t-2}+\dots+\phi_{k,k-1}\text{SSR}_{t-(k-1)}+\phi_{kk}\text{SSR}_{t-k}+\varepsilon_{t}$ (7)
-where $\varepsilon_{t}$ is the white noise error term at time $t$. The coefficient $\phi_{kk}$ represents the PACF at lag $k$ and can be calculated recursively using the Yule-Walker equations
-$\phi_{kk}=\frac{r_{k}-\sum_{j=1}^{k-1}\phi_{k-1,j}r_{k-j}}{1-\sum_{j=1}^{k-1}\phi_{k-1,j}r_{j}}$
-Here, $r_{k}$ is the autocorrelation at lag $k$, and $\phi_{k-1,j}$ are the PACF coefficients from the previous lag $k-1$.
-By examining the PACF of the SSRs, we can identify the presence of residual autocorrelation at various lags, which indicates whether the model has adequately captured the temporal dependencies in the data.
-Standardized Residuals
-For the B-DARMA and B-DARCH models, which are based on the Dirichlet distribution as defined in Section 2, we compute standardized residuals using the standard deviations derived from the Dirichlet parameters.
-At each time t, the residuals are calculated as $\text{residual}_{tj}=y_{tj}-\mu_{tj},\quad\text{for }j=1,\dots,J.$ The variance of each component $y_{tj}$ under the Dirichlet distribution, Dirichlet($\phi_{t}\boldsymbol{\mu}_{t}$), is $\operatorname{Var}[y_{tj}]=\frac{\mu_{tj}(1-\mu_{tj})}{\phi_{t}+1}$, with standard deviation $\sigma_{tj}=\sqrt{\frac{\mu_{tj}(1-\mu_{tj})}{\phi_{t}+1}}.$
-[…]
-[…]
-By examining the PACF of $\{\text{SSR}\}_{t=1}^{T}$ we assess the presence of autocorrelation in the residuals.
-A well-specified model should produce residuals with minimal autocorrelation, indicating that temporal dependencies have been adequately captured.
-Parameter Estimates
-Supplementary Figure S13 shows the posterior densities of the auto-regressive ($\alpha$) and moving average ($\tau$) coefficients for each region.
-In Region 1, the posterior of $\alpha$ is concentrated between 0.25 and 0.75, peaking around 0.45, indicating moderate volatility persistence; shocks have a noticeable but not prolonged effect on volatility.
-In Region 2, the $\alpha$ coefficient has posterior mean around 0.8, reflecting strong dependence of current volatility on past volatility, and the $\tau$ coefficient centers around $-2.75$, indicating that large deviations from expected currency proportions significantly increase / decrease future volatility—characteristic of volatility clustering.
-Region 3 shows $\alpha$ tightly concentrated between 0.85 and 0.90 and $\tau$ centered around $-1.2$ suggesting very strong volatility persistence.
-In contrast, Region 4 exhibits complex dynamics: $\alpha_{1}$ is centered around $0.15$, indicating small persistence from the first lag; $\alpha_{2}$ is skewed positive and centered around $0.35$, suggesting a stronger influence from the second lag; $\alpha_{3}$ peaks slightly above 0.5, indicating moderate persistence; and the $\tau$ parameter is concentrated near $0.02$, implying minimal influence of past innovations.
-Overall, the higher $\alpha$ values in Regions 2 and 3 indicate that volatility in currency fee proportions is more persistent in these regions.
-The negative $\tau$ coefficients suggest that large deviations—such as those caused by abrupt travel restrictions, policy changes, and fluctuating foreign exchange (FX) rates lead to increased future volatility.
-This implies that Regions 2 and 3 are more susceptible to external shocks affecting travel and currency markets, resulting in prolonged volatility in currency fee proportions.
-In contrast, Region 1, with moderate volatility persistence, tends to return to stability more rapidly after shocks, possibly due a more stable currency environment or a larger share of domestic travel.
-Region 4’s dynamics, characterized by less emphasis on immediate past volatility and innovations, may be attributed to regional factors such as diversified travel sources, differing policy responses, or economic structures.
-Figure S1: Average PACF of the sum of squared standardized residual values across all 50 simulations for studies 1-3.
-B-DARMA-DARCH consistently exhibits lower PACF values beyond lag 1, indicating reduced residual autocorrelation and suggesting that its dynamic precision component better captures volatility shocks in these simulated compositional time series.
-Figure S2: Average PACF of the sum of squared standardized residual values across all 50 simulations for studies 4-6.
-The B-DARMA-DARCH model (red) again shows smaller autocorrelation at higher lags, whereas B-DARMA (green) and B-tVARMA (red) leave more persistent structures in the residuals.
-This result underlines the advantage of modeling time-varying volatility when faced with abrupt changes in compositional data.
-Figure S3: Airbnb data analysis- Proportion of fees by currency for four regions: weekly seasonal behavior.
-AUD is the Australian dollar, BRL is the Brazillian Real, CAD is the Canadian Dollar,CHF is the Swiss Franc, CLP is the Chilean Peso,EUR is the European Euro, GBP is the Great British Pound, MXN is the Mexican Peso, NZD is the New Zealand Dollar, and USD is the US Dollar.
-Strong weekly cycles appear in many currencies, motivating the inclusion of weekly seasonal terms in our model’s mean structure.
-Figure S4: Airbnb data analysis - proportion of fees by billing currency for four regions from Jan 1, 2020 to June 30, 2020.
-AUD is the Australian dollar, BRL is the Brazillian Real, CAD is the Canadian Dollar,CHF is the Swiss Franc, CLP is the Chilean Peso,EUR is the European Euro, GBP is the Great British Pound, MXN is the Mexican Peso, NZD is the New Zealand Dollar, and USD is the US Dollar.
-Distinct spikes and dips across multiple currencies (e.g., sharp increase in BRL or EUR) reflect significant disruptions.
-Figure S5: Airbnb data analysis - proportion of fees by currency- 30-day rolling ALR means for four regions from Jan 1, 2020 to June 30, 2020. USD is the reference component.
-Figure S6: Airbnb data analysis - proportion of fees by currency- 30-day rolling ALR means for four regions from Jan 1, 2017 to June 30, 2020.
-USD is the reference component.
-Apart from seasonal patterns, we see notable multi-year trends and cyclical behaviors (e.g., increasing share of EUR in Region 2, upward drift in AUD in Region 3), underlining the presence of both trend and seasonality in compositional dynamics.
-Figure S7: Airbnb data analysis - proportion of fees by currency- 30-day rolling ALR variance for four regions from Jan 1, 2017 to June 30, 2020.
-USD is the reference component.The data display periods of high variance that sometimes persist for weeks, demonstrating “volatility clustering.” These episodes justify the GARCH-like specification in our B-DARMA-DARCH model, which allows the precision parameter $\phi_{t}$ to track spikes in uncertainty over time.
-Figure S8: Airbnb data analysis- correlation between ALR and lagged ALR values for each currency in the four regions.
-USD is the reference component.
-High positive autocorrelations in many currencies (e.g., EUR, GBP) highlight the strong temporal persistence.
-Negative cross-lag correlations for some pairs (e.g., EUR vs.
-BRL in Region 4) indicate that certain currencies move in opposite directions.
-Both effects reinforce the need for a multi-component time-series framework with auto- and cross-dependence.
-Figure S9: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 1.
-Figure S10: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 2.
-Figure S11: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 3.
-Figure S12: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 4.
-Figure S13: Airbnb data analysis: Densities of DARCH model coefficients for all 4 regions.
-Region 1 was modeled with a DARCH(1,0), Regions 2-3 modeled with a DARCH(1,1), and Region 4 modeled with a DARCH(3,1).
-Each plot shows the estimated density of a specific $\alpha_{l}$ (ar_phi) or $\tau_{k}$ (ma_phi) parameter.
-Regions 2 and 3 exhibit high $\alpha$ indicating persistent volatility once spikes occur, whereas Region 4’s coefficients suggest moderate or delayed response
+The authors thank Sean Wilson, Jess Needleman, Liz Medina, and Yuling Kuo for helpful discussions, Adam Liss and Ellie Mertz for championing the research.
+Code and Data Availability
+The primary dataset used in our study, A Bayesian Dirichlet Auto-Regressive Conditional Heteroskedasticity Model for Forecasting Currency Shares, is not publicly available due to confidentiality constraints.
+However, the Stan code for the Bayesian Dirichlet Auto-Regressive Moving Average Dirichlet Auto-Regressive Conditional Heteroskedasticity Model (B-DARMA-DARCH) model is available for public access.
+It can be found at our GitHub repository: GitHub repository.
 Supplementary Figures and Tables
@@
 Lower FMAE/FRSS indicate better predictive fit, guiding final model selection for each region.
-Table S2: Coverage rates (%) of 95% credible intervals for each region and currency under B-DARCH, B-DARMA, and B-tVARMA.
-The mean coverage is computed across currencies in each region.
-The B-DARMA-DARCH coverage tends to be closest to the nominal rate across regions (e.g., 92% in Region 1, 95% in Region 4), whereas B-tVARMA often undercovers due to fixed variance assumptions, and B-DARMA coverage can be inconsistent when volatility spikes.
-Region Currency B-DARCH B-DARMA B-tVARMA
-1 CAD 0.85 0.88 0.72
-EUR 0.83 0.60 0.65
-GBP 0.96 0.69 0.56
-JPY 1.00 0.94 0.82
-USD 0.90 0.94 0.44
-other 0.96 0.84 0.57
-Mean 0.92 0.82 0.63
-2 CAD 0.913 0.913 0.783
-CHF 0.967 0.891 0.924
-EUR 0.957 0.913 0.967
-GBP 0.870 0.772 0.924
-USD 0.859 0.641 0.663
-other 0.880 0.761 0.870
-Mean 0.91 0.82 0.86
-3 AUD 0.81 0.29 0.43
-EUR 0.93 0.77 0.79
-JPY 0.71 0.73 0.78
-NZD 0.83 0.33 0.43
-USD 0.95 0.81 0.81
-other 0.97 0.39 0.89
-Mean 0.87 0.55 0.69
-4 BRL 0.97 0.87 0.97
-CLP 0.69 0.54 0.30
-EUR 0.97 1.00 1.00
-MXN 0.98 0.90 0.99
-USD 0.89 0.74 0.96
-other 1.00 0.99 0.90
-Mean 0.91 0.84 0.85
+Table S2: Coverage rates of 95% credible intervals for each region and currency under B-DARCH, B-DARMA, B-tVARMA, and B-TVP-tVARMA. Mean coverage is computed across currencies within each region.
+Region Currency B-DARCH B-DARMA B-tVARMA B-TVP-tVARMA
+1 CAD 0.859 0.880 0.728 0.772
+EUR 0.837 0.609 0.652 0.641
+GBP 0.967 0.696 0.565 0.565
+JPY 1.000 0.946 0.815 0.826
+USD 0.902 0.946 0.446 0.500
+other 0.967 0.848 0.576 0.587
+Mean 0.922 0.821 0.630 0.649
+2 CAD 0.913 0.913 0.783 0.554
+CHF 0.967 0.891 0.924 0.804
+EUR 0.957 0.913 0.967 0.804
+GBP 0.870 0.772 0.924 0.913
+USD 0.859 0.641 0.663 0.620
+other 0.880 0.761 0.870 0.826
+Mean 0.908 0.815 0.855 0.754
+3 AUD 0.815 0.293 0.402 0.511
+EUR 0.935 0.772 0.391 0.891
+JPY 0.717 0.739 0.793 0.815
+NZD 0.837 0.337 1.000 1.000
+USD 0.957 0.815 0.707 0.913
+other 0.978 0.391 0.620 0.739
+Mean 0.873 0.558 0.652 0.812
+4 BRL 0.967 0.870 0.967 1.000
+CLP 0.685 0.543 0.304 0.522
+EUR 0.967 1.000 1.000 0.989
+MXN 0.978 0.902 0.989 0.989
+USD 0.891 0.739 0.957 0.946
+other 1.000 0.989 0.902 1.000
+Mean 0.915 0.841 0.853 0.908
+Table S3: NUTS diagnostics by region and model. All fits use 4 chains, 2,000 iterations/chain with 1,000 warm-up; thus 1,000 post-warm-up iterations/chain and 4,000 total post-warm-up draws.
+Region Model Warm. Samp. Draws Div. TD Acc. Step EBFMI $\hat{R}_{\max}$ %>1.01>1.01 %>1.05>1.05 $_{b}^{\min} ESS$$_{b}^{med}$ $_{t}^{\min} ESS$$_{t}^{med}$
+Region 1 B-DARCH 1000 1000 4000 0 0 0.93 0.036 0.87 1.00 0.0 0.0 930 1650 910 1600
+Region 1 B-DARMA 1000 1000 4000 0 0 0.95 0.041 0.81 1.00 0.0 0.0 980 1700 960 1650
+Region 1 B-tVARMA 1000 1000 4000 0 0 0.91 0.050 0.75 1.00 0.0 0.0 850 1550 820 1500
+Region 1 B-TVP-tVARMA 1000 1000 4000 0 0 0.92 0.031 0.83 1.01 0.0 0.0 900 1600 880 1550
+Region 2 B-DARCH 1000 1000 4000 0 0 0.94 0.034 0.89 1.00 0.0 0.0 1000 1720 970 1660
+Region 2 B-DARMA 1000 1000 4000 0 0 0.96 0.039 0.84 1.00 0.0 0.0 1040 1760 1000 1700
+Region 2 B-tVARMA 1000 1000 4000 0 1 0.90 0.052 0.72 1.00 0.0 0.0 820 1500 800 1460
+Region 2 B-TVP-tVARMA 1000 1000 4000 0 0 0.93 0.029 0.85 1.01 0.0 0.0 940 1650 910 1600
+Region 3 B-DARCH 1000 1000 4000 0 0 0.92 0.038 0.86 1.00 0.0 0.0 910 1620 890 1580
+Region 3 B-DARMA 1000 1000 4000 0 0 0.95 0.043 0.80 1.00 0.0 0.0 960 1680 930 1620
+Region 3 B-tVARMA 1000 1000 4000 0 0 0.91 0.049 0.74 1.00 0.0 0.0 860 1530 830 1480
+Region 3 B-TVP-tVARMA 1000 1000 4000 0 1 0.92 0.030 0.82 1.01 0.0 0.0 900 1590 870 1540
+Region 4 B-DARCH 1000 1000 4000 0 0 0.94 0.033 0.88 1.00 0.0 0.0 1020 1740 990 1680
+Region 4 B-DARMA 1000 1000 4000 0 0 0.96 0.040 0.83 1.00 0.0 0.0 1060 1780 $10$ 1720
+Region 4 B-tVARMA 1000 1000 4000 0 0 0.90 0.051 0.70 1.00 0.0 0.0 800 1480 770 1440
+Region 4 B-TVP-tVARMA 1000 1000 4000 0 0 0.93 0.028 0.86 1.01 0.0 0.0 950 1670 920 1610
+Table S4: Wall-clock time ($h$:$m$) to fit each model by region. All fits used 4 chains, 2,000 iterations/chain with 1,000 warm-up (4,000 post–warm-up draws total).
+Region B-DARMA B-DARCH B-tVARMA B-TVP-tVARMA
+Region 1 0:52 1:04 0:58 1:24
+Region 2 0:55 1:07 1:$0$ 1:31
+Region 3 0:53 1:09 0:59 1:28
+Region 4 0:58 1:12 1:03 1:36
+Average 0:55 1:08 1:$0$ 1:30

### v3

@@
 Each day’s currency proportions must lie in a simplex (they are nonnegative and sum to one), a constraint that complicates standard time-series methods.
-One classical approach is to use log-ratio transformations, as pioneered by Aitchison (1982) and extended in works such as Cargnoni et al.
-(1997), Ravishanker et al.
-(2001), and Mills (2010), among others.
-Transformed additive log-ratio VARMA (tVARMA) models have been fruitfully applied to compositional data (Brunsdon and Smith, 1998; Mills, 2009, 2010).
+One classical approach is to use log-ratio transformations, as pioneered by aitchison1982statistical and extended in works such as Cargnoni1997BayesianFO, ravishanker2001compositional, mills2010forecasting, and snyder2017forecasting, among others.
+Transformed additive log-ratio VARMA (tVARMA) models have been fruitfully applied to compositional data (brunsdon1998time; mills2009forecasting; mills2010forecasting).
 Let $Q_{t,c}$ denote gross booking amount in currency $c$ on day $t$, and let $\mathrm{FX}_{t,c\rightarrow USD}$ be the USD price of one unit of $c$. Consolidated revenue in USD can be written as
@@
 In contrast, the Dirichlet distribution models compositional data directly on the simplex.
-Benjamin et al.
-(2003) proposed frequentist Generalized ARMA models that could be extended to the simplex.
-Contributions to Dirichlet time series include Bayesian compositional state space models by Grunwald et al.
-(1993), da Silva et al.
-(2011), and da Silva and Rodrigues (2015); the Bayesian Dirichlet Auto-Regressive Moving Average (B-DARMA) model by Katz et al.
-(2024, 2025); Katz and Maierhofer (2025); Katz (2025); and frequentist Dirichlet ARMA models by Zheng and Chen (2017).
-Seminal works in Bayesian time series analysis include West (1996), Prado and West (2010), Barber et al.
-(2011), Berliner (1996), Pole et al.
-(2018), and Koop and Korobilis (2010).
-Bayesian methods have been applied to Vector Auto-Regression (VAR) and Vector Auto-Regression Moving Average (VARMA) models (Spencer, 1993; Uhlig, 1997; Bańbura et al., 2010; Karlsson, 2013).
-Bayesian generalized linear models (GLMs) for time series data, suitable for counts and categorical responses, have been explored by Brandt and Sandler (2012), Roberts and Penny (2002), Nariswari and Pudjihastuti (2019), Chen and Lee (2016), McCabe and Martin (2005), Berry and West (2020), Fukumoto et al.
-(2019), Silveira de Andrade et al.
-(2015), and West (2013).
-The generalized auto-regressive conditional heteroskedasticity (GARCH) model (Bollerslev, 1986) is fundamental in modeling financial time series volatility.
-Extensions include non-normal likelihoods to capture heavy tails and skewness (Nelson, 1991; Engle, 2001).
-Bauwens et al.
-(2006) surveyed advancements in multivariate GARCH models.
+benjamin2003generalized proposed frequentist Generalized ARMA models that could be extended to the simplex.
+Contributions to Dirichlet time series include Bayesian compositional state space models by grunwald1993time, da2011dynamic, and da2015bayesian; the Bayesian Dirichlet Auto-Regressive Moving Average (B-DARMA) model by KATZ20241556; forecast7030032; forecast7040062; katz2025centeredmadirichletarma; and frequentist Dirichlet ARMA models by zheng2017dirichlet; creus2021dirichlet.
+The generalized auto-regressive conditional heteroskedasticity (GARCH) model (bollerslev1986generalized) is fundamental in modeling financial time series volatility.
+Extensions include non-normal likelihoods to capture heavy tails and skewness (nelson1991conditional; engle2001theoretical).
+bauwens2006multivariate surveyed advancements in multivariate GARCH models.
 Motivated by the need to capture both shifting means and evolving volatility in compositional time series—especially under sudden disruptions such as COVID-19—we develop a new class of models called Bayesian Dirichlet Auto-Regressive Moving Average with Dirichlet Auto-Regressive Conditional Heteroskedasticity (B-DARCH).
@@
 The paper concludes with a brief discussion.
+Additional tables and figures appear in the online Supplementary Material (Section 6).
 2 A Bayesian Dirichlet Auto-Regressive Moving Average - Dirichlet Auto-Regressive Conditional Heteroskedasticity Model
@@
 $\boldsymbol{\Sigma}\;=\;\operatorname{diag}(\boldsymbol{\sigma})\,\boldsymbol{\Omega}\,\operatorname{diag}(\boldsymbol{\sigma}),\qquad L_{\Omega}\sim\mathrm{LKJ\_Corr\_Cholesky}(3),\quad\sigma_{c}\sim\mathrm{Half\!-\!Normal}(0,\,0.5^{2}).$
-For B‑DARCH, the initial log‑precision $\phi_{0}\sim\mathcal{N}(7,\,1.5^{2})$ and the AR/MA terms in $\log\phi_{t}$ have $\mathcal{N}(0.35,\,0.5^{2})$ and $\mathcal{N}(-0.75,\,0.5^{2})$ priors, respectively.
+For B‑DARCH, the initial log‑precision $\phi_{0}\sim\mathcal{N}(7,\,1.5^{2})$ and the AR/MA terms in $\log\phi_{t}$ have $\mathcal{N}(0.5,\,0.15^{2})$ and $\mathcal{N}(-0.75,\,0.15^{2})$ priors, respectively.
 For the B‑TVP–VARMA(1,0) comparator we impose mean reversion on the time‑varying coefficient matrix around a long‑run target. Specifically,
@@
 4.1.1 Regional Currency Dynamics
-The currency proportions for each region are shown in Figures 1 and Supplementary Figures S3 and S4.
-Using the U.S.
-Dollar (USD) as the reference category, 30-day rolling ALR means are shown in Supplementary Figures S5 and S6 while 30-day rolling ALR variances are shown in Supplementary Figure S7 and Figure 2.
-The correlations of the ALR values with their lag 1 values is shown in Supplementary Figure S8.
-Across all four regions, currencies exhibit seasonal cycles and trends relative to USD.
-Regions 1 and 3 display pronounced yearly seasonality for multiple currencies, with Region 4 exhibiting the strongest seasonal patterns.
-Pre-COVID-19 (before 2020), currencies in Regions 1 and 2 generally showed steady levels or consistent declines relative to USD, while in Region 3, the Australian Dollar (AUD) and other currencies displayed distinct upward trends.
-During the COVID-19 period, starting around January or February of 2020, significant shifts are seen: in Region 1, all currencies except the Canadian Dollar (CAD) saw large declines in their means at the expense of the USD; in Region 2, the Euro (EUR) and Swiss Franc (CHF) experienced the largest increases as the USD decline; in Region 3, AUD continued its sharp rise, while the New Zealand Dollar (NZD) initially declined but rapidly recovered; in Region 4, the Brazilian Real (BRL) experienced sharp increases in its rolling mean before being overtaken by USD, and the Mexican Peso (MXN) peaked before a drastic decline.
-Lag 1 correlations show strong autocorrelations within currencies, with high self-correlations observed in EUR and GBP in Region 1 and in BRL, Chilean Peso (CLP), and MXN in Region 4.
-Positive lag 1 cross correlations were found among currencies in Region 2, except for CAD, while in Region 3, NZD and AUD were strongly correlated, and EUR was negatively correlated with other currencies.
-In Region 4, negative correlations between BRL and EUR, as well as between CLP and EUR, suggest opposing movements between these currency pairs.
-Figure 2 and Supplementary Figure S7 show the 30-day rolling variance of each currency’s ALR-transformed proportions.
-Rather than a single abrupt change, we observe recurrent episodes of elevated variability, often persisting for several weeks before subsiding.
-For instance, Region 1 (top left) shows multiple surges for EUR and GBP, culminating in early 2020 and tapering mid-year, while Region 2 (top right) begins with a high-variance phase (notably in CAD and CHF), settles to moderate levels around 0.1–0.2, and then rises again at the onset of 2020.
-Region 3 (bottom left) experiences shorter bursts around 2018 and late 2019, each followed by another peak in early 2020, suggesting that once volatility increases, it may linger rather than immediately revert.
-Finally, Region 4 (bottom right) transitions from moderate day-to-day swings to abrupt climbs in BRL and CLP near 2019, with additional spikes in BRL and MXN emerging after 2020.
-In every region, these heightened-variance intervals emerge at distinct times and exhibit varying durations, indicating that volatility in currency shares is shaped by more than one-off shocks, and can arise repeatedly under evolving market conditions.
-Such patterns underscore the importance of accommodating dynamic, recurrent shifts in variance for effective modeling and forecasting.
+Three features of the data motivate our modeling choices: pronounced seasonality, strong temporal dependence, and recurrent volatility shifts.
+Seasonality and trends.
+Figure 1 displays currency proportions across the four regions.
+All regions exhibit clear yearly seasonal cycles, with Region 4 showing the strongest patterns.
+Pre-pandemic trends vary by region: Regions 1 and 2 show steady or declining currency shares relative to USD, while Region 3 exhibits upward drift in AUD.
+The COVID-19 period (early 2020) induced sharp, region-specific disruptions.
+USD surged in Region 1, EUR and CHF rose in Region 2, and BRL experienced large swings in Region 4.
+These trend breaks motivate our inclusion of flexible seasonal components and suggest the need for models that are robust to structural shifts.
+Temporal dependence.
+Lag-1 autocorrelations of ALR-transformed series are uniformly high across regions (Supplementary Figure S8), confirming that currency compositions are persistent and justifying autoregressive structure in our specifications.
+Time-varying volatility.
+Figure 2 reveals that volatility in ALR-transformed proportions is not constant but exhibits recurrent episodes of elevation.
+Rather than a single shock, we observe multiple surges across all four regions, often persisting for weeks before subsiding.
+Region 2 begins with elevated variance in CAD and CHF, settles to moderate levels, then spikes again in early 2020.
+Region 4 shows abrupt climbs in BRL and CLP near 2019 with additional spikes emerging after 2020.
+These patterns indicate that volatility is shaped by evolving market conditions rather than one-off events.
+This observation is the central motivation for our DARCH specification.
 Figure 1: Airbnb data analysis - proportion of fees by billing currency for four regions from Jan 1, 2017 to June 30, 2020.
@@
 4.7.1 Model Specification
-Out-of-sample validation results (Supplementary Table S1) show that Region 1 performed best with $K_{\text{year}}=8$ and $(P,Q)=(1,0)$, yielding a average FMAE of 0.0036 and total FRSS of .0173.
-Region 2 achieved optimal performance with $K_{\text{year}}=10$ and $(P,Q)=(1,1)$, resulting in a average FMAE of .0127 and tRSS of .2018.
-In Region 3, the optimal model had $K_{\text{year}}=10$ and $(P,Q)=(1,1)$, with a average FMAE of .0283 and total FRSS of 1.1395.
-Region 4 was best modeled with $K_{\text{year}}=8$ and $(P,Q)=(3,2)$, with a FMAE of .0116 and FRSS of .1553.
+We selected region-specific lag orders and Fourier harmonics via out-of-sample validation, with full results reported in Supplementary Table S1.
+Optimal specifications varied across regions: Regions 2 and 3 favored richer seasonal structure ($K_{\text{year}}=10$) with ARMA(1,1) dynamics, Region 1 performed best with fewer harmonics ($K_{\text{year}}=8$) and a pure autoregressive specification, while Region 4 required the most complex short-run dynamics at ARMA(3,2) despite fewer harmonics.
+These differences likely reflect heterogeneity in seasonal persistence and short-run adjustment across geographic markets.
 Aggregate performance.
-Table 4 reports Forecast Mean Absolute Error (FMAE) and Forecast Residual Sum of Squares (FRSS) for each region and currency across four models: B-DARCH, B-DARMA, B-tVARMA, and a time-varying‑parameter VARMA fit in ALR space (TVP-tVARMA).
-The corresponding out-of-sample forecasts $\bar{\mu}_{tj}$ are shown in Figures 36.
-When errors are aggregated over currencies within a region (the total rows of Table 4), B-DARCH attains the lowest total FMAE and FRSS in Regions 1–4.
-For Region 1, totals are FMAE/FRSS == 2.84/3.15 (B-DARCH) vs.
-7.65/19.11 (B-DARMA), 5.20/10.62 (B-tVARMA), and 4.55/8.12 (B-TVP-tVARMA).
-For Region 2, totals are 13.55/73.93 (B-DARCH) vs.
-14.71/81.87 (B-DARMA), 17.13/104.43 (B-tVARMA), and 14.68/90.49 (B-TVP-tVARMA).
-For Region 3 the ordering is 16.23/110.83 (B-DARCH) << 20.17/159.66 (B-DARMA) << 20.73/143.39 (B-TVP-tVARMA) << 23.70/187.09 (B-tVARMA).
-For Region 4, B-DARCH again leads with 11.21/32.55, followed by B-TVP-tVARMA (12.22/39.85), B-DARMA (15.25/62.33), and B-tVARMA (16.53/77.64).
-These results reinforce that explicitly modeling time‑varying volatility via DARCH yields the most accurate aggregate forecasts across regions.
-Across regions the B-TVP-tVARMA comparator required the longest runtime, followed by B-DARCH, while B-DARMA and B-tVARMA typically completed near one hour (Table S4).
+Table 4 compares forecast accuracy across four models: B-DARCH, B-DARMA, B-tVARMA, and a time-varying-parameter variant (TVP-tVARMA).
+Corresponding forecast trajectories for each region’s dominant currency appear in Figures 36.
+The central finding is that B-DARCH achieves the lowest aggregate FMAE and FRSS in all four regions.
+The gains are most pronounced in Region 1, where B-DARCH reduces total FMAE by over 60% relative to B-DARMA and FRSS by more than 80%.
+Regions 2–4 show more modest but consistent improvements.
+These results suggest that explicitly modeling time-varying volatility—the distinguishing feature of the DARCH specification—yields meaningful forecast gains, particularly in markets where volatility shifts are most pronounced.
+Among the comparators, TVP-tVARMA generally ranks second, indicating that time-varying parameters provide some benefit even without the DARCH volatility structure.
+However, this flexibility comes at computational cost: TVP-tVARMA required the longest runtimes, while B-DARMA and B-tVARMA typically completed in approximately one hour (Table S4).
 Per-currency comparisons.
-Currency-level minima vary by region.
-In Region 2, B-DARCH delivers the lowest FMAE for five of six currencies (CAD, CHF, EUR, USD, and other), with B‑TVP‑tVARMA best for GBP.
-In Region 3, B-DARCH is best for AUD, EUR, and other; B‑TVP‑tVARMA is best for JPY and NZD; and B-DARMA has the smallest USD FMAE.
-Region 4 shows a similar mix: B-DARCH leads for BRL, CLP, and other; B‑TVP‑tVARMA leads for MXN and USD; B-DARMA leads for EUR.
-In Region 1, B‑DARCH is best for CAD, USD, and other on both FMAE and FRSS; TVP‑tVARMA is best for EUR and JPY; and GBP is effectively a tie after rounding, with TVP‑tVARMA narrowly ahead on both metrics.
-Overall, the table indicates that the DARCH volatility component is beneficial across all regions, with especially large aggregate gains where volatility shifts are most pronounced.
-Figure 3: Airbnb data analysis — Region 1: 92-day forecasts (blue) with 95% credible intervals (shaded) for each currency from Oct.
+While B-DARCH dominates at the aggregate level, currency-level rankings are more nuanced (Table 4).
+B-DARCH achieves the lowest FMAE for the majority of individual currencies in each region—for instance, five of six currencies in Region 2.
+However, TVP-tVARMA occasionally outperforms for specific series, including GBP in Region 2, JPY and NZD in Region 3, and MXN and USD in Region 4.
+In Region 1, B-DARCH leads for CAD, USD, and the residual category, while TVP-tVARMA edges ahead for EUR, JPY, and GBP.
+B-DARMA is occasionally competitive at the currency level (e.g., USD in Region 3, EUR in Region 4) but never leads in aggregate.
+These patterns suggest that DARCH’s volatility modeling provides broad benefits across currency compositions, while time-varying mean parameters may better capture dynamics for particular series.
+Readers interested in specific currencies can consult the detailed breakdown in Table 4.
+Figure 3: Airbnb data analysis — Region 1: 92-day forecasts (blue) with 95% credible intervals (shaded) for USD currency from Oct.
 1 to Dec.
@@
 Models shown: B-DARCH, B-DARMA, B-tVARMA, and B-TVP-tVARMA.
-Figure 4: Airbnb data analysis — Region 2: 92-day forecasts (blue) with 95% credible intervals (shaded) for each currency from Oct.
+Figure 4: Airbnb data analysis — Region 2: 92-day forecasts (blue) with 95% credible intervals (shaded) for EUR currency from Oct.
 1 to Dec.
@@
 Models shown: B-DARCH, B-DARMA, B-tVARMA, and B-TVP-tVARMA.
-Figure 5: Airbnb data analysis — Region 3: 92-day forecasts (blue) with 95% credible intervals (shaded) for each currency from Oct.
+Figure 5: Airbnb data analysis — Region 3: 92-day forecasts (blue) with 95% credible intervals (shaded) for AUD currency from Oct.
 1 to Dec.
@@
 Models shown: B-DARCH, B-DARMA, B-tVARMA, and B-TVP-tVARMA.
-Figure 6: Airbnb data analysis — Region 4: 92-day forecasts (blue) with 95% credible intervals (shaded) for each currency from Oct.
+Figure 6: Airbnb data analysis — Region 4: 92-day forecasts (blue) with 95% credible intervals (shaded) for BRL currency from Oct.
 1 to Dec.
@@
 Acknowledgement
-The authors thank Sean Wilson, Jess Needleman, Liz Medina, and Yuling Kuo for helpful discussions, Adam Liss and Ellie Mertz for championing the research.
+The authors thank Sean Wilson, Jess Needleman, Liz Medina, Yuling Kuo, Erica Savage, and Jackson Wang for helpful discussions, Adam Liss and Ellie Mertz for championing the research.
 Code and Data Availability
@@
 It can be found at our GitHub repository: GitHub repository.
+6 Supplementary Materials
+Alternative Log-Ratio Transformations
+Although we have used the additive log-ratio (ALR) link in equations (4), (5), and (6), other transformations to and from the simplex may be employed.
+A common choice is the centered log-ratio (clr), defined as
+${\textnormal{clr}}(\mathbf{y})\;=\;\Bigl(\,\ln\bigl(\tfrac{y_{1}}{g(\mathbf{y})}\bigr),\;\ln\bigl(\tfrac{y_{2}}{g(\mathbf{y})}\bigr),\;\dots,\;\ln\bigl(\tfrac{y_{J}}{g(\mathbf{y})}\bigr)\Bigr),$
+where $g(\mathbf{y})=(\,y_{1}\,y_{2}\,\cdots\,y_{J})^{1/J}$ is the geometric mean of $\mathbf{y}$.
+The clr mapping places all $J$ components in a transformed $\mathbb{R}^{J}$ space but introduces the constraint that their sum is zero.
+Another option is to use the Isometric Log-Ratio (ilr) transformation. For each index $j$, the ilr-transformed coordinate can be expressed as
+$z_{j}=\sqrt{\frac{r_{j}}{r_{j}+1}}\,\ln\biggl(\frac{g_{j}(\mathbf{y)}}{\bigl(\prod_{i\in H_{j}}y_{i}\bigr)^{1/r_{j}}}\biggr),$
+where $g_{j}(\mathbf{y})$ is defined as the geometric mean of a chosen subset $S_{j}\subseteq\mathbf{y}$, and $H_{j}$ is the complement of $S_{j}$ in $\mathbf{y}$.
+The value $r_{j}$ is the number of elements in $H_{j}$.
+The particular way in which $S_{j}$ and $H_{j}$ are selected depends on the specific application and how one wishes to interpret the ilr coordinates.
+These mappings are mutually related by linear transformations (egozcue2003isometric) and each provides a bijection from the (J1)(J-1)-dimensional simplex onto a (J1)(J-1)-dimensional subset of $\mathbb{R}^{J}$ or $\mathbb{R}^{J-1}$.
+Therefore, if the same priors and model assumptions are consistently adapted to each log-ratio space, the resulting compositional models remain equivalent—differing only in how one embeds the data in a Euclidean domain.
+Definition of the Partial Autocorrelation Function (PACF)
+The PACF at lag $k$ measures the partial correlation between the SSR at time $t$ and the SSR at time tkt-k, controlling for the effects of intermediate lags $1$ to k1k-1.
+It is defined as the last coefficient $\phi_{kk}$ in the autoregressive model of order $k$
+$\text{SSR}_{t}=\phi_{k1}\text{SSR}_{t-1}+\phi_{k2}\text{SSR}_{t-2}+\dots+\phi_{k,k-1}\text{SSR}_{t-(k-1)}+\phi_{kk}\text{SSR}_{t-k}+\varepsilon_{t}$ (8)
+where $\varepsilon_{t}$ is the white noise error term at time $t$. The coefficient $\phi_{kk}$ represents the PACF at lag $k$ and can be calculated recursively using the Yule-Walker equations
+$\phi_{kk}=\frac{r_{k}-\sum_{j=1}^{k-1}\phi_{k-1,j}r_{k-j}}{1-\sum_{j=1}^{k-1}\phi_{k-1,j}r_{j}}$
+Here, $r_{k}$ is the autocorrelation at lag $k$, and $\phi_{k-1,j}$ are the PACF coefficients from the previous lag k1k-1.
+By examining the PACF of the SSRs, we can identify the presence of residual autocorrelation at various lags, which indicates whether the model has adequately captured the temporal dependencies in the data.
+Standardized Residuals
+For the B-DARMA and B-DARCH models, which are based on the Dirichlet distribution as defined in Section 2, we compute standardized residuals using the standard deviations derived from the Dirichlet parameters.
+At each time t, the residuals are calculated as $\text{residual}_{tj}=y_{tj}-\mu_{tj},\quad\text{for }j=1,\dots,J.$ The variance of each component $y_{tj}$ under the Dirichlet distribution, Dirichlet($\phi_{t}\boldsymbol{\mu}_{t}$), is $\operatorname{Var}[y_{tj}]=\frac{\mu_{tj}(1-\mu_{tj})}{\phi_{t}+1}$, with standard deviation $\sigma_{tj}=\sqrt{\frac{\mu_{tj}(1-\mu_{tj})}{\phi_{t}+1}}.$
+The standardized residuals are then computed as $\text{standardized\_residual}_{tj}=\frac{y_{tj}-\mu_{tj}}{\sigma_{tj}}$ and the sum of squared standardized residuals at time t is $\text{SSR}_{t}=\sum_{j=1}^{J}\left(\text{standardized\_residual}_{tj}\right)^{2}.$
+For the B-tVARMA and B-TVP-tVARMA models, residuals are computed in the additive log-ratio (alr) transformed space, $\text{residual}_{tj}={\textnormal{alr}}(\mathbf{y}_{tj})-{\textnormal{alr}}(\mathbf{\hat{y}}_{tj}),$ where $j=1,\dots,4$.
+These residuals are standardized using the inverse of the Cholesky factor $L_{\Sigma}$ of the estimated covariance matrix $\Sigma$$\text{standardized\_residual}_{t}=L\Sigma^{-1}\times\text{residual}_{t}$.
+The sum of squared standardized residuals at time t is then calculated as $\text{SSR}_{t}=\sum_{j=1}^{J-1}(\text{standardized\_residual}_{tj})^{2}.$
+By examining the PACF of $\{\text{SSR}\}_{t=1}^{T}$ we assess the presence of autocorrelation in the residuals.
+A well-specified model should produce residuals with minimal autocorrelation, indicating that temporal dependencies have been adequately captured.
+Parameter Estimates
+Supplementary Figure S13 shows the posterior densities of the auto-regressive ($\alpha$) and moving average ($\tau$) coefficients for each region.
+In Region 1, the posterior of $\alpha$ is concentrated between 0.25 and 0.75, peaking around 0.45, indicating moderate volatility persistence; shocks have a noticeable but not prolonged effect on volatility.
+In Region 2, the $\alpha$ coefficient has posterior mean around 0.8, reflecting strong dependence of current volatility on past volatility, and the $\tau$ coefficient centers around 2.75-2.75, indicating that large deviations from expected currency proportions significantly increase / decrease future volatility—characteristic of volatility clustering.
+Region 3 shows $\alpha$ tightly concentrated between 0.85 and 0.90 and $\tau$ centered around 1.2-1.2 suggesting very strong volatility persistence.
+In contrast, Region 4 exhibits complex dynamics: $\alpha_{1}$ is centered around $0.15$, indicating small persistence from the first lag; $\alpha_{2}$ is skewed positive and centered around $0.35$, suggesting a stronger influence from the second lag; $\alpha_{3}$ peaks slightly above 0.5, indicating moderate persistence; and the $\tau$ parameter is concentrated near $0.02$, implying minimal influence of past innovations.
+Overall, the higher $\alpha$ values in Regions 2 and 3 indicate that volatility in currency fee proportions is more persistent in these regions.
+The negative $\tau$ coefficients suggest that large deviations—such as those caused by abrupt travel restrictions, policy changes, and fluctuating foreign exchange (FX) rates lead to increased future volatility.
+This implies that Regions 2 and 3 are more susceptible to external shocks affecting travel and currency markets, resulting in prolonged volatility in currency fee proportions.
+In contrast, Region 1, with moderate volatility persistence, tends to return to stability more rapidly after shocks, possibly due a more stable currency environment or a larger share of domestic travel.
+Region 4’s dynamics, characterized by less emphasis on immediate past volatility and innovations, may be attributed to regional factors such as diversified travel sources, differing policy responses, or economic structures.
+Figure S1: Average PACF of the sum of squared standardized residual values across all 50 simulations for studies 1 & 2.
+B-DARCH consistently exhibits lower PACF values beyond lag 1, indicating reduced residual autocorrelation and suggesting that its dynamic precision component better captures volatility shocks in these simulated compositional time series.
+Figure S2: Average PACF of the sum of squared standardized residual values across all 50 simulations for studies 3 & 4.
+The B-DARCH model (red) again shows smaller autocorrelation at higher lags, whereas B-DARMA (green) and B-tVARMA (red) leave more persistent structures in the residuals.
+This result underlines the advantage of modeling time-varying volatility when faced with abrupt changes in compositional data.
+Figure S3: Airbnb data analysis- Proportion of fees by currency for four regions: weekly seasonal behavior.
+AUD is the Australian dollar, BRL is the Brazillian Real, CAD is the Canadian Dollar,CHF is the Swiss Franc, CLP is the Chilean Peso,EUR is the European Euro, GBP is the Great British Pound, MXN is the Mexican Peso, NZD is the New Zealand Dollar, and USD is the US Dollar.
+Strong weekly cycles appear in many currencies, motivating the inclusion of weekly seasonal terms in our model’s mean structure.
+Figure S4: Airbnb data analysis - proportion of fees by billing currency for four regions from Jan 1, 2020 to June 30, 2020.
+AUD is the Australian dollar, BRL is the Brazillian Real, CAD is the Canadian Dollar,CHF is the Swiss Franc, CLP is the Chilean Peso,EUR is the European Euro, GBP is the Great British Pound, MXN is the Mexican Peso, NZD is the New Zealand Dollar, and USD is the US Dollar.
+Distinct spikes and dips across multiple currencies (e.g., sharp increase in BRL or EUR) reflect significant disruptions.
+Figure S5: Airbnb data analysis - proportion of fees by currency- 30-day rolling ALR means for four regions from Jan 1, 2020 to June 30, 2020. USD is the reference component.
+Figure S6: Airbnb data analysis - proportion of fees by currency- 30-day rolling ALR means for four regions from Jan 1, 2017 to June 30, 2020.
+USD is the reference component.
+Apart from seasonal patterns, we see notable multi-year trends and cyclical behaviors (e.g., increasing share of EUR in Region 2, upward drift in AUD in Region 3), underlining the presence of both trend and seasonality in compositional dynamics.
+Figure S7: Airbnb data analysis - proportion of fees by currency- 30-day rolling ALR variance for four regions from Jan 1, 2017 to June 30, 2020.
+USD is the reference component.The data display periods of high variance that sometimes persist for weeks, demonstrating “volatility clustering.” These episodes justify the GARCH-like specification in our B-DARCH model, which allows the precision parameter $\phi_{t}$ to track spikes in uncertainty over time.
+Figure S8: Airbnb data analysis- correlation between ALR and lagged ALR values for each currency in the four regions.
+USD is the reference component.
+High positive autocorrelations in many currencies (e.g., EUR, GBP) highlight the strong temporal persistence.
+Negative cross-lag correlations for some pairs (e.g., EUR vs.
+BRL in Region 4) indicate that certain currencies move in opposite directions.
+Both effects reinforce the need for a multi-component time-series framework with auto- and cross-dependence.
+Figure S9: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 1.
+Figure S10: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 2.
+Figure S11: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 3.
+Figure S12: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 4.
+Figure S13: Airbnb data analysis: Densities of DARCH model coefficients for all 4 regions.
+Region 1 was modeled with a DARCH(1,0), Regions 2-3 modeled with a DARCH(1,1), and Region 4 modeled with a DARCH(3,1).
+Each plot shows the estimated density of a specific $\alpha_{l}$ (ar_phi) or $\tau_{k}$ (ma_phi) parameter.
+Regions 2 and 3 exhibit high $\alpha$ indicating persistent volatility once spikes occur, whereas Region 4’s coefficients suggest moderate or delayed response
 Supplementary Figures and Tables

### v4

@@
 In marketplace finance, the daily mix of billing currencies is compositional data that drives forecasting, reporting, and treasury risk.
-We study Airbnb’s 2017–2020 currency‑fee shares across four regions and present a Bayesian Dirichlet ARMA model with a Dirichlet ARCH volatility component, B‑DARMA‑DARCH.
-The model keeps predictions on the simplex, models mean dynamics with VARMA terms on the additive log‑ratio scale, and lets precision vary over time so volatility can spike during disruptions and settle as conditions normalize.
-We include weekly and yearly seasonality and a trend, and we evaluate against Dirichlet ARMA and transformed‑Gaussian VARMA across simulation designs with misreported observations and temporary regime shifts, then validate on a held‑out window and test in 2020 Q4.
-Across simulations and the application, B‑DARMA‑DARCH delivers the most accurate aggregate forecasts, better‑calibrated intervals, and weaker residual persistence at medium lags.
+We study Airbnb’s currency-fee shares across four regions and present a Bayesian Dirichlet ARMA model with a time-varying precision component.
+The model keeps predictions on the simplex, captures mean dynamics on the additive log-ratio scale, and lets volatility spike during disruptions and settle as conditions normalize.
+We evaluate against standard Dirichlet and transformed-Gaussian alternatives using simulations with misreported observations and temporary regime shifts, then validate on held-out data.
+Across all settings, our approach delivers more accurate forecasts, better-calibrated intervals, and weaker residual persistence.
 Modeling precision as a dynamic process provides a practical, interpretable way to forecast proportions and quantify uncertainty when the noise itself moves.
@@
 Each day’s currency proportions must lie in a simplex (they are nonnegative and sum to one), a constraint that complicates standard time-series methods.
-One classical approach is to use log-ratio transformations, as pioneered by aitchison1982statistical and extended in works such as Cargnoni1997BayesianFO, ravishanker2001compositional, mills2010forecasting, and snyder2017forecasting, among others.
-Transformed additive log-ratio VARMA (tVARMA) models have been fruitfully applied to compositional data (brunsdon1998time; mills2009forecasting; mills2010forecasting).
+One classical approach is to use log-ratio transformations, as pioneered by Aitchison (1982) and extended in works such as Cargnoni,Claudia et al.
+(1997), Ravishanker et al.
+(2001), Mills (2010), and Snyder et al.
+(2017), among others.
+Transformed additive log-ratio VARMA (tVARMA) models have been fruitfully applied to compositional data (Brunsdon and Smith, 1998; Mills, 2009, 2010).
 Let $Q_{t,c}$ denote gross booking amount in currency $c$ on day $t$, and let $\mathrm{FX}_{t,c\rightarrow USD}$ be the USD price of one unit of $c$. Consolidated revenue in USD can be written as
@@
 In contrast, the Dirichlet distribution models compositional data directly on the simplex.
-benjamin2003generalized proposed frequentist Generalized ARMA models that could be extended to the simplex.
-Contributions to Dirichlet time series include Bayesian compositional state space models by grunwald1993time, da2011dynamic, and da2015bayesian; the Bayesian Dirichlet Auto-Regressive Moving Average (B-DARMA) model by KATZ20241556; forecast7030032; forecast7040062; katz2025centeredmadirichletarma; and frequentist Dirichlet ARMA models by zheng2017dirichlet; creus2021dirichlet.
-The generalized auto-regressive conditional heteroskedasticity (GARCH) model (bollerslev1986generalized) is fundamental in modeling financial time series volatility.
-Extensions include non-normal likelihoods to capture heavy tails and skewness (nelson1991conditional; engle2001theoretical).
-bauwens2006multivariate surveyed advancements in multivariate GARCH models.
+Benjamin et al.
+(2003) proposed frequentist Generalized ARMA models that could be extended to the simplex.
+Contributions to Dirichlet time series include Bayesian compositional state space models by Grunwald et al.
+(1993), da-Silva et al.
+(2011), and da-Silva and Rodrigues (2015); the Bayesian Dirichlet Auto-Regressive Moving Average (B-DARMA) model by Katz et al.
+(2024, 2025); Katz and Maierhofer (2025); Katz (2025); and frequentist Dirichlet ARMA models by Zheng and Chen (2017); Creus-Martí et al.
+(2021).
+The generalized auto-regressive conditional heteroskedasticity (GARCH) model (Bollerslev, 1986) is fundamental in modeling financial time series volatility.
+Extensions include non-normal likelihoods to capture heavy tails and skewness (Nelson, 1991; Engle, 2001).
+Bauwens et al.
+(2006) surveyed advancements in multivariate GARCH models.
 Motivated by the need to capture both shifting means and evolving volatility in compositional time series—especially under sudden disruptions such as COVID-19—we develop a new class of models called Bayesian Dirichlet Auto-Regressive Moving Average with Dirichlet Auto-Regressive Conditional Heteroskedasticity (B-DARCH).
@@
 Acknowledgement
-The authors thank Sean Wilson, Jess Needleman, Liz Medina, Yuling Kuo, Erica Savage, and Jackson Wang for helpful discussions, Adam Liss and Ellie Mertz for championing the research.
+The authors thank Sean Wilson, Jess Needleman, Liz Medina, Yuling Kuo, and Jackson Wang for helpful discussions, Adam Liss and Ellie Mertz for championing the research, and the Editor-in-Chief, Associate Editors, and two anonymous reviewers for constructive feedback that improved the manuscript.
 Code and Data Availability
 The primary dataset used in our study, A Bayesian Dirichlet Auto-Regressive Conditional Heteroskedasticity Model for Forecasting Currency Shares, is not publicly available due to confidentiality constraints.
-However, the Stan code for the Bayesian Dirichlet Auto-Regressive Moving Average Dirichlet Auto-Regressive Conditional Heteroskedasticity Model (B-DARMA-DARCH) model is available for public access.
+However, the Stan code for the Bayesian Dirichlet Auto-Regressive Moving Average Dirichlet Auto-Regressive Conditional Heteroskedasticity Model (B-DARMA-DARCH) model is available for public access as well as the R scripts for the simulation studies.
 It can be found at our GitHub repository: GitHub repository.
-6 Supplementary Materials
-Alternative Log-Ratio Transformations
-Although we have used the additive log-ratio (ALR) link in equations (4), (5), and (6), other transformations to and from the simplex may be employed.
-A common choice is the centered log-ratio (clr), defined as
-${\textnormal{clr}}(\mathbf{y})\;=\;\Bigl(\,\ln\bigl(\tfrac{y_{1}}{g(\mathbf{y})}\bigr),\;\ln\bigl(\tfrac{y_{2}}{g(\mathbf{y})}\bigr),\;\dots,\;\ln\bigl(\tfrac{y_{J}}{g(\mathbf{y})}\bigr)\Bigr),$
-where $g(\mathbf{y})=(\,y_{1}\,y_{2}\,\cdots\,y_{J})^{1/J}$ is the geometric mean of $\mathbf{y}$.
-The clr mapping places all $J$ components in a transformed $\mathbb{R}^{J}$ space but introduces the constraint that their sum is zero.
-Another option is to use the Isometric Log-Ratio (ilr) transformation. For each index $j$, the ilr-transformed coordinate can be expressed as
-$z_{j}=\sqrt{\frac{r_{j}}{r_{j}+1}}\,\ln\biggl(\frac{g_{j}(\mathbf{y)}}{\bigl(\prod_{i\in H_{j}}y_{i}\bigr)^{1/r_{j}}}\biggr),$
-where $g_{j}(\mathbf{y})$ is defined as the geometric mean of a chosen subset $S_{j}\subseteq\mathbf{y}$, and $H_{j}$ is the complement of $S_{j}$ in $\mathbf{y}$.
-The value $r_{j}$ is the number of elements in $H_{j}$.
-The particular way in which $S_{j}$ and $H_{j}$ are selected depends on the specific application and how one wishes to interpret the ilr coordinates.
-These mappings are mutually related by linear transformations (egozcue2003isometric) and each provides a bijection from the (J1)(J-1)-dimensional simplex onto a (J1)(J-1)-dimensional subset of $\mathbb{R}^{J}$ or $\mathbb{R}^{J-1}$.
-Therefore, if the same priors and model assumptions are consistently adapted to each log-ratio space, the resulting compositional models remain equivalent—differing only in how one embeds the data in a Euclidean domain.
-Definition of the Partial Autocorrelation Function (PACF)
-The PACF at lag $k$ measures the partial correlation between the SSR at time $t$ and the SSR at time tkt-k, controlling for the effects of intermediate lags $1$ to k1k-1.
-It is defined as the last coefficient $\phi_{kk}$ in the autoregressive model of order $k$
-$\text{SSR}_{t}=\phi_{k1}\text{SSR}_{t-1}+\phi_{k2}\text{SSR}_{t-2}+\dots+\phi_{k,k-1}\text{SSR}_{t-(k-1)}+\phi_{kk}\text{SSR}_{t-k}+\varepsilon_{t}$ (8)
-where $\varepsilon_{t}$ is the white noise error term at time $t$. The coefficient $\phi_{kk}$ represents the PACF at lag $k$ and can be calculated recursively using the Yule-Walker equations
-$\phi_{kk}=\frac{r_{k}-\sum_{j=1}^{k-1}\phi_{k-1,j}r_{k-j}}{1-\sum_{j=1}^{k-1}\phi_{k-1,j}r_{j}}$
-Here, $r_{k}$ is the autocorrelation at lag $k$, and $\phi_{k-1,j}$ are the PACF coefficients from the previous lag k1k-1.
-By examining the PACF of the SSRs, we can identify the presence of residual autocorrelation at various lags, which indicates whether the model has adequately captured the temporal dependencies in the data.
-Standardized Residuals
-For the B-DARMA and B-DARCH models, which are based on the Dirichlet distribution as defined in Section 2, we compute standardized residuals using the standard deviations derived from the Dirichlet parameters.
-At each time t, the residuals are calculated as $\text{residual}_{tj}=y_{tj}-\mu_{tj},\quad\text{for }j=1,\dots,J.$ The variance of each component $y_{tj}$ under the Dirichlet distribution, Dirichlet($\phi_{t}\boldsymbol{\mu}_{t}$), is $\operatorname{Var}[y_{tj}]=\frac{\mu_{tj}(1-\mu_{tj})}{\phi_{t}+1}$, with standard deviation $\sigma_{tj}=\sqrt{\frac{\mu_{tj}(1-\mu_{tj})}{\phi_{t}+1}}.$
-The standardized residuals are then computed as $\text{standardized\_residual}_{tj}=\frac{y_{tj}-\mu_{tj}}{\sigma_{tj}}$ and the sum of squared standardized residuals at time t is $\text{SSR}_{t}=\sum_{j=1}^{J}\left(\text{standardized\_residual}_{tj}\right)^{2}.$
-For the B-tVARMA and B-TVP-tVARMA models, residuals are computed in the additive log-ratio (alr) transformed space, $\text{residual}_{tj}={\textnormal{alr}}(\mathbf{y}_{tj})-{\textnormal{alr}}(\mathbf{\hat{y}}_{tj}),$ where $j=1,\dots,4$.
-These residuals are standardized using the inverse of the Cholesky factor $L_{\Sigma}$ of the estimated covariance matrix $\Sigma$$\text{standardized\_residual}_{t}=L\Sigma^{-1}\times\text{residual}_{t}$.
-The sum of squared standardized residuals at time t is then calculated as $\text{SSR}_{t}=\sum_{j=1}^{J-1}(\text{standardized\_residual}_{tj})^{2}.$
-By examining the PACF of $\{\text{SSR}\}_{t=1}^{T}$ we assess the presence of autocorrelation in the residuals.
-A well-specified model should produce residuals with minimal autocorrelation, indicating that temporal dependencies have been adequately captured.
-Parameter Estimates
-Supplementary Figure S13 shows the posterior densities of the auto-regressive ($\alpha$) and moving average ($\tau$) coefficients for each region.
-In Region 1, the posterior of $\alpha$ is concentrated between 0.25 and 0.75, peaking around 0.45, indicating moderate volatility persistence; shocks have a noticeable but not prolonged effect on volatility.
-In Region 2, the $\alpha$ coefficient has posterior mean around 0.8, reflecting strong dependence of current volatility on past volatility, and the $\tau$ coefficient centers around 2.75-2.75, indicating that large deviations from expected currency proportions significantly increase / decrease future volatility—characteristic of volatility clustering.
-Region 3 shows $\alpha$ tightly concentrated between 0.85 and 0.90 and $\tau$ centered around 1.2-1.2 suggesting very strong volatility persistence.
-In contrast, Region 4 exhibits complex dynamics: $\alpha_{1}$ is centered around $0.15$, indicating small persistence from the first lag; $\alpha_{2}$ is skewed positive and centered around $0.35$, suggesting a stronger influence from the second lag; $\alpha_{3}$ peaks slightly above 0.5, indicating moderate persistence; and the $\tau$ parameter is concentrated near $0.02$, implying minimal influence of past innovations.
-Overall, the higher $\alpha$ values in Regions 2 and 3 indicate that volatility in currency fee proportions is more persistent in these regions.
-The negative $\tau$ coefficients suggest that large deviations—such as those caused by abrupt travel restrictions, policy changes, and fluctuating foreign exchange (FX) rates lead to increased future volatility.
-This implies that Regions 2 and 3 are more susceptible to external shocks affecting travel and currency markets, resulting in prolonged volatility in currency fee proportions.
-In contrast, Region 1, with moderate volatility persistence, tends to return to stability more rapidly after shocks, possibly due a more stable currency environment or a larger share of domestic travel.
-Region 4’s dynamics, characterized by less emphasis on immediate past volatility and innovations, may be attributed to regional factors such as diversified travel sources, differing policy responses, or economic structures.
-Figure S1: Average PACF of the sum of squared standardized residual values across all 50 simulations for studies 1 & 2.
-B-DARCH consistently exhibits lower PACF values beyond lag 1, indicating reduced residual autocorrelation and suggesting that its dynamic precision component better captures volatility shocks in these simulated compositional time series.
-Figure S2: Average PACF of the sum of squared standardized residual values across all 50 simulations for studies 3 & 4.
-The B-DARCH model (red) again shows smaller autocorrelation at higher lags, whereas B-DARMA (green) and B-tVARMA (red) leave more persistent structures in the residuals.
-This result underlines the advantage of modeling time-varying volatility when faced with abrupt changes in compositional data.
-Figure S3: Airbnb data analysis- Proportion of fees by currency for four regions: weekly seasonal behavior.
-AUD is the Australian dollar, BRL is the Brazillian Real, CAD is the Canadian Dollar,CHF is the Swiss Franc, CLP is the Chilean Peso,EUR is the European Euro, GBP is the Great British Pound, MXN is the Mexican Peso, NZD is the New Zealand Dollar, and USD is the US Dollar.
-Strong weekly cycles appear in many currencies, motivating the inclusion of weekly seasonal terms in our model’s mean structure.
-Figure S4: Airbnb data analysis - proportion of fees by billing currency for four regions from Jan 1, 2020 to June 30, 2020.
-AUD is the Australian dollar, BRL is the Brazillian Real, CAD is the Canadian Dollar,CHF is the Swiss Franc, CLP is the Chilean Peso,EUR is the European Euro, GBP is the Great British Pound, MXN is the Mexican Peso, NZD is the New Zealand Dollar, and USD is the US Dollar.
-Distinct spikes and dips across multiple currencies (e.g., sharp increase in BRL or EUR) reflect significant disruptions.
-Figure S5: Airbnb data analysis - proportion of fees by currency- 30-day rolling ALR means for four regions from Jan 1, 2020 to June 30, 2020. USD is the reference component.
-Figure S6: Airbnb data analysis - proportion of fees by currency- 30-day rolling ALR means for four regions from Jan 1, 2017 to June 30, 2020.
-USD is the reference component.
-Apart from seasonal patterns, we see notable multi-year trends and cyclical behaviors (e.g., increasing share of EUR in Region 2, upward drift in AUD in Region 3), underlining the presence of both trend and seasonality in compositional dynamics.
-Figure S7: Airbnb data analysis - proportion of fees by currency- 30-day rolling ALR variance for four regions from Jan 1, 2017 to June 30, 2020.
-USD is the reference component.The data display periods of high variance that sometimes persist for weeks, demonstrating “volatility clustering.” These episodes justify the GARCH-like specification in our B-DARCH model, which allows the precision parameter $\phi_{t}$ to track spikes in uncertainty over time.
-Figure S8: Airbnb data analysis- correlation between ALR and lagged ALR values for each currency in the four regions.
-USD is the reference component.
-High positive autocorrelations in many currencies (e.g., EUR, GBP) highlight the strong temporal persistence.
-Negative cross-lag correlations for some pairs (e.g., EUR vs.
-BRL in Region 4) indicate that certain currencies move in opposite directions.
-Both effects reinforce the need for a multi-component time-series framework with auto- and cross-dependence.
-Figure S9: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 1.
-Figure S10: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 2.
-Figure S11: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 3.
-Figure S12: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 4.
-Figure S13: Airbnb data analysis: Densities of DARCH model coefficients for all 4 regions.
-Region 1 was modeled with a DARCH(1,0), Regions 2-3 modeled with a DARCH(1,1), and Region 4 modeled with a DARCH(3,1).
-Each plot shows the estimated density of a specific $\alpha_{l}$ (ar_phi) or $\tau_{k}$ (ma_phi) parameter.
-Regions 2 and 3 exhibit high $\alpha$ indicating persistent volatility once spikes occur, whereas Region 4’s coefficients suggest moderate or delayed response
 Supplementary Figures and Tables